A Complete Bibliography of Publications in Journal of Fish Diseases (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.04

Title word cross-reference

14 [MSGA95]. _{1} [SB94]. α [LWHD97, SNA^{+}92]. α_{2} [SNØ^{+}93]. β [BWC92, DIB^{+}95, DS95, DISS96, DBIS96, DMH^{+}98, SD97b]. × [PCKS94].

-D-glucan [DIB^{+}95, DS95, DISS96, DBIS96, DMH^{+}98]. -galactosidase [BWC92]. -glucan [SD97b]. -like [SB94]. -macroglobulin [SNØ^{+}93]. -malachite [MSGA95]. -naphthyl [LWHD97]. -polymerase [LLGJS94].

001 [ZFW97].

1 [CHL99]. 16s [MHM95]. 18S [HYHA98]. 1990 [Bro91].

2-macroglobulin [SNA^{+}92]. 20E [BEGA93, GCO91]. 21-hemisuccinate [HM90]. 214 [WHC^{+}98]. 299 [HCW93].

37^{o} [Aus93].

47- [VW96]. 470 [HK98].

6-year [SQTF90].

70 [BWC92, WDCM92]. 70-kDa [BWC92, WDCM92].

A-layer [BENJ92, TGK94]. A-protein [SABòJ99]. A. [HYHA98, NE91]. abalone [DHB91]. abandoned [HLJ^{+}91]. abdominal [PMS98, VD99]. ability [Rit97]. Abramis [KT93a]. absence [FDS92]. Absorption [SD97b, DIHS94, DLB95]. acanthias [FDS92]. Acanthopagrus [Rou94b, Rou94a]. accompanying [LGL^{+}98]. Accumulation [DIB^{+}95, DISS96, Bro90]. acetate [LWHD97]. acetonide [HM90]. acetylcholinesterase [RFN93]. acheilognathi [HN94]. achromogenes [GJS^{+}97]. acid [LH90, SE99, WHHL95, WFS^{+}95]. Acquired [SB93, BL96, FM98, HM90]. Actin [LGL^{+}98]. actinosporean [MEMSB99, YDOW97]. activated [WCS90]. activation [EBL^{+}94]. active [AMW96, LWHD97]. activities [KGW90, LH90, SJH^{+}93]. activity [AOES99, BAHL91, CW93, DB96, HCM92, ISM^{+}98, NDBL94, OSG95]. aculeatus [JK90]. acutely [GWG99]. acyltransferase [DB96, EBL^{+}94, LWHD97, RSTE93, SNA^{+}92]. Adeleorina [DEA94]. Adherence [DSB90, KSPO99]. adhesin [FLT^{+}98]. adhesins [LYGS97]. adhesion [DHV^{+}99, DHTC99, YIM^{+}97]. adhesions [CSP98, VD99]. adjuvanted [LLP92, SEMB^{+}99]. adjuvants [CYA^{+}98]. administered [CMAM96, KEY91]. administering [SLJ97]. administration [BMEC97, CRM^{+}96, CA92, DBIS96, ISM^{+}98, IKU^{+}96, Kak98, PMD90, PRL95, PERL96, SOAK93, SYAK95, YKI95]. adult [JBWT92, SDS95, WHHL95]. advantage [MF91]. Aegean [DAKD98]. aequans [CSBCV98]. Aeromona [EA91]. aeromonads [DSB90]. Aeromonas [AHTY92, AH91, ABJE93, AEJ95, AE99, Aus93, ASR^{+}95, BAHL91, BLHA92, BHA94, BWC92, Ber90, Ber91, BENJ92, BHP^{+}94, BBS96, CDPS90, CCH^{+}94, CT91, CCD95, DMH^{+}98, DSA^{+}98, DWV94, DES92, DPPE96, DPPE97, DVB^{+}98, DB96, DAKD98, EA94, EBL^{+}94, Ell91, EGTH92, FLT^{+}98, FH91, Gat91, GRR92, GJS^{+}97, HMH96, HHE94, HWGV91, HCM92, IFMR91, IR91, IYBF93, IMR93, IRM^{+}96, KSS91, KRK91, KON90, LWHD97, LTGC99, LE91a, LYGS97, LLLS95, LYLS95, LLLS96, LH90, LSAB99, LGL^{+}98, LJHE91, MGE90, MF91, Møy93, MLB95, NHOJ93, Ner96, NDBL94, NE91, NRV^{+}98, NBWO91, PRCC96, PBC^{+}90, RT94, RV91, RFN93, REM90, RSTE93, SNØ91, SNA^{+}92, SNØ^{+}93, SH92, SKQ^{+}99, SD93, SABòJ99, SDB99, Tat91, TGK94, TL94, WLHD97, WDCM92, WDV93, Wik90, YIA92]. Aeromonashydrophila [FLT^{+}98]. aeruginosa [CRG^{+}97]. aetiological [AL95, CFS93, DFS94]. aetiology [SRRN94]. affected [Cai92, WRC95]. affecting [AR95, AA91, BLSC93, CNN^{+}94, OBL^{+}99, SH92]. aflatoxin [SB94]. AFP [SB94]. Africa [CGA93]. African [DHB91, PBP94, WGFF98, YKI95]. after [ARX95, DBIS96, DF94, LSR90, MSGA95, PMD90]. against [AMW96, ABJE93, BAR97b, BAHL91, BSBD99, BSHJ91, BR99, BCD90, CW93, CFS93, DMH^{+}98, DWV94, DAL90, FW96, FAAEC99, FMY94, GG97, GJS^{+}97, GMHK93, IRV^{+}91, IRM^{+}96, IKU^{+}96, LYD94, LH90, LW97, LSR90, MC94, NS96, NVSB94, NRV^{+}98, OSG95, PV93, SBN^{+}91, SBB^{+}93, SDS90, SK97, SBW94, TW90b, TSL^{+}94, VW96, WCS90, WR92, YMM91, YLSW99, ZH96, ZFW97]. AGD [FM98]. age [MCRF94, SDS90]. agent [ARD92, AL95, AMN^{+}93, BAR97b, BHB90, DFS94, MOBF98, SD97a, SCBB93, VC93, ZBM^{+}99]. agents [CW93, IR91, PI93, Rod91]. agglomerans [HRO90]. Aggregata [ERSA98]. Aggregatidae [ERSA98]. akaara [CLK^{+}97]. al [LS93]. albidus [LT92]. Alcohol [vdIOK96]. Alcohol-soluble [vdIOK96]. algal [RTEC94]. alginates [FMY94]. alginolyticus [ASR^{+}95, ASR^{+}93]. alimentary [DMOR97]. alpha [SB94]. alpinus [KRL96]. Alterations [BLHA92, SEMB^{+}99]. alternate [MEMSB99]. altivelis [KEY91]. alum [MLB95]. aluminium [MLB95]. alvei [RGA^{+}98]. amago [TMMA96b, TMMA96a]. Amazonian [AMM95]. Ambassis [HR95]. America [BDM91]. americana [BP91]. aminopeptidolytic [SJH^{+}93]. Ammodytidae [LPO^{+}99]. amoeba [LPO^{+}98, PFO^{+}98]. amoebic [BLW95, CN99, FM98]. among [ASR^{+}93, BHP^{+}94, DTS^{+}92]. amoxicillin [IMR93]. Amoxycillin [BHA94]. amplification [MEJ^{+}94, TIM^{+}96]. amplified [MHM95]. amylase [CDPS90]. Amyloidosis [MCL97]. anaemia [DFK93, DF94, DFS94, HNWE94, MF95, NHWE95, NKKH97, SM96, TW90a]. analogue [HK98]. analyses [KRSE93]. Analysis [KEY91, BSH^{+}90, DGR97, Dav91b, DPPE96, LSAB99, LTKC93, MCRF94, MC94, RM95b, SA94]. Anarhichas [HMH96]. anfuilla [EA91]. Angas [AAL95]. angelfish [LHS94, RKMF97]. anguiliarum [GRR92]. Anguilla [AMNH99, BAE^{+}91, BRW92, CD98, DMOR97, EA91, GMHK93, HAH92, HCW93, HLTK96, HLTK97a, HLTK97b, LKM99, SB93, ZPVH97]. anguillae [LKM99]. anguillarum [ASR^{+}95, MB92, BSH^{+}90, BSHJ91, ERJ91, GCO91, HLTK96, HLTK97b, KRSE93, LPD94, OJW^{+}96, OJW^{+}98, PLLT94, SJH^{+}93]. anguillicaudatus [CW91]. Anguillicola [HYHA98, HAH92]. anguilliseptica [DFGG^{+}99]. anion [KSS91]. ANNOUNCEMENT [Ano90a, Ano91a]. Anoplodiscidae [RQW92]. Anoplodiscus [RQW92, WR98]. Anoplopoma [BHB90]. Anorexia [TW92]. Anostraca [DVB^{+}98]. anovulatory [KRL96]. antagonized [BHA95]. Antarctic [FMDC93]. anterior [Ain93, DBIS96]. anti [FW98, GMHK93]. anti- [GMHK93]. anti-thymocyte [FW98]. antibacterial [PI93]. antibacterials [BHA95, DHB^{+}95]. Antibiotic [IFMR91, IRM^{+}96, AHTY92, AB90, BAR95, HLJ^{+}91, IYBF93, NDBL94]. antibiotic-producing [AB90]. antibiotic-resistant [HLJ^{+}91]. antibodies [ARD92, BAR97b, CND90, FW98, FAAEC99, HAC^{+}92, LAB92, MRD98, MCRF94, MC94, NS96, SBB^{+}93, SA94, SBW94, WLHD97]. Antibody [BSHJ91, HLTK97a, HLTK97b, RMR95, BR99, CCH^{+}94, CFS93, DES92, FW96, GJS^{+}97, GMHK93, HIMK93, HM90, OCC^{+}95, SAK91, VW96, WBA91, WCS90, ZH96, ZFW97, dMMMK96]. antibody-secreting [DES92, WBA91]. Antigen [PRL95, ARX95, DLB95, FW96, GL91, RMR95, TGK94, VW96, WPRM94]. antigen-capture [VW96]. Antigenic [DTS^{+}92, MCRF94, MC94, THBL96]. antigenically [AKK96]. antigenicity [BWC92, CATR97]. antigens [BSH^{+}90, BLW95, LES^{+}93, LJHE91, Tat91]. antimicrobial [CW93, IR91, OSG95, Rod91, SCBB93]. antimicrobials [SSM95]. Antioxidant [CGøL^{+}95]. antiperoxidase [JHL^{+}91]. antisera [FH91]. antithrombin [SNA^{+}92, SNØ^{+}93]. aorta [BMEC97]. Aphanomyces [CRWP95, FCC92, LR97, WR94, WRC95]. API [BEGA93, GCO91]. Apicomplexa [DEA94, ED91, ERSA98, Jon90, PI91]. apoptosis [HLHW98]. appearance [TRR96]. Application [CFS93, BR99, F.92]. appraisal [Ell91]. approach [LSL91]. Aquaculture [BHA95, APR99, HSP97, HLJ^{+}91, MB97, RRS93, SRL^{+}96]. aquareovirus [LHS94]. aquarium [ASR^{+}93, PI91]. aquarium-reared [PI91]. aquatic [BSM^{+}99, MCRF94, MC94, PI93]. arches [DHTC99]. Arctic [KRL96]. area [SBAP90]. areas [ALVS94]. arginine [NHOJ93]. Argulus [RMR95]. Armitage [RQW92]. arterial [FDS92]. arteriosclerosis [DT96]. artificially [STP^{+}98, TOK^{+}94]. ascites [NS96]. ascorbic [WFS^{+}95]. ascorbyl [WFS^{+}95]. Asia [LYLS95]. Asian [RWC93, WR94]. aspects [GBG^{+}93, RWC93, Rou94a]. asperrima [HOF96]. assay [CATR97, DES92, DB96, RPT^{+}93, SBW94, VW96, WBA91]. assays [OGFL92]. assess [BSM^{+}99]. assessed [SHB95, WBS^{+}96]. Assessment [BHP^{+}94, MO95, BB97, IRV^{+}91, SSJ97, WB94]. associated [AM92, BB94, Ber95, BGMC96, BBW98, CDPS90, CCH^{+}94, CLK^{+}97, CRWP95, DMOR97, DDGP^{+}96, FCC92, GHdB90, JRV95, JSB90, JBWT92, KKS^{+}98, LNFT95, LPD94, LGSO97, LPO^{+}98, LF94b, NG99, OMF97, PSB^{+}92, PM91, RHC^{+}93, RV91, RTEC94, SAAH90, SSJ97, SOF^{+}92, SY94, SBD^{+}92, SAMP97, SBJ^{+}98, SDM^{+}98]. Association [BLH^{+}90, DFGG^{+}99, IKU98]. Astacus [SQTF90]. astaxanthin [CGøL^{+}95]. asteroides [Che92b, Che92a, CW93]. Astronotus [MHM95, PGLH96]. asymptomatic [DLP^{+}94, SA95]. Atherinidae [RMRR96]. Atlantic [AFJ^{+}97, ABJE93, BK96, BLP96, BHG91, BB96a, BB96b, BSHJ91, BARH90, BSWR93, Bru92, BR92, BMO^{+}98, CGøL^{+}95, CFHR98, CNH^{+}97, CN99, CSP98, DIHS94, DIB^{+}95, DS95, DLB95, DISS96, DBIS96, DFK93, DF94, DFS94, Dav98, DRP92, EGTH92, FHMG95, FM98, GG99, GHB91, GG97, GJS^{+}97, HEH^{+}90, HNWE94, HCM92, IFMR91, IRV^{+}91, IYBF93, JGOB95, JRV95, JS95, JSB90, JBWT92, JLR93, KSPO99, LTGC99, LMH^{+}97, LSR90, LLP92, LJHE91, MD90, MHB94, MRD98, MSW94, MF95, Møy93, MLB95, MRD^{+}92, NLTN92, NVSB94, NR97, NRV^{+}98, NC99, NHWE95, OCC^{+}95, OGFL92, PSB^{+}92, PMS98, PRL95, PERL96, RT94, Rit97, RRER90, RD93, RR98, RDMW98, SN91, SNØ^{+}93, SRRN94, SNó95, SLJ97, SE99, SPC90, SBD^{+}92, SMBM95, SM96, SSD^{+}97, SBMB95, SAMP97, SHT96, SJGB98, SABòJ99]. Atlantic [SHB98, SCBB93, SSS^{+}99, SD97b, SOP96, SDB99, TRSD98, TT99, Tat91, TL94, TRR96, VD98a, VD98b, VD99, WGSH94, WBS^{+}96, vdIOK96]. Attachment [CH92, BB94, BBW98, OLM^{+}97]. Atypical [LTGC99, OBL^{+}99, Wik90, Aus93, GG97, HMH96, HWGV91, RV91, WDCM92, WDV93]. Auerbach [AS93a]. Augmentation [YKI95]. aurafus [SLL96]. aurata [APAFSP95, Ath92, DFGG^{+}99, FTB90, SbFSAP92]. auratus [CDGM99, DHM97, HN97, JM95, MS91, RHC^{+}93, RQW92, SCE94, WB94, WR98]. aureus [LLLS95]. Australia [CRG^{+}97, CGA93, FCC92, HSP97, KDC95, NC99, RQW92, Rou94a, Vir92]. Australian [AL95, JR98, VPSL93]. australis [HB98, Rou94b, Rou94a]. Autoagglutination [Ber90]. autosomal [Mai92]. auxidis [Jon90]. avirulent [WT92]. axenic [EM93]. ayu [KEY91]. azo [LWHD97].

B [BR92, SB94]. Babesiosoma [CD98]. bacilliform [TIM^{+}96]. Bacteria [Ber95, AS91, BSB91, HLJ^{+}91, IR91, KBE^{+}99, RK91, SSM95]. Bacterial [CSBP^{+}99, LS93, SH92, ABS92, BHB90, FMY94, KWJ90, LE91b, OGFL92, OSG95, OBL^{+}99, SOAK93, SRL^{+}96, SFB^{+}91b, SFB^{+}91a, VCC94, YMM91, Bro91]. bactericidal [BAHL91]. bacterin [MLB95, SYAK95, SEMB^{+}99]. bacterium [AHA92, CGA93, DGK99]. bacuiovirus [LTC96]. baculo [LTBL94]. baculo-like [LTBL94]. baculovirus [KKS^{+}98, LTKC93, LTKC95, TIM^{+}96, VPSL93]. bagrid [MW94]. Baltic [BRB^{+}94, GRR92, HAH92, MB98]. Bangladesh [AR95]. Barbus [MBM96]. Barral [AEL^{+}99]. barramundi [CSM93, GHdB90]. Barrier [HB96]. based [Dav91b, HYHA98, LLGJS94, WPRM94, WBA91]. bass [BR99, CSBCV98, Che92b, DAL90, DAKD98, LGSO97, SBAP90, TAD99, TNB^{+}93]. Bate [TIM^{+}96, WTKC97]. bath [DWV94, SLJ97, SAMP97, WDV93]. bath-administering [SLJ97]. Bay [HOF96, Rou94a, VCC94, BLH^{+}90]. Bayliss [HB98]. Bean [KGW90]. Benedeniella [Tho90]. benzoate [SSS^{+}99]. beta [KGW90]. Betta [LDTC93]. bettencourti [CD98]. between [DEA94, ERSA98, IKU98, MF95, TOK^{+}94]. Bidyanus [FTN97]. Bifidobacterium [IKU^{+}96]. bigemina [DEA94, ED91]. bimaculoides [FHBN91]. binding [BWC92, Col95]. bioassay [KN90b]. biocellatum [MCL97]. Biochemical [DMS99, NBWO91, BAR95]. biochemistry [CMAM96]. bioencapsulation [DHB^{+}95]. Biological [FW96]. biopsy [DHM97]. biosynthesis [DTMD99]. biotype [BAE^{+}91]. biotypes [BEGA93, Dav91b]. birnavirus [MCRF94]. birnaviruses [HSK94, HSK96, MC94]. Bivalvulida [SbFSAP92]. Bizarre [LMTW95]. black [CPLR95, DVB^{+}98, RK91, SA90]. black-tip [SA90]. bladder [VC93]. bleached [SCE94]. Bleeker [CW91, MW94]. Blenniidae [ED91]. Blennius [ED91]. Bloch [AMN^{+}93, BBW98, CSM93, CRWP95, DHM97, GHdB90, KBMGL90, PBP94, RVA98, RM95a, RQW92, Sub93, WR98]. blood [Dav98, FMDC93, JS95, MS90, RHL^{+}94, WGSH94]. bloom [RTEC94]. blot [Ber91, LHS94, OGFL92]. blue [KDL95, RTEC94]. blue-eyed [KDL95]. blue-green [RTEC94]. bodies [RR98]. body [TOK^{+}94]. bonariensis [HLTS93]. bonasus [Tho90]. bond [FAAEC99]. Bonnaterre [FDS92, HB96, HB98]. Bonne [MA99]. Book [Ano90c, Ano90b, Ano91b, Ano91c, Ano92a, Ano92b, Ano93, Ano94a, Ano94b, Ano95c, Ano95a, Ano95b, Ano96, Ano98]. Boone [LTBL94]. booster [PV93]. borchgreviniki [FMDC93]. borreli [SKK90]. Boschma [KDC95]. Bothriocephalus [HN94]. Boulenger [FMDC93]. boundaries [BSBD99]. bovine [Kak98, SOAK93]. Brachydanio [RMT^{+}92]. brama [KT93a]. Branchial [SAMP97, SSJ97, SFB^{+}91b]. Branchiomyces [KLW^{+}98]. branchiophilum [DOF99, OMLF95, OLM^{+}97]. Branchiura [AO94]. bream [APAFSP95, Ath92, DFGG^{+}99, FTB90, KT93a, SbFSAP92]. brevifilum [LES^{+}93]. Brevoortia [SPC90]. British [RR98]. broodfish [BHG91, GHB91]. broodstock [OCC^{+}95]. Brook [AFW94, CC91, LS93, MO95, RM92, SDS90, SDS95, SEMB^{+}99, SA95, SHB98]. Broussonet [BLP96]. brown [BMEC97, CNN^{+}94, DBCA97, FH91, RTEC94, RGA^{+}98, ZPVH99, MMB93]. brown-spotted [CNN^{+}94]. bubalis [Dav90]. Bucephalidae [HOF96]. Buchanan [RMT^{+}92]. buffers [KR95]. bulbocordis [MBM96]. Burchell [AO94, PCKS94, YKI95]. Buron [AEL^{+}99]. Burrishoole [ZPVH99]. butyricum [SYAK95]. bycatch [ZPVH99].

C [Aus93, CR91, CCC93, FWM94, MSGA95, SE99, WVG^{+}98]. C. [PCKS94]. cadmium [RMRR96]. cage [LGSO97, LOF96]. cage-reared [LGSO97]. caged [CNH^{+}97]. calcarifer [GHdB90, CSM93]. Caligidae [HR95, Rit97, Rou94b]. Caligidea [WS90]. Caligus [EMM90, HR95, Rou94b]. callinectes [RRJ^{+}96]. Calyptospora [AMM95]. Calyptosporidae [Jon90]. can [ZH96]. Canada [CC91, LAB92, VCC94]. Canadian [THBL96]. canal [DMOR97]. Capillaria [SMG96]. Capillariidea [SMG96]. capsid [WHC^{+}98]. captive [BLP96, PBP94, RQW92, Rou94b]. captivity [KBMGL90]. capture [VW96, WPRM94]. Carassius [CDGM99, HN97, JM95, MS91, RHC^{+}93, SCE94, SLL96, WB94]. Carbohydrate [Ain93, WGSH94]. carbon [AM92]. Carcharhinus [BGMC96, SA90]. carchariae [BGMC96]. carcinoma [FCS91, HIH^{+}96, MB94, SB94]. Cardiac [RVA98, CD95]. Carnobacterium [JOW^{+}97, TNB^{+}93]. carp [Cai92, CMAM96, CRM^{+}96, CRG^{+}97, CM92, DWV94, FMY94, HN94, HM90, HWGV91, IHKS90, IKU98, KN90a, Kak98, KR95, KKP92, LLLS96, MB94, MS90, NVBK93, RHL^{+}94, RPT^{+}93, SS90, SI91, Söv92, SKK90, VHCCGMOS98, VC93, WDV93, YMM91, YLSW99]. carpio [CMAM96, CRM^{+}96, CRG^{+}97, CM92, FMY94, HN94, HM90, HWGV91, IHKS90, IKU98, KN90a, Kak98, KR95, KKP92, MB94, MS90, RHL^{+}94, RPT^{+}93, SS90, SI91, Söv92, SKK90, VHCCGMOS98, VC93, WDV93, YMM91, YLSW99]. carps [KRK91]. Carrier [MF91, DLP^{+}94, SJMPP91, SDS95, SA95]. carriers [DLB95]. Caryophyllaeus [KT93b]. Caryophyllidae [KT93b]. case [BFPW90, BS98, FOM^{+}95, Góm98, MOBF98]. caseinolytic [SJH^{+}93]. cases [DMS99, PGLH96]. cataracts [WBS^{+}96]. catchment [ZPVH99]. catfish [AM92, Ain93, ARX95, BLSC93, BCD90, CT91, CP96, CA92, DGK99, GWG99, KLW^{+}98, LYD94, MW94, PCKS94, PV93, PS98, PS93, SK97, WBA91, YKI95, ZH96]. cations [BHA95]. Caudal [Mai92]. caught [ZPVH99]. caulleryi [LPO^{+}99]. causative [ARD92, AMN^{+}93, BAR97b, MAR99, SD97a, VC93, ZBM^{+}99]. cause [ASR^{+}93, vdIOK96]. caused [ASR^{+}95, AO94, CTC^{+}94, DGK99, FNTF95, LOF96, Rou94b, YLSW99]. causes [DVB^{+}98]. causing [BRAP99, CNS91, PM91, RGA^{+}98, SRL^{+}96]. CCIV [SS90]. cDNA [DHS94]. CDS [Mai92]. cecropin [KBE^{+}99]. Cell [SLL96, BAR97a, BBCT^{+}94, BARH90, CATR97, CHL99, DLP^{+}94, DPPE97, DSB90, DM90, FLT^{+}98, FCS91, FMDC93, Fre96, GWFE97, JLR93, KSS91, LYGS97, MB94, MS91, RRER90, RSTE93, TW90b]. cell-culture [DLP^{+}94]. cell-free [KSS91]. Cell-mediated [SLL96, TW90b]. cells [CH92, DES92, FAC99, FNTF95, GSC^{+}99, HLHW98, IHKS90, IKU98, JHL^{+}91, LLLS96, LGL^{+}98, LTC96, PR90, PM91, SOP96, WBA91, WHC^{+}98, YIM^{+}97]. Cellular [DES91, HN97, DSA^{+}98, LJHE91, RHL^{+}94]. central [VRK92]. centres [RM95a]. Centrocestus [VHCCGMOS98]. Centropomus [KBMGL90]. Cephalic [OMF97]. cephalopods [ERSA98]. cephalus [Vir92]. Ceratomyxa [IGH92]. cerevisiae [RRER90]. CFLB [SSM95]. Chaetoceros [YA95]. Chaetodipterus [BLP96]. chain [GWG99, KKS^{+}98, LLGJS94, LTKC93, MHM95, SD97a, TIM^{+}96]. chair [KRL96]. challenge [MO95, NR97, SB93, SDB99, WDV93]. challenged [OGFL92, SKA99]. Changes [Møy93, AM92, BF93, Cai92, FWM94, LLLS96, MBM96, TMMA96b, vdIOK96]. Channa [Che92b, CLC90, CRWP95]. Channel [PS98, AM92, Ain93, ARX95, BLSC93, BCD90, CT91, CP96, CA92, DGK99, GWG99, KLW^{+}98, LYD94, PV93, SK97, WBA91, ZH96]. char [RM92]. Characterisation [LSAB99, AE99]. Characteristics [CGA93, DSB90, NBWO91]. Characterization [DHV^{+}99, MA99, VW96, WST96, AH91, AEJ95, BAR97b, BENJ92, BRB^{+}94, CT91, CHL99, FW96, LYGS97, LTC98, RM95a, SSD^{+}97, SRL^{+}96, SJH^{+}93, WLHD97, ZBM^{+}99]. charr [AFW94]. Chelon [MB95b]. chelonae [MHM95]. chemistry [WGSH94]. chemotherapeutic [RRS93]. Chemotherapy [SSM95]. Cherax [JR98, JBRM98, Lan91a, Lan91b, LT92]. Chesapeake [BLH^{+}90]. Chilean [CNS91]. Chinook [BFPW90, SKA99]. Chlamys [HOF96]. chloramine [SSJ97]. chloramine-T [SSJ97]. chloride [PS93]. cholerae [RHC^{+}93]. cholesterol [DB96, EBL^{+}94, LWHD97, RSTE93, SNA^{+}92]. chondrosarcomas [CF91]. chromaffin [YA95]. chronic [BGMC96]. CHSE [WHC^{+}98]. CHSE-214 [WHC^{+}98]. CHV [MS90]. CHV-induced [MS90]. cichild [MCL97]. Cichlasoma [CR91, MCL97, SASG94, SMG96]. cichlasomae [SMG96]. cichlid [CR91, MHM95, SMG96]. cichlids [WGFF98]. Ciliophora [LDCA92, MB95b]. ciprofloxacin [LH90]. circulating [FW98]. cirrusspiralis [RQW92, WR98]. Citrobacter [KKP92]. Clarence [Vir92]. Clarias [AO94, AA91, PCKS94, YKI95]. Clark [LT92]. classification [EHH91]. cleaner [LTGC99, TL94]. clearance [DOF99]. cleavage [WHC^{+}98]. clinical [BLHA92, SBD^{+}92, SSM95]. Clonal [Dav91b]. cloned [DHS94]. Cloning [AH91]. closed [LSL91]. Clostridium [DHB91, SYAK95]. Clubbing [CNH^{+}97]. clupeoid [LTF92]. CNG [CNH^{+}97]. COA [TT99]. coagglutination [TT99]. coagulation [SNØ91, SN91, SNó95]. coagulopathy [SN91, SNØ^{+}93]. coast [AEL^{+}99, FTB90, GRR92]. coated [WFS^{+}95]. coccidian [AMM95, ERSA98]. Coccidiosis [PI91]. coccidium [JK90]. cod [DISS96, ERJ91, KRSE93, MB98, PI91, SSD^{+}97, SJGB98, WHHL95]. cohabitant [EGTH92]. cohabitation [TRSD98]. Coho [BDM91, GCS95, SAK91, TOK^{+}94]. coioides [CHL99]. Coitus [Dav90]. cold [NRV^{+}98]. collagen [ZW98]. collagenolytic [SJH^{+}93]. collected [NHWE95]. collecting [PGLH96]. Colonization [JOW^{+}97, LS93]. colony [WGFF98]. coloured [YLSW99]. columnare [DHV^{+}99, DHTC99]. columnaris [CW91, DHV^{+}99, Wak91]. comb [CD98, LPO^{+}99]. combination [OSG95]. combinations [FWM94]. combined [WVG^{+}98]. Commensal [BSB91]. comments [PGLH96]. commercial [AL95, HOF96]. commercialis [AL95, AAL95]. common [CSBP^{+}99, DM90, FMY94, HMH96, KR95, KKP92, KBMGL90, MS91, Söv92, VHCCGMOS98, VC93]. commonly [Che92a]. communication [DAKD98, ERSA98]. communicationGill [VHCCGMOS98]. communicationMeasurement [DPPE97]. communicationRecovery [ISM^{+}98]. communicationUnusual [Góm98]. Comparative [AKK96, BB97, WT92, dMMMK96, AS93a, BL96]. comparing [LR97]. Comparison [AMW96, FWM94, GHB91, LAB92, LF94b, NR97, BK96, BAR95, CATR97, EMM90, GBH93, HSK96, WFS^{+}95]. competitive [SD93]. complement [FW98]. completely [SKA99]. complex [CT91]. complexed [RSTE93]. components [BAR97a, LSAB99, PERL96, vdIOK96]. composition [Møy93, SKQ^{+}99, WHHL95]. compounds [HHE94]. concavicornis [YA95]. Concentration [MGE90]. concert [SNA^{+}92]. condition [BDM91]. conditioned [BB99]. conditions [BAR97a, Kak98, SJGB98, Wak91]. conductimetry [SGOH94]. confirmation [CNS91]. conformation [FAAEC99]. consumptive [SN91, SNØ^{+}93]. containing [BB97, GMHK93, JHL^{+}91]. contaminated [VFCV99]. contejeani [SQTF90]. continuous [CHL99]. contribution [TW92]. control [BSWR93, RRJ^{+}96, Rod91, RRS93, Tat91]. controlled [SK97, Sub93]. controlling [LSL91]. controls [HK98]. Coomassieagar [Ber90]. copepod [BBW98]. Copepoda [BB94, HR95, Rit97, Rou94b, WS90]. copper [RMT^{+}92, Tho90]. coral [BB94]. corneas [BBW98]. Coronary [DT96, FJS90, FDS92]. correlation [CDPS90, NDBL94]. cortex [MMB93]. corticosteroids [HM90]. cortisol [HM90, YA95]. Coryphaena [HRO90, Lan91c, LMTW95]. Coryphaenidae [LMTW95]. Cottidae [Dav90]. Couch [RRJ^{+}96]. counteracted [SNA^{+}92]. County [ZPVH97, ZPVH99]. course [WFS^{+}95, WB94]. covertly [TRSD98]. cownose [Tho90]. crab [KDC95]. crabs [RVSK99]. cracked [AA91]. cracked-skull [AA91]. Crassostrea [ABC^{+}98]. crassus [HYHA98, HAH92]. crayfish [JR98, JBRM98, SQTF90]. creatine [WST96]. Crenicichla [AMM95]. creosote [VFCV99]. creosote-contaminated [VFCV99]. Cross [LYD94, LSR90, BL96, FH91, GG97, HIMK93, ERSA98]. Cross-protection [LSR90, BL96]. cross-reacting [FH91]. cross-reactive [HIMK93]. Cross-reactivity [ERSA98]. Crustacea [AO94, DVB^{+}98]. crustacean [RMR95]. cryopreserved [IGH92]. cryosections [KSPO99]. Cryptobia [AFW94, FW96, FW98, KBW95, SBW94, TW90a, TW90b, TW92, VW96, WT92, ZPVH99, ZFW97, ZW98]. cryptobiosis [AFW94, LW97, SBW94]. Cryptocaryon [MMB93, MB95b]. Cryptogonimidae [SASG94]. Ctenolabrus [BBS96, CA95, GSS98]. Ctenopharyngodon [Cai92]. Culaea [LS93]. Culaeatrema [LS93]. Cultivation [WB94, BHG91, SSD^{+}97]. culturability [PRCC96]. culture [APAFSP95, DTMD99, DLP^{+}94, GWFE97, LR97, LSL91, PSM^{+}96, SK90, SJGB98]. cultured [ASN90, Ber91, BFCH91, CSBP^{+}99, DDGP^{+}96, DAKD98, FNTF95, GSC^{+}99, JRV95, JBRM98, KEY91, LGL^{+}98, LOF96, MOBF98, PSL^{+}93, RRJ^{+}96, SAAH90, SAK91, SBAP90, SbFSAP92, TWC99, TLMM94, UUNY91, WTKC97, YIM^{+}97]. cultures [BBCT^{+}94, Fre96]. Cura [NG99]. Cura-d [NG99]. Current [RRS93]. cutaneous [Vir92]. cuticle [JR98]. cuticular [JBRM98]. Cuvier [APR99, Cai92, DAKD98, HLTS93, MHM95, PGLH96, WS90]. Cyanobacterial [RTEC94, BMEC97]. cycle [CSBCV98, GCS95, MMB93, MSWB98, MHM95]. cyprinacea [WS90]. cyprini [SMSA91, SSSH92, SMHS93]. Cyprinus [CMAM96, CRM^{+}96, CRG^{+}97, CM92, FMY94, HN94, HM90, HWGV91, IHKS90, IKU98, KN90a, Kak98, KR95, KKP92, MB94, MS90, RHL^{+}94, RPT^{+}93, SS90, SI91, Söv92, SKK90, VHCCGMOS98, VC93, WDV93, YMM91, YLSW99]. cysts [KRL96]. cytolysin [LE91a]. cytometry [SJMPP91]. cytopathology [LF94b]. Cytophaga [CSM93, SSM95]. Cytophaga/Flexibacter [SSM95]. Cytophaga/Flexibacter-like [SSM95]. Cytophage [AHA92]. Cytophage-like [AHA92]. cytoskeleton [PS98]. Cytotoxin [YIA92].

d [NG99, DIB^{+}95, DS95, DISS96, DBIS96, DMH^{+}98, TMMA96b]. dab [BF93, DM90]. damage [AO94, BB94, BMEC97]. damsela [ZBM^{+}99, FNTF95]. damselfish [NG99]. danilewskyi [BB97, BB99, WB94]. data [AMM95, CA95]. DCH [HK98]. dead [AS91]. Decapoda [HHS93, Lan91a, Lan91b]. defence [ABC^{+}98, AMNH99, CA92, DIB97]. deficiency [CR91, FWM94, ISM^{+}98, TMMA96a]. deficient [CCC93, PMS98]. defined [DMTM95, ERJ91]. deformities [VD98a]. deformity [Mai92]. Delivery [DHB^{+}95]. demonstrate [MB94]. Demonstration [FHMG95, HCW93]. Dempsey [MCL97]. dentex [CSBP^{+}99, CSBP^{+}99, AMN^{+}93]. depends [SOS94]. depletion [MHB94]. deposits [JR98]. derangement [ISM^{+}98]. derivatives [DIHS94]. derived [CHL99, IKU^{+}96]. derjavini [LB98]. Dermal [CDGM99]. Description [BDM91, CC91, Lan91a, MBM96]. Desseria [CD98]. detectable [Dav91a]. detecting [BR99]. Detection [GWG99, GBH93, JHL^{+}91, KN90b, LTKC95, SJMPP91, BB96a, BB96b, DLP^{+}94, HEHR91, KKS^{+}98, LWHD97, LLGJS94, LTL96, MGE90, PI93, RPT^{+}93, SAK91, SSSH92, SBW94, TT99, VW96, VPSL93, WBA91]. deteriorating [Kak98]. determinants [AH91]. determination [YDOW97]. determined [SDM^{+}98]. develop [DF94]. developing [WBS^{+}96]. Development [ARD92, AL95, BK96, BBCT^{+}94, F.92, HAC^{+}92, KKS^{+}98, LTKC93, LTL96, WLHD97, CSP98, FTB90, GG99, MD90, RMT^{+}92, SI91, SFB^{+}91a, TGK94]. developmental [DM90, IGH92]. Dexamethasone [LB98]. diacetate [YDOW97]. diagnosis [AAL95, DHS94, PMS98]. diagnostic [AL95, ASRT^{+}94, FH91, PMS98]. diatom [YA95]. Dicentrarchus [BR99, CSBCV98, DAL90, DAKD98, LGSO97, SBAP90, TAD99]. dichlorvos [JSW92]. diclina [RM92]. diet [TMMA96b]. Dietary [TW90a, WGSH94, WHHL95]. diets [CGøL^{+}95]. Differences [HWGV91, WDV93]. different [APAFSP95, AS93b, CRWP95, DT93, GHB91, LSR90, MSWB98, MF91, NE91, NR97, PERL96, RFN93, SBN^{+}91, WLHD97]. differential [F.92]. Digenea [LS93, SASG94, SI91, WCS90]. Digenean [HOF96]. digenetic [DMOR97]. dimorpha [AEL^{+}99]. Diplectanidae [CSBCV98]. Diplectanum [CSBCV98]. Diplostomum [WCS90]. Diplozoon [KT93a]. Direct [DPPE96, MHM95]. Discovery [AEL^{+}99]. discrimination [BEGA93]. Disease [WDV93, ARD92, ASRT^{+}94, AMN^{+}93, ASR^{+}93, AA91, BHB90, BL96, BDM91, BRAP99, BFCH91, Cai92, CLK^{+}97, CFHR98, CNN^{+}94, CN99, CNS91, DM90, DVB^{+}98, DFGG^{+}99, FOM^{+}95, FM98, FMDC93, FCC92, GG99, GMHK93, KN90a, LPO^{+}98, MOBF98, MRD98, MAR99, MRD^{+}92, NHWE95, OGFL92, OBL^{+}99, PCKS94, RRER90, RDMW98, SAAH90, SD97a, SBD^{+}92, SFB^{+}91b, SFB^{+}91a, SBJ^{+}98, TLMM94, Vir92, WGSH94, WVG^{+}98, WTKC97, CNN^{+}94]. diseased [AHTY92, AGS^{+}92, BAE^{+}91, CPLR95, LYLS95, MA99, NG99]. Diseases [APR99, Bro91, ASR^{+}95, SRL^{+}96]. disinfectants [Che92a]. display [BAHL91]. Disseminated [SN91]. dissolution [SDM^{+}98]. Dissonidae [BB94]. Dissonus [BB94]. distal [BK96, vdIOK96]. distinct [TAD99]. Distribution [DM90, DSA^{+}98, PMD90]. disulphide [FAAEC99]. diversity [BHP^{+}94, THBL96]. DNA [CMMM95, FAAE^{+}99, GWG99, LTKC93, MEJ^{+}94, MAR99, TIM^{+}96]. Do [SNØ91]. does [Dav91a]. Dogiel [Jon90]. Dolops [AO94]. dolphin [HRO90]. dorsal [BMEC97, TRR96]. dose [MHB94, SE99]. doses [GACHF91]. dot [Ber91]. double [GBH93]. double-sandwich [GBH93]. doughboy [HOF96]. drum [CCC93, CD95]. Duboscquodinida [HHS93]. ducts [PGLH96]. due [FWM94, ISM^{+}98, Lan91b, Lan91c, LT92, LDTC93, NRV^{+}98, PI91, RVA98, VHCCGMOS98]. during [AS91, AS92a, EGTH92, GCS95, IRM^{+}96, Møy93, NHWE95, RMP^{+}94, SFB^{+}91a, SDM^{+}98, VC93]. dye [LWHD97]. Dykova [MEMSB99]. dynamics [HR95, WS90].

early [Ber95, SNó95, SDM^{+}98]. early-maturing [SNó95]. East [TLMM94, LYLS95]. eastern [FCC92, LAB92, THBL96]. eaused [UUNY91]. eberthi [ERSA98]. ecological [Rou94a]. ectodermal [TIM^{+}96]. ectoparasites [RMR95]. ectoparasitic [FHBN91]. Ecuador [JBRM98]. Edwardsiella [Ain93, ARX95, GMHK93, PV93, SK97]. eel [AMNH99, BAE^{+}91, BRW92, DMTM95, DTMD99, EA91, GMHK93, HAH92, HCW93, HLTK96, LPO^{+}99, RM95a, SB93, ZPVH97]. eels [CD98, HLTK97a, HLTK97b, LKM99]. Effect [KR95, KWJ90, Söv92, SBD99, Wak91, BAR97a, DS95, DBIS96, DMH^{+}98, FMY94, HCM92, JMK^{+}95, MS90, RMT^{+}92, TNAF99, WFS^{+}95]. effective [ASR^{+}95]. effectiveness [GHB91]. Effects [ABC^{+}98, AMNH99, DAL90, DOF99, KBE^{+}99, MLB95, PFO^{+}98, Sub93, CMAM96, FW98, GACHF91, HM90, KGW90, KDC95, MSWB98, PS93, RMRR96, SNØ^{+}93, SDS90, SMG96, Tho90]. efficacies [SSM95]. efficacious [SKA99]. Efficacy [YLSW99, DHB^{+}95, IRV^{+}91, LW97, NVSB94, NR97, NRV^{+}98, SLJ97, SSS^{+}99]. effluent [HAH92, MB97, SCE94]. EGCs [SOP96]. egg [GMHK93]. eggs [BSB91, BHT92, MSGA95, NG99, RM92, RHL^{+}94, SMG96]. EHNV [WPRM94]. Eigenmann [KDL95, RMRR96]. Eimeria [Dav90]. elaboration [YIA92]. elasmobranchs [FDS92]. electroimmunotransfer [OGFL92]. electron [DMOR97, JK90, MSW94, SHB95, TWC99, TRR96, WTKC97]. electrophoresis [BHP^{+}94, LSAB99, WST96]. Electrophoretic [BSH^{+}90]. elements [ISM^{+}98]. Elevated [SB94]. Elevation [YA95]. elicit [FAAEC99]. elimination [Fre96]. ELISA [BR99, DLP^{+}94, GBH93, HEHR91, RPT^{+}93, SBW94, WBA91, WPRM94]. ELISA-based [WBA91]. ELISPOT [DES92]. elongata [BBW98]. elongatus [EMM90]. emamectin [SSS^{+}99]. embryonic [HLHW98]. encoding [Ner96]. endogenous [Dav90]. endothelial [IKU98]. energy [CCD95]. England [AS91, AS92a, Aus93, DL99]. enhanced [SPC90]. Enhancement [IKU^{+}96, NK93, RRER90, SOAK93, SYAK95]. enrofloxacin [SCBB93]. enteric [AS92b, DGK99, SK97]. entering [SNØ91]. enteritis [BK96]. Enterobacter [HRO90]. enterococcosis [AMW96]. Enterococcus [AKK96, CGA93, IKU^{+}96, NDQT95]. Enterococcus-like [CGA93]. entry [DRL94, LGL^{+}98, OJW^{+}96]. environment [Bru92, EA94, MB97, PI93, VFCV99]. Environmental [BLSC93, Vir92, CSP98, Kak98, Wak91]. Enzyme [RM92, RPT^{+}93, BHP^{+}94, CATR97, LTL96, PRL95, SBW94, VW96].

Enzyme-linked [RPT^{+}93, CATR97, SBW94, VW96]. enzymes [KGW90]. Eoacanthocephala [Tar90]. eosinophilic [SOP96]. epaulette [HB96, HB98]. eperlanus [AAHY90]. epidemicus [HR95, Rou94b]. epidemiological [F.92]. Epidemiology [WPRM94]. epidermis [PR90, SHB98]. Epinephelus [CLK^{+}97, CHL99, CNN^{+}94]. epithelial [FLT^{+}98, LYGS97, LLLS96]. Epitheliocystis [FTN97, LMBS92, NC99]. epithelium [JBRM98]. Epitope [SA94]. Epizootic [EHH91, HEHR91, AR95, CCH^{+}94, CRWP95, HAC^{+}92, LF94b, LR97, RWC93, Sub93, WPRM94, WR94, WRC95, YIA92]. Epizootics [DGK99]. erosion [SCE94]. Erratum [Ano92c, Ano99]. erythrocytes [LE91b]. erythrocytic [LNFT95, RR98, TOK^{+}94]. erythrodermatitis [DWV94, Söv92]. esmarkii [PM91]. Esox [BRB^{+}94, Wik90]. Establishment [CHL99]. ester [SE99]. esterolytically [LWHD97]. Estimation [DGR97]. Estonian [BRB^{+}94]. estuarine [FCC92, VCC94]. etropli [RVA98]. Etroplus [RVA98]. Eucestoda [KT93b]. Euphrasen [Dav90]. European [AAHY90, AMNH99, BAE^{+}91, BRW92, EA91, HAH92, ZPVH97]. EUS [CCH^{+}94, CRWP95, LF94b, LR97, RWC93, Sub93, WR94]. evaluate [NR97]. Evaluating [BSBD99]. Evaluation [BEGA93, GSC^{+}99, GG97, NRV^{+}98, PSM^{+}96, TT99, NVSB94]. Evidence [HSK94, SNó95, SPC90, SD93, SN91, WTKC97]. Examination [PMS98]. excised [DT93]. exclusion [SD93]. exhibit [DFK93]. exhibiting [KRL96]. exogenous [NHOJ93, SNØ^{+}93]. exoprotease [NDBL94]. exoproteases [WLHD97]. Exotoxin [SNØ^{+}93]. Exotoxin-induced [SNØ^{+}93]. experiment [EGTH92]. Experimental [AFJ^{+}97, BHT92, BB96a, BB96b, BRW92, Gat91, IGH92, LT92, MEMSB99, RVSK99, SBMB95, SHT96, BHB90, Bri95, CP96, DGR97, DMH^{+}98, GSS98, MF95, SK90, SHB98, SDB99]. experimentally [CRWP95, DFK93, DFS94, OGFL92, OMLF95, SKA99, SR99, SKK90, TNB^{+}93]. Experiments [DMTM95, HR95, WR98, ERJ91, SS90]. exposed [CRM^{+}96, CRG^{+}97, DTMD99, SCE94]. exposure [BSBD99, DWV94, SK97, TRSD98]. express [ZH96]. expressed [BWC92]. expression [BSBD99]. external [DT93]. Extracellular [CAR97, EBL^{+}94, LF94a, AEJ95, AE99, BAR97a, Bro90, CATR97, CYA^{+}98, CCD95, Ell91, FNTF95, HCM92, KON90, LE91a, LJHE91, MA99, NE91, SNØ91, SJH^{+}93, Tat91]. extract [RMR95]. extracted [EMM90]. extracts [JOW^{+}97]. extrasporogonic [MSW94]. exudates [AOES99]. eyed [KDL95].

faber [BLP96]. Fabricius [CPLR95, LTKC95, RRJ^{+}96, RK91, TIM^{+}96, WTKC97]. facilities [MB97, SRL^{+}96]. factor [SNØ91]. Factors [AR95, BLSC93, CSP98, JGOB95, VD98a, VD98b, VD99, Vir92]. faecal [JOW^{+}97]. faeces [OJW^{+}98]. fails [SDS95]. fairy [DVB^{+}98]. Fallowing [BSWR93]. family [HHS93]. farm [DGK99, KLW^{+}98, TL94]. farm-raised [DGK99]. farm-reared [KLW^{+}98]. farmed [AR95, ALVS94, BDM91, BRAP99, Bru92, BS98, CSM93, CFHR98, HMH96, LTGC99, LKM99, MRD98, NDQT95, NC99, PCKS94, Rit97, RR98, RDMW98, SN91, SBD^{+}92, SMBM95, TRR96, WPRM94]. farms [BSWR93, MRD^{+}92, RMP^{+}94, VRK92]. fasciatus [YI90]. Fast [MEJ^{+}94]. fasted [BK96]. fate [CM92]. fatty [PBP94, WHHL95]. Faust [HN97]. FDA [YDOW97]. features [Góm98]. fecundity [WR98]. fed [CGøL^{+}95, STP^{+}98, TMMA96b, WHHL95]. feeding [Söv92]. feral [CRG^{+}97, SY94, SCE94]. fibrinolysis [SNó95]. fibroblast [BB99]. fibroblast-conditioned [BB99]. fibroma [CDGM99]. fibromyxoma [MB95a]. Field [CN99, LLP92, LSAB99, MC94, NVSB94]. fighting [LDTC93]. filament [LKM99]. fimbria [BHB90]. Fin [SCE94, TRR96]. findings [WPRM94]. fingerlings [KKP92]. Finland [RV91, VRK92]. First [BAE^{+}91, GCS95]. Fish [Bro91, GBG^{+}93, AR95, AMW96, AHA92, AAAR98, AHTY92, AB90, AS91, ABS92, AS92a, AMM95, BSH^{+}90, BK96, BAR97a, BAR97b, BLH^{+}90, BTL^{+}92, BSM^{+}99, BB97, BA91, BQLC97, BSHJ91, BFCH91, CRG^{+}97, CA95, CATR97, CAR97, CM92, CMMM95, DGR97, DIB97, DSA^{+}98, DLP^{+}94, DSB90, DMS99, DTS^{+}92, EA94, FLT^{+}98, FAC99, FAAE^{+}99, FCC92, Fre96, GRR92, Góm98, GSC^{+}99, GWFE97, HR95, HRO90, HLHW98, HSK94, HSK96, IR91, JHL^{+}91, JOW^{+}97, KT93b, KWJ90, KBE^{+}99, KN90b, LTGC99, LPD94, LYGS97, LLLS95, LYLS95, LMBS92, LSL91, LDCA92, LDTC93, LGL^{+}98, LLN94, MO95, MCRF94, NE91, NBWO91, OJW^{+}96, OSG95, RMRR96, SNØ91, SMHS93, SD93, SSM95, SGOH94, TOK^{+}94, TL94, VCC94, VRK92, Wak91, WR98, WR92, WR94, WRC95, WS90, YIA92, YIM^{+}97]. fish-pathogenic [AHA92, GRR92, IR91, SGOH94]. fishes [DEA94, RWC93]. Fishing [MB98]. five [FDS92]. fixing [FW98]. flagellate [UUNY91]. Flavobacterium [DHV^{+}99, DHTC99, DOF99, OMLF95, OLM^{+}97]. flesh [Lan91c]. Flexibacter [CW91, DHV^{+}99, HSP97, LOF96, OMF97, PSM^{+}96, Wak91]. Flexibacter-like [SSM95]. Florfenicol [IRV^{+}91, NVSB94, NRV^{+}98]. flounder [UUNY91]. flow [SJMPP91]. fluke [RHL^{+}94]. fluorescein [YDOW97]. fluorescens [BQLC97, CSP98, FH91]. fluorescent [FAAE^{+}99, OCC^{+}95, SD93]. fluoroquinolone [SCBB93]. Fluoroquinolones [BAHL91]. focusing [WST96]. foetoprotein [SB94]. foliaceus [RMR95]. following [BMEC97, DWV94, PRL95, PERL96, SK97, SDB99]. fontinalis [MO95, AFW94, CC91, RM92, SDS90, SDS95, SEMB^{+}99, SA95, SHB98]. foreign [ZH96]. form [MB94, OBL^{+}99]. formalin [DOF99, NLTN92, SAMP97]. formalin-killed [DOF99]. formation [CCD95, SNA^{+}92]. Formosa [Che92b]. formosanus [VHCCGMOS98]. forms [DTMD99]. Forskål [KDC95]. Forsskål [CNN^{+}94, SAAH90]. fossilis [PS93]. Fouquet [BCD90, HM90, LYD94]. four [Dav91a, ERJ91]. Fowler [CLC90]. fragment [BWC92]. Franca [CD98]. France [BAR95]. free [HCM92, KSS91, NVBK93, TMMA96b, WB94]. frequencies [BAHL91]. fresh [LW97, NBWO91]. freshwater [AFJ^{+}97, ASN90, GCS95, LSL91, LTC98, NBWO91, RMRR96, RKMF97, TWC99, WRC95]. freshwater-raised [AFJ^{+}97]. Freundii [KKP92]. Fry [SDS90, AS91, AS92a, MSGA95]. fumagillin [Söv92, YLSW99, HK98]. fumagillin-DCH [HK98]. fumatus [HOF96]. Fundulus [VFCV99]. fungal [SPC90, WR94, WRC95]. fungi [LR97]. fungicides [WR92]. fungus [CRWP95, SGOH94]. Further [FM98, BENJ92, BFCH91]. furunculosis [Bri95, GG97, IRV^{+}91, IRM^{+}96, JGOB95, LE91a, LLP92, MF91, NVSB94, NR97, NRV^{+}98, PRL95, PERL96, SGNG93, SH92, SD93, SCBB93, Tat91, TGK94]. fuscogutatus [CLK^{+}97]. fusion [FAC99].

G [FAAEC99]. Gadus [DISS96, ERJ91, KRSE93, MB98, SSD^{+}97, SJGB98, WHHL95]. Gaimard [SA90]. gairdneri [NLS^{+}90, PR90, SK90]. galactosidase [BWC92]. Galicia [RMP^{+}94]. gariepinus [AO94, PCKS94, YKI95]. Gasc [AEL^{+}99]. gasterosteid [VCC94]. Gasterosteus [JK90]. gavage [CMAM96, CRM^{+}96]. GCAT [Ner96, RSTE93]. GCAT-LPS [RSTE93]. gear [MB98]. gear-induced [MB98]. gel [LSAB99]. gels [LWHD97]. gene [MHM95, Ner96, WDCM92, ZH96]. generated [BSBD99, KSS91]. genes [ZH96]. Genetic [MMJ^{+}97, BHP^{+}94, LTC98, WDV93]. genogroup [LHS94]. Genogrouping [HSK96]. genome [HSK96, SSSH92, SMHS93]. genomes [TAD99]. genus [MSW94]. German [GRR92]. GF [CHL99]. GF-1 [CHL99]. giant [ASN90, PM91, TWC99, WTKC97]. giganteum [CA95]. Gilchrist [Lan91c, LTF92]. Gill [LKM99, SK90, ASR^{+}93, BRW92, CNH^{+}97, CN99, DHV^{+}99, DHTC99, FM98, FMDC93, KR95, LPO^{+}98, OBL^{+}99, SS90, SB93, SFB^{+}91b, SFB^{+}91a, SBJ^{+}98, VC93]. gills [BB94, BBCT^{+}94, DOF99, OLM^{+}97, PFO^{+}98, RMRR96, VCC94]. gilt [FTB90, SbFSAP92]. gilt-head [FTB90]. gilthead [APAFSP95]. glass [HLTK97b]. glauca [FDS92]. globiceps [HYHA98]. Glucan [CA92, DIB^{+}95, DS95, DISS96, DBIS96, DMH^{+}98, JLR93, RRER90, SD97b]. Glugea [LPO^{+}99]. Glycerophospholipid [RSTE93, EBL^{+}94, LWHD97, SNA^{+}92]. glycerophospholipids [DB96]. Glycogen [CCD95]. glycoprotein [ZFW97]. glycyrrhizin [JMK^{+}95]. Gmelin [ABC^{+}98, BP91]. GNNV [CHL99]. Goitre [WGFF98]. goldfish [CDGM99, HN97, JM95, MS91, RHC^{+}93, SCE94, SLL96, WB94]. goldsinny [BBS96, GSS98]. gonads [MW94, RM95a]. good [LYLS95]. Goode [KGW90]. Goussia [JK90, Jon90, PI91]. Grant [BBW98]. granule [SOP96]. granulomas [CTC^{+}94]. granulosum [DMTM95, DTMD99, ZPVH97]. grass [Cai92]. gray [SW90]. grayling [DL99]. Great [HB96]. greater [LPO^{+}99]. Greece [Ath92, APR99, BAR95]. green [ACH93, FAAE^{+}99, GACHF91, MSGA95, RTEC94]. Greenland [BBW98]. griseus [SW90]. Gross [TRR96]. Grouper [CNN^{+}94, CLK^{+}97, CHL99]. groups [WBS^{+}96]. grow [Aus93]. grown [GWFE97]. Growth [PLLT94, SGOH94, BAR97a, Ber95, Ber90, BB97, CCD95, DMTM95, GHB91, NHOJ93, SI91, SDM^{+}98, TOK^{+}94, BB99]. gulf [KGW90]. Gunther [PCKS94, CR91, CW91, HR95, MBM96, Rou94b, Rou94a, SASG94, SMG96]. gut [DES91]. Gyrodactylus [CMMM95, DL99, LB98, SHB98].

habitats [AS93b]. haem [HHE94]. haemagglutination [DSB90]. haemangiopericytoma [MS91]. Haematological [HAH92, SKK90, TMMA96b]. haematology [WHHL95]. haematopoietic [CFS93, DRL94, EHH91, HEHR91, JM95, PSL^{+}93, Söv92, WPRM94]. haemoflagellate [KBW95, SBW94, ZFW97, ZW98]. Haemogregarina [CD98, DEA94, ED91]. haemogregarines [MD90]. haemolysin [AH91]. haemolysins [ABJE93]. haemolysis [NDBL94]. haemolytic [RSTE93]. haemopoietic [PSB^{+}92]. Haemorrhagic [ALVS94, BMO^{+}98, Cai92, Col95, DT93, SBB^{+}93, YLSW99]. Hafnia [RGA^{+}98]. halibut [BHT92, Ber95, BB96a, BB96b, JRV95, SLJ97, SE99, SD97b]. Haliotis [DHB91]. Hamilton [CHL99, RMT^{+}92]. Hamilton-Buchanan [RMT^{+}92]. harboured [MF91]. harmful [YA95]. harveyi [AAAR98, KBMGL90, MA99]. hatcheries [WR92]. hatchery [BR99, CLK^{+}97, MGE90, YI90]. hatchery-reared [CLK^{+}97, YI90]. head [FTB90, IKU98, JS95, JLR93, LTBL94, LTL96, SbFSAP92, SM96, SSD^{+}97, SJGB98]. Health [SRL^{+}96, CRG^{+}97, MCRF94]. healthy [AHTY92]. heart [DISS96, DT96, MBM96]. Heckel [AMM95]. held [Bro91, Kak98, KBMGL90, SE99]. Helostoma [WS90]. Hematodinium [HHS93]. Hemiscyllium [HB96, HB98]. hemisuccinate [HM90]. Henneguy [TSL^{+}94]. Henneguya [Rou94a]. Hepatic [PBP94, STP^{+}98, KGW90]. Hepatoblastomas [VFCV99]. hepatocellular [SB94]. hepatocytes [Ain93]. hepatopancreas [JR98]. hepatotoxicity [AM92]. hepatotoxin [BMEC97]. Herpesviral [JM95]. Herpesvirus [LKM99, SMSA91, SSSH92, SMHS93, ZH96]. herpesviruses [HIMK93]. heteroclitus [VFCV99]. Heterogeneity [NE91]. Heteropneustes [PS93]. Heterosporis [LDTC93]. Hewitt [Rou94b]. High [UUNY91, DHTC99]. Hippoglossus [BHT92, Ber95, BB96a, BB96b, JRV95, SLJ97, SE99, SD97b]. hippurus [HRO90, Lan91c, LMTW95]. histochemical [DBCA97, MB94, RM95a, SSJ97]. histochemistry [IHKS90]. histologic [VFCV99]. Histological [TWC99, RM95a, TRR96]. histology [AO94]. Histopathological [BF93, FNTF95, RMRR96, Rou94a, TMMA96b, MBM96, TLMM94]. Histopathology [BSM^{+}99, CP96, CR91, CRWP95, DMOR97, PS93, Rou94b, SAAH90, SASG94, TNB^{+}93, BL96, BP91, CRM^{+}96, JSB90, JBWT92, RVSK99]. hoferi [SGOH94, SHB95]. Hoffman [HN97]. homestead [AA91]. homogenates [RPT^{+}93]. horizontal [AFJ^{+}97]. hormone [KRL96].

Hoshina [HLTK96, HLTK97b]. Host [KT93a, KT93b, Tar90, ABC^{+}98, AFW94, AEL^{+}99, FHBN91, LTF92, LMBS92, RVSK99, Rit97]. Host-parasite [KT93a, KT93b, Tar90]. hosts [LTF92, MEMSB99]. hovorkai [YLSW99]. human [DMS99]. Humoral [HLTK96, LES^{+}93, LJHE91, ERJ91, GJS^{+}97, HM90, TW90a]. hybrid [BWC92, FCS91, PCKS94]. hybridization [BB96a, BB96b, LTKC95, LHS94, MAR99, SSSH92, SMHS93, SDM^{+}98]. hybrids [DT93]. hydrocortisone [HM90]. hydrogen [DOF99, MSWB98]. hydrophila [FLT^{+}98, AHTY92, AH91, CCH^{+}94, CT91, DVB^{+}98, DAKD98, Gat91, KRK91, LYGS97, LLLS95, LYLS95, LLLS96, LGL^{+}98, NE91, NBWO91, RFN93, YIA92]. hydrophobicity [DPPE97]. Hyperoplus [LPO^{+}99]. hypertrophic [YA95]. hypoxia [ABC^{+}98].

Ichthyobodo [BLP96, Bru92, TSL^{+}94, UUNY91]. Ichthyobodo-like [BLP96]. Ichthyophoniasis [Ath92]. Ichthyophonus [SGOH94, SHB95]. Ichthyophthiriasis [CM92, LSL91]. Ichthyophthiriidae [MMB93]. ichthyophthiriosis [WFS^{+}95]. Ichthyophthirius [BCD90, EM93, HM90, LYD94, LDCA92, SLL96]. Ichthyosporidium [CA95]. Ictafurus [CP96]. ictaluri [Ain93, ARX95, PV93, SK97]. Ictalurid [ZH96]. Ictalurus [AM92, Ain93, ARX95, BLSC93, BCD90, CT91, CA92, DGK99, GWG99, KLW^{+}98, LYD94, PV93, SK97, WBA91]. idellus [Cai92]. Identification [Ber91, CMMM95, DHB91, Fre96, LTC98, LHS94, MHM95, BEGA93, CT91, GCO91, HAC^{+}92, MEJ^{+}94, MEMSB99, NG99, WDV93]. identified [BLHA92]. IgY [GMHK93]. ihlei [KDC95]. imaging [PMS98]. immersion [CMAM96, CRM^{+}96, FAAE^{+}99, KEY91, PV93, SDS90]. immobilization [LYD94]. Immune [ARX95, AHA92, CA92, CYA^{+}98, ERJ91, HLTK96, LES^{+}93, LJHE91, SLL96, TNAF99, TW90a, TW90b]. immunity [CGøL^{+}95, HM90, LYD94, SK97, WGSH94, WHHL95, WVG^{+}98]. Immunization [BCD90, SDS95, AMW96, LSR90, SDS90]. immunized [CM92, HEH^{+}90, RMR95]. immuno [MB94]. immuno-histochemical [MB94]. immunoassay [Ber91, LTL96]. immunoblotting [CATR97]. Immunochemical [KRSE93, BSH^{+}90]. Immunocytochemical [SOP96, GBG^{+}93]. immunodepression [TW92]. Immunodiagnosis [CCH^{+}94]. immunoelectron [HEHR91]. immunoelectron-microscopy [HEHR91]. Immunogenicity [BLW95, KON90]. immunoglobulins [HEH^{+}90]. Immunohistochemical [BRB^{+}94, ASRT^{+}94, DRL94, JHL^{+}91, VFCV99]. immunohistochemistry [BB96a, BB96b, HEHR91, SDM^{+}98]. Immunohistological [BARH90]. Immunological [ABJE93, KRK91]. immunomodulating [SD97b]. immunomodulatory [DIHS94, DIB^{+}95, DS95, DISS96, DBIS96]. immunoreactivity [SB94]. immunosorbent [CATR97, RPT^{+}93, SBW94, VW96]. immunostimulants [YKI95]. Immunosuppression [HM90]. Impact [KBW95]. impaired [DFK93]. implications [CRG^{+}97]. important [LYGS97]. impression [VPSL93].

Improved [WR94]. Incidence [RMP^{+}94]. includes [BWC92]. inclusion [RR98, TOK^{+}94]. inconstans [LS93]. increases [Col95, LB98]. incubating [BSB91]. Incubation [BHG91, PFO^{+}98]. India [RVSK99]. Indian [KRK91, RRJ^{+}96]. indicator [LYLS95]. indirect [Ber91, OCC^{+}95]. individual [WDV93]. Indonesia [TLMM94]. induce [SNØ91, ZH96]. Induced [FAC99, SBJ^{+}98, BK96, MB98, SNØ^{+}93, SB94, SEMB^{+}99, TOK^{+}94, YMM91, MS90]. inducible [SD93]. Induction [HLHW98, TRSD98, Bri95, SHT96]. inermis [RHL^{+}94, SI91]. infected [AS92b, CRWP95, DFK93, DF94, DFS94, FW98, FAC99, GWG99, HK98, KT93a, LKM99, LLLS96, LTC96, MB92, OMLF95, SMHS93, SASG94, SR99, SKK90, TRSD98, Tar90, TW90a, TW92, TNB^{+}93, VW96]. infecting [CA95, PFO^{+}98]. Infection [EC91, HMH96, HB98, HHS93, LTBL94, RQW92, AFJ^{+}97, BHT92, BB96a, BB96b, BRAP99, BP91, CP96, CW91, DGR97, DHB91, DFGG^{+}99, DAKD98, DRP92, EGTH92, FMY94, Gat91, GSS98, HR95, HLTS93, HLTK97b, IGH92, IKU^{+}96, KKP92, KLW^{+}98, LTGC99, LNFT95, LDTC93, LTKC95, MB95b, MF95, MF91, Møy93, NDQT95, RVSK99, RRJ^{+}96, RHC^{+}93, RKMF97, SOAK93, SbFSAP92, SI91, TL94, TWC99, WB94, WPRM94, YMM91]. infections [ABC^{+}98, AS93b, BHB90, DMH^{+}98, DHS94, HOF96, LB98, Rou94a, SB93, SSM95, Sub93, ZPVH97, ZPVH99]. Infectious [CND90, EGTH92, GSS98, MB97, PSL^{+}93, SOS94, SBD^{+}92, BB96a, BB96b, CFS93, CSP98, DFK93, DF94, DFS94, DLP^{+}94, DHS94, DRL94, HEH^{+}90, HLHW98, HNWE94, HCW93, JGOB95, JS95, LAB92, LLGJS94, MO95, MC94, MF95, NHWE95, NKKH97, RM95b, SJMPP91, SOF^{+}92, SDS90, SDS95, SA94, SY94, SM96, SHT96, TRSD98, TT99, THBL96, WHC^{+}98]. Infectivity [DFS94, SS90, BBS96, NK93, Wak91]. infestation [HAH92]. infestations [RRS93, SSS^{+}99]. infiltration [PBP94]. inflammation [NVBK93, VC93]. inflammatory [AOES99]. inflicting [AO94]. Influence [CSBCV98, DHTC99, WGSH94, WVG^{+}98, BR92, SHB98]. Influences [SKQ^{+}99]. influencing [Vir92]. inhabiting [VCC94]. inhibit [Ber95]. Inhibition [AB90, ABS92, BQLC97, OLM^{+}97, FLT^{+}98, PLLT94, ZFW97]. inhibitors [DTMD99, SH92]. inhibitory [JOW^{+}97, SNØ^{+}93]. iniae [DMS99]. initial [CNH^{+}97, EGTH92]. injectable [PRL95]. injected [JLR93, NLTN92]. injection [TRSD98]. inoculated [SBW94]. inoculation [DF94]. inoculations [FNTF95]. Instability [GL91]. integrated [DLP^{+}94]. Integumental [AO94]. intensively [OMF97, OBL^{+}99]. inter [YA95]. inter-renal [YA95]. Interaction [LLLS95, MF95, PS98]. interactions [Ain93]. interface [KT93a, KT93b, Tar90]. intermittent [SAMP97]. internal [DFS94]. internalization [GSC^{+}99]. International [Bro91]. intertidal [DEA94]. Intestinal [DIHS94, BRAP99, DLB95, HN94, JK90, JOW^{+}97, SOP96]. intestine [BK96, OJW^{+}96, OJW^{+}98, SASG94, vdIOK96]. intestines [BK96]. intracellular [JBRM98, VC93]. intraleucocytic [MD90]. intraperitoneal [ARX95, CMAM96, CRM^{+}96, DBIS96, PMD90]. intraperitoneally [JLR93, NLTN92]. intravascular [SN91]. invaderis [WRC95]. invasion [FLT^{+}98, LYGS97, RM92]. invertebrate [FHBN91]. Investigation [CNN^{+}94, ED91]. Investigations [DEA94, BFCH91, CN99, SD97a]. involved [LR97]. Involvement [WHC^{+}98, RM92]. involving [MOBF98]. iodide [YDOW97]. Iodophor [SKA99]. IPN [SBD^{+}92, SMBM95, TRSD98]. IPN-Sp [SMBM95]. IPNV [HSK94, JS95, LAB92, LLGJS94, MF95, SDS95, SA94, SY94, TRSD98, TT99]. Iredale [AL95]. Ireland [AGS^{+}92, MRD^{+}92, RD93, ZPVH97, ZPVH99]. iridovirus [MOBF98, RKMF97, SS90]. iridovirus-like [MOBF98]. iridoviruses [MMJ^{+}97]. Iron [JR98, BAR97a, Dav91a, SEMB^{+}99]. iron-regulated [Dav91a]. irritans [MMB93, MB95b]. ISA [DFK93, DF94, DFS94, MF95]. ISAV [NHWE95, NKKH97]. Isles [SMBM95, RR98]. isoelectric [WST96]. Isoenzymatic [Cai92]. isoenzymes [WST96]. isoforms [WST96]. isolate [AB90, BLHA92, MA99, SBMB95]. isolated [AHTY92, BAR95, BHS98, CPLR95, DES91, DHTC99, DMS99, DTS^{+}92, HSK96, HCM92, IFMR91, IYBF93, JS95, KRSE93, LAB92, LYLS95, LTC98, NBWO91, RK91, SA94, SR99, Wik90, YIA92, EA91]. isolates [Aus93, BHA94, BHP^{+}94, HSK94, MO95, MCRF94, NE91, SBB^{+}93, THBL96, TAD99]. Isolation [AAHY90, AE99, CFHR98, FH91, HRO90, LYGS97, MSM^{+}90, NG99, NLS^{+}90, RDMW98, SSD^{+}97, ZBM^{+}99, BENJ92, CNS91, EM93, F.92, GCS95, GHB91, GBH93, WPRM94, WR94]. Isometric [KN90a]. Isurus [FDS92]. Italian [DBCA97]. Italy [BAR95].

jack [AMN^{+}93, MCL97]. Janthinobacterium [AGS^{+}92]. Japan [BAR95, MMJ^{+}97]. Japanese [GMHK93, HSK94, HCW93, HLTK96, HLTK97a, LKM99, UUNY91, YI90]. japonica [GMHK93, HCW93, HLTK96, HLTK97a, HLTK97b, LKM99]. japonicus [TIM^{+}96, WTKC97]. Jastrzebski [JK90]. Jaundice [PCKS94]. Jaw [LMH^{+}97]. Jervis [HOF96]. johnsonae [CSM93]. Jordan [TMMA96b, TMMA96a]. junction [HSK96]. June [Bro91]. juvenile [ASR^{+}93, BP91, CSM93, CCC93, HM90, JSB90, OJW^{+}98, SBW94, SBMB95, UUNY91]. juveniles [YI90].

K1 [JOW^{+}97]. Kabata [BB94]. Katz [AFW94, FW96, FW98, KBW95, SBW94, TW90a, TW90b, VW96, ZFW97, ZW98]. Kaup [ZBM^{+}99]. kDa [BWC92, PBC^{+}90, VW96, WDCM92, ZW98]. kidney [ARD92, Ain93, BHB90, BFCH91, BMO^{+}98, CC91, DIB^{+}95, DISS96, DBIS96, GBH93, IKU98, JS95, JLR93, MAR99, OGFL92, SD97a, SM96, SSD^{+}97, SJGB98, SABòJ99]. killed [BCD90, DOF99]. killer [ISM^{+}98]. killer-like [ISM^{+}98]. killing [KSS91]. kinase [WST96]. Kinetoplastida [SKK90, TW90b, TW92, WT92]. kisutch [BDM91, GCS95, SAK91, TOK^{+}94]. Koch [CNS91]. Korea [PSL^{+}93]. kraft [SCE94]. Kroyer [BSWR93, EMM90, MSWB98, SSS^{+}99]. Kudoa [BP91, Lan91c, LTF92]. kuruma [WTKC97].

l [DS95]. L. [AAHY90, AHA92, AFJ^{+}97, APAFSP95, ALVS94, AMNH99, AEL^{+}99, ABJE93, Ath92, AS93a, AS93b, ASR^{+}93, BK96, BHG91, BHT92, Ber95, BB96a, BB96b, BAE^{+}91, BSHJ91, BRB^{+}94, BRAP99, BARH90, BR99, BBS96, BSWR93, Bru92, BRW92, BF93, BR92, BMEC97, BMO^{+}98, CMAM96, CRM^{+}96, CRG^{+}97, CA95, CSBCV98, CP96, CTC^{+}94, CGøL^{+}95, CFHR98, CNH^{+}97, CN99, CCC93, CD95, CSP98, CSBP^{+}99, CDGM99, CM92, CD98, DIHS94, DIB^{+}95, DS95, DLB95, DISS96, DBIS96, DFK93, DF94, DFS94, Dav98, DBCA97, DAL90, DL99, DMOR97, DM90, DHB91, DDGP^{+}96, DFGG^{+}99, DAKD98, DRP92, DHB^{+}95, ED91, EGTH92, ERJ91, EA91, FDS92, FTB90, FHMG95, FM98, FNTF95, FMY94, Gat91, GSS98, GG99, GHB91, GG97, GJS^{+}97, HEH^{+}90, HRO90, HMH96, HAH92, HN94, HM90, HWGV91, HNWE94]. L. [HN97, HCM92, IHKS90, IKU98, IFMR91, IRV^{+}91, IYBF93, JGOB95, JK90, JRV95, JS95, JSB90, JBWT92, JLR93, JM95, KT93a, KN90a, Kak98, KR95, KKP92, KRL96, KRSE93, KSPO99, LTGC99, LNFT95, Lan91c, LMTW95, LGSO97, LES^{+}93, LPO^{+}98, LSR90, LLP92, LJHE91, MD90, Mai92, MB94, MB95a, MHB94, MOBF98, MRD98, MSW94, MF95, MB98, MS91, MS90, Møy93, MLB95, NDQT95, NLTN92, NVSB94, NR97, NRV^{+}98, NC99, NHWE95, OCC^{+}95, OJW^{+}96, OGFL92, OJW^{+}98, PSB^{+}92, PFO^{+}98, PMS98, PRL95, PERL96, RHC^{+}93, RT94, RHL^{+}94, Rit97, RRER90, RPT^{+}93, RD93, RTEC94, RR98, RGA^{+}98, RDMW98, RMR95, SN91, SNØ^{+}93, SRRN94, SNó95, SLJ97, SE99, SBN^{+}91, SCE94, SLL96, SBAP90, SbFSAP92, SQTF90, SB93, SBD^{+}92, SMBM95, SW90, SM96, SI91, SSD^{+}97, Söv92, SBMB95, SAMP97]. L. [SHT96, SKK90, SJGB98, SABòJ99, SHB98, SCBB93, SSS^{+}99, SD97b, SOP96, SDB99, TNAF99, TRSD98, TT99, Tat91, TAD99, TL94, TRR96, VD98a, VD98b, VD99, VHCCGMOS98, Vir92, VFCV99, VC93, WGSH94, WHHL95, WBS^{+}96, WB94, WDV93, Wik90, WS90, YMM91, YLSW99, ZPVH97, ZPVH99, vdIOK96]. Labbé [ERSA98]. labrax [BR99, CSBCV98, DAL90, DAKD98, LGSO97, SBAP90, TAD99]. Labridae [GSS98, LTGC99, TL94]. labrosus [MB95b]. labyrinth [LTC96]. Lacépède [Che92b, BB94, Che92b]. lacking [TGK94]. lactoferrin [Kak98, SOAK93]. Lagenidium [RRJ^{+}96]. lake [SA94, SY94, CRG^{+}97, SQTF90]. Lamarck [HOF96]. lamellar [FMDC93]. laminaran [DIHS94, DIB^{+}95, DS95, DISS96, DBIS96, DMH^{+}98]. Lamna [FDS92]. lanceolatus [LPO^{+}99]. large [SW90]. largemouth [Che92b]. larger [SOF^{+}92, SOS94]. larvae [BHT92, BB96a, BB96b, Gat91, JRV95, SD97b, YI90]. larval [ASN90, BB94, GHdB90, MA99, RRJ^{+}96]. Lasee [LS93]. latent [IRM^{+}96]. lateral [KN90a]. Lates [CSM93, GHdB90]. laticeps [KT93b]. latipes [HHKE90, HIH^{+}96]. Latrobe [SPC90]. Laveran [BB97, BB99, DEA94, DMTM95, DTMD99, ED91, SKK90, WB94]. Lawrence [LMH^{+}97]. layer [BENJ92, TGK94]. layers [BARH90]. leading [ASR^{+}93]. Lectin [DBCA97, IHKS90, Ain93, dMMMK96]. Léger [AS93a]. lenticular [CCC93]. leopardus [BB94]. Lepeophtheirus [BSWR93, EMM90, JSB90, JSW92, JBWT92, MSWB98, Rit97, SSS^{+}99]. lepidota [AMM95]. Lernea [WS90]. lesions [AS92b, BHS98, BBW98, CCC93, FJS90, FDS92, FNTF95, NDBL94, SFB^{+}91b, SFB^{+}91a, SAMP97, VD98b, VCC94, VHCCGMOS98]. lethal [HCM92, LE91a, Mai92, RFN93]. lethality [SMSA91]. leucocyte [JMK^{+}95]. Leucocytes [DFK93, Ain93, DBIS96, DF94, DES91, EC91, FW98, ISM^{+}98, JS95, RT94]. level [ASR^{+}93, BR99]. levels [SGNG93, SB94]. Leven [RTEC94]. lice [BSWR93, RRS93, SSS^{+}99]. Lichtenstein [RKMF97]. life [Ber95, CSBCV98, GCS95, MMB93, MSWB98]. life-cycle [MSWB98]. Light [JK90, MSW94, WTKC97, DMOR97, SHB95]. like [AAHY90, AHA92, BLP96, CGA93, CTC^{+}94, CNS91, GHdB90, ISM^{+}98, KDL95, LDCA92, LTBL94, MOBF98, RD93, SB94, SSM95]. Limanda [BF93, DM90]. limitation [BAR97a]. line [BBCT^{+}94, CHL99, Fre96]. linked [CATR97, RPT^{+}93, SOS94, SBW94, VW96]. lipid [WHHL95]. lipoidosis [PBP94]. Lipophrys [AEL^{+}99]. lipopolysaccharide [AHA92, DSA^{+}98, HCM92, RSTE93, SKQ^{+}99, SABòJ99]. lipopolysaccharide-free [HCM92]. lipopolysaccharides [LE91b]. liposomes [PMD90]. lipped [MB95b]. liquefaciens [AB90]. liquefaction [PM91]. listeriosis [RMT^{+}92]. live [BCD90, FNTF95, SBW94, TGK94]. Liver [BMEC97, AMM95, DHM97, SMG96]. livers [BF93]. lividum [AGS^{+}92]. loach [CW91]. load [EGTH92]. lobsters [RVSK99]. Localization [SABòJ99, BARH90, DLB95, SOP96, SDB99]. Loch [RTEC94]. Lom [MEMSB99]. Loma [AEL^{+}99, BSBD99, SKA99, SBJ^{+}98, SDM^{+}98, SBD99]. lomi [PI91]. long [YKI95]. long-term [YKI95]. longevity [WR98]. Loubés [AEL^{+}99]. louse [JSW92, MSWB98]. low [ASR^{+}93, BAHL91, BHS98, DHTC99, Col95]. low-level [ASR^{+}93]. Low-pH [Col95]. Loxothylacus [KDC95]. LPS [DS95, DSA^{+}98, RSTE93]. LR [BMEC97]. lucius [BRB^{+}94, Wik90]. lupus [HMH96]. luteus [AS92a]. Lutjanus [SW90]. Lymphoblastic [HHKE90]. lymphocystis [CD95]. lymphocytes [MS90]. Lymphoma [BS98, BRB^{+}94, HHKE90]. lysins [AMNH99]. lysis [LE91b]. lysophospholipids [RSTE93]. lysozyme [SOP96]. lytic [KWJ90].

MAb [ZFW97]. MAb-001 [ZFW97]. Maccullochella [PI91]. mackerel [Dav98, MD90]. Macrobrachium [ASN90, LTC98, TWC99]. macrocephalus [PCKS94]. macroglobulin [SNA^{+}92, SNØ^{+}93]. Macrophage [BB99, PSB^{+}92, RM95a]. Macrophages [PM91, DS95, SM96, SSD^{+}97, SJGB98, WCS90]. Macropsobrycon [RMRR96]. macropterus [MW94]. maculata [Che92b]. mahi [Lan91c, LMTW95]. maintenance [EM93]. major [FLT^{+}98, GL91, KRK91, PBC^{+}90, SEMB^{+}99]. Malachite [ACH93, GACHF91, MSGA95]. Malaysia [AHTY92]. Malaysian [ASN90]. malignant [BRB^{+}94]. Malmberg [SHB98]. mammalian [BB97]. Man [ASN90, LTC98, TWC99]. manipulation [SBD99]. manteri [SASG94, BB94]. marcescens [BTL^{+}92]. marianus [HR95]. marine [APR99, BSWR93, Bru92, CA95, DSA^{+}98, DHM97, EA94, FHBN91, HR95, HSK94, HSK96, LMBS92, STP^{+}98, WR98]. marini [HN97]. marinus [ABC^{+}98]. maritimus [HSP97, PSM^{+}96]. marker [SGNG93]. markers [IHKS90]. marmoratus [CF91, RM95a]. marron [Lan91a, Lan91b, LT92]. Marteilia [AAL95]. Martens [JR98, JBRM98]. Mass [ASN90, CLK^{+}97, GHdB90, JRV95, LGSO97]. Matthews [LES^{+}93]. maturation [WHC^{+}98]. maturing [SNó95]. Maurant [AEL^{+}99]. maximus [DDGP^{+}96, AHA92, ALVS94, ASR^{+}93, BRAP99, DAL90, DHB^{+}95, FNTF95, Gat91, LNFT95, LES^{+}93, LPO^{+}98, NDQT95, OJW^{+}96, OJW^{+}98, PFO^{+}98, SBN^{+}91, TNAF99]. Mayo [ZPVH97, ZPVH99]. MBV [LTKC93, LTKC95, LTC96, VPSL93]. McConnaughey [FHBN91]. McGregor [TMMA96b, TMMA96a]. meal [BK96]. means [WST96]. measure [AOES99]. measured [SGOH94]. mebendazole [BRW92]. mebendazole-resistant [BRW92]. mechanism [BR92, RSTE93]. mechanisms [ABC^{+}98, CA92, DIB97, KEY91]. medaka [HHKE90, HIH^{+}96]. media [DMTM95, EM93, GHB91, PSM^{+}96]. mediated [IYBF93, SLL96, TW90b]. Mediterranean [SBAP90]. medium [BB97, BB99, F.92, GBH93, WB94]. melanin [JHL^{+}91]. melanin-containing [JHL^{+}91]. Melano [RM95a, PSB^{+}92]. Melano-macrophage [RM95a, PSB^{+}92]. melanopterus [SA90]. member [LHS94]. membrane [BLHA92, Ber90, BA91, CDPS90, Col95, Dav91b, Dav91a, FW96]. membrane-associated [CDPS90]. membranes [RSTE93]. membranous [LTC96]. menhaden [SPC90]. mercuric [PS93]. Mesnil [BB97, BB99, DEA94, DMTM95, DTMD99, ED91, SKK90, WB94]. mesodermal [TIM^{+}96]. metabolism [KBW95]. metacercariae [HN97, SASG94, VHCCGMOS98]. metallo [ZW98]. metallo-protease [ZW98]. metalloproteases [AEJ95, AE99, SJH^{+}93]. Metastatic [FCS91]. methionine [NHOJ93]. method [BR99, Bri95, BSWR93, KEY91, LWHD97, LLGJS94, VPSL93]. methodology [WR94]. methods [NR97]. Mexico [SMG96, VHCCGMOS98]. mice [CATR97]. microbioassay [PI93]. microcephalus [BBW98]. Micrococcus [AS92a]. microcystin [BMEC97]. microcystin-LR [BMEC97]. microcystins [CMAM96, CRM^{+}96]. Microcystis [CRG^{+}97]. Microgadus [LMH^{+}97]. Microgemma [LPO^{+}99]. Micropterus [Che92b]. microscope [MSW94]. microscopic [DMOR97, JK90, TWC99, TRR96, WTKC97]. microscopy [HEHR91, SGOH94, SHB95]. Microspheres [DLB95]. Microspora [FTB90, HLTK96, HLTK97b, Lan91a, LES^{+}93]. Microsporidia [SKA99]. Microsporidian [DRP92, CA95, LPO^{+}99, PM91, SBJ^{+}98].

Microsporidiosis [Lan91b, LT92]. midae [DHB91]. migration [SI91, WR98]. Mikailov [LB98]. Mikrocytos [AL95]. milky [Lan91c]. mill [SCE94]. Mimachlamys [HOF96]. mirror [MB94]. Misgurnus [CW91]. Mitchell [FTN97, MSM^{+}90, PI91, SDS90]. Mitchill [AFW94, CC91, MO95, RM92, SDS95, SEMB^{+}99, SA95, SHB98, Tho90]. mitogen [EC91]. mitogen-stimulated [EC91]. mitogens [DFK93]. MMV [TWC99]. Mn [ISM^{+}98]. mobile [SK90]. model [FAAE^{+}99, SBD99, SK90]. Modified [Tat91, ASN90]. modulation [TW90a]. Mokoan [CRG^{+}97]. Molecular [AAL95, DMS99]. molecules [SABòJ99]. monitor [OGFL92]. Monoclonal [SBB^{+}93, ARD92, BAR97b, CCH^{+}94, CFS93, CND90, ERJ91, FW96, HAC^{+}92, HIMK93, LAB92, MCRF94, MC94, NS96, SA94, VW96, WLHD97, ZFW97, dMMMK96]. monoclonal-defined [ERJ91]. monodon [RRJ^{+}96, CPLR95, LTKC93, LTKC95, LTC96, RK91, TIM^{+}96, VPSL93, WTKC97]. monodon-type [LTKC93, LTKC95, VPSL93]. Monogenea [CSBCV98, CMMM95, KT93a, SHB98, WR98]. monogenean [RQW92, Tho90]. monogeneans [BRW92, SB93]. monophasic [EM93]. mordax [MSM^{+}90]. Moreton [Rou94a]. morhua [DISS96, ERJ91, KRSE93, MB98, SSD^{+}97, SJGB98, WHHL95]. moribund [AS91, AS92a]. Morone [BP91, TNB^{+}93]. Morphogenesis [LTC96, GWFE97]. Morphological [LLLS96, HIH^{+}96, vdIOK96]. Morphology [SHB95, BP91]. Morphometric [SSJ97]. morphopathological [GBG^{+}93, Góm98]. mortahty [BSB91]. Mortalities [PSB^{+}92, RHC^{+}93, ASN90, ASR^{+}93, BLH^{+}90, CLK^{+}97, GHdB90, LGSO97]. mortality [AL95, CNS91, JGOB95, JRV95, RGA^{+}98, SBD^{+}92, UUNY91, VC93]. mortem [Lan91c]. mossambicus [SR99]. motile [DSB90]. mouth [ALVS94]. mouuscan [MO95]. mucus [BARH90, FOM^{+}95, JOW^{+}97]. mucus-secreting [BARH90]. mud [KDC95]. Mugil [Vir92]. mullet [MB95b, Vir92]. multifiliis [HM90, BCD90, EM93, LYD94, LDCA92, SLL96]. multilocus [BHP^{+}94]. Multiple [MHB94]. Multiple-dose [MHB94]. Multiplication [JS95]. multivalent [SEMB^{+}99]. Multivalvulida [BP91, HB96, Lan91c]. mummichog [VFCV99]. Murray [PI91]. Muscle [LDTC93, HB96, KN90a, PM91, SA90, TWC99]. mutation [BAHL91, Mai92]. Mycobacteriosis [CLC90, GBG^{+}93]. Mycobacterium [CATR97, CAR97, CYA^{+}98, HLTS93, MHM95]. Mycoplasma [SK90]. mycoplasmas [Fre96]. Mycotic [RWC93]. mykiss [AOES99, ACH93, Ath92, AS91, AGS^{+}92, AS92a, AS92b, BSBD99, BBCT^{+}94, BS98, BLW95, BMEC97, CGA93, CH92, DES91, DES92, DT96, DOF99, DT93, DRL94, FW98, FOM^{+}95, FWM94, GCS95, IGH92, ISM^{+}98, JMK^{+}95, KR95, KBW95, LW97, LB98, LOF96, MB92, MSGA95, NKKH97, OLM^{+}97, PMD90, RMR95, SJMPP91, SOAK93, SYAK95, SSJ97, SBN^{+}91, SB94, SEMB^{+}99, SBW94, SAMP97, SBJ^{+}98, SDM^{+}98, TW90a, TW90b, TW92, TSL^{+}94, TNB^{+}93, VW96, WFS^{+}95, WVG^{+}98, WPRM94, WCS90, WST96, YA95]. Myoliquefaction [Lan91c]. myoliquefactive [LTF92]. myositis [LOF96].

Mystus [MW94]. Myxidium [AS93a]. Myxobolidae [MBM96]. Myxobolus [MBM96, RVA98, Rou94a]. Myxosporea [BP91, DBCA97, HB96, Lan91c, MBM96, MEMSB99, Rou94a, SBAP90, SbFSAP92, YLSW99]. Myxosporean [BRAP99, AS93b, CC91, HB96, IGH92, LTF92, MSW94, YDOW97]. myxosporeans [AS93a]. myxosporidian [SA90]. myxosporiosis [RVA98]. Myxozoa [YLSW99].

n [CD98]. N1 [MC94]. naive [Sub93, WS90]. namaycush [SA94, SY94]. naphthyl [LWHD97]. Nardo [BGMC96]. nasus [FDS92]. nasuta [FDS92]. native [FAAEC99]. natural [ISM^{+}98, SFB^{+}91a]. Naturally [ZPVH97, HK98, KT93a, Tar90]. nature [HCM92]. necator [TSL^{+}94]. necrosis [AMN^{+}93, BB96a, BB96b, CLK^{+}97, CHL99, CFS93, CND90, CNH^{+}97, CSP98, DLP^{+}94, DHS94, DRL94, EHH91, GSS98, HEH^{+}90, HLHW98, HCW93, HEHR91, JGOB95, JS95, JM95, LGSO97, LAB92, LKM99, LLGJS94, MB97, MO95, MC94, MF95, PSL^{+}93, RM95b, SJMPP91, SOF^{+}92, SOS94, SDS90, SDS95, SA94, SY94, SS90, SBD^{+}92, SHT96, TRSD98, TT99, THBL96, TAD99, WPRM94, WHC^{+}98, YI90]. Necrotic [LOF96, OMF97]. Needle [DHM97]. Nematoda [SMG96]. nematode [HB98]. Neoechinorhynchus [Tar90]. neoplasm [PGLH96]. Nephrocalcinosis [JRV95]. nerka [HK98]. nervous [AMN^{+}93, CLK^{+}97, CHL99, LGSO97, TAD99, YI90]. neutralizing [FAAEC99]. neutrophil [AOES99]. nifurpirinol [KGW90]. Nigeria [AA91]. Nile [CTC^{+}94, CYA^{+}98]. nilotica [CYA^{+}98]. niloticus [CP96, CTC^{+}94, Mai92, MOBF98]. Nilsson [PM91]. nitrocellulose [LTL96]. nitrocellulose-enzyme [LTL96]. No [FOM^{+}95]. No-mucus [FOM^{+}95]. noble [SQTF90, IGH92]. Nocardia [Che92b, Che92a, CW93]. Nodavirus [SR99, BR99]. nodular [VCC94]. Non [CYA^{+}98, DIB97, Aus93, PRCC96, RHC^{+}93, RRER90, WVG^{+}98, YKI95]. non-01 [RHC^{+}93]. non-culturability [PRCC96]. non-salmonids [Aus93]. Non-specific [CYA^{+}98, DIB97, RRER90, WVG^{+}98, YKI95]. Nordmann [EMM90]. norfloxacin [LH90]. normal [AO94]. north [AEL^{+}99, DM90, SRL^{+}96, TLMM94]. north-west [AEL^{+}99]. Northeast [MD90]. northern [BRB^{+}94, KDC95, VRK92, AGS^{+}92]. Northwest [MD90]. Norway [CND90, CFHR98, NVSB94, PM91, SQTF90]. Norwegian [MCRF94, MC94]. notifiable [SRL^{+}96]. nov [FTB90, Lan91a, LPO^{+}99, MBM96, RVA98, SBAP90, WRC95]. Nova [CC91, VCC94, DWV94]. novel [CNN^{+}94]. NS [HSK96, WHC^{+}98]. nucleoprotein [CFS93]. nucleotide [Ner96]. numbers [DWV94]. Nutritional [Bro90, WBS^{+}96]. NW [RMP^{+}94].

O [TGK94]. O-antigen [TGK94]. obliterates [FMDC93]. Observation [HNWE94, RD93]. Observational [RR98]. observations [CD98, DM90, IKU98, MSW94, RHL^{+}94, SKK90]. occur [Møy93].

Occurrence [AA91, DL99, VRK92, BLP96, BFCH91, HLJ^{+}91, SH92, WDCM92]. occurring [BDM91, ZPVH97]. occurs [SDM^{+}98]. ocellatum [HB96, HB98]. ocellatus [CCC93, CD95, MHM95, PGLH96]. octopiana [ERSA98]. Octopus [FHBN91]. Ocular [BBW98, CF91]. Odonthestes [HLTS93]. off [HIH^{+}96, WBS^{+}96]. oil [SEMB^{+}99]. oil-adjuvanted [SEMB^{+}99]. oligochaete [MEMSB99]. Oligogonotylus [SASG94]. olivaceus [UUNY91]. Ommatokoita [BBW98]. oncogenicity [SMSA91]. Oncorhynchus [AOES99, ACH93, Ath92, AS91, AGS^{+}92, AS92a, AS92b, BFPW90, BSBD99, BBCT^{+}94, BDM91, BS98, BLW95, BMEC97, CGA93, CH92, DES91, DES92, DT96, DOF99, DT93, DRL94, FW98, FOM^{+}95, FWM94, GCS95, HK98, IGH92, ISM^{+}98, JMK^{+}95, KR95, KBW95, LW97, LB98, LOF96, MB92, MSGA95, NKKH97, OLM^{+}97, PMD90, RMR95, SJMPP91, SAK91, SOAK93, SYAK95, SSJ97, SBN^{+}91, SB94, SKA99, SEMB^{+}99, SBW94, SAMP97, SBJ^{+}98, SDM^{+}98, TOK^{+}94, TMMA96b, TMMA96a, TW90a, TW90b, TW92, TSL^{+}94, TNB^{+}93, VW96, WFS^{+}95, WVG^{+}98, WPRM94, WCS90, WST96, YA95]. one [MF91, RMP^{+}94]. one-year [RMP^{+}94]. ongoing [SNó95]. onset [SDM^{+}98, TW92]. Oomycotina [RM92]. Ophicephalus [Sub93]. Oplegnathus [YI90]. opportunistic [KBMGL90, NLS^{+}90, RGA^{+}98]. Opsanus [KGW90]. Optimising [SJGB98]. oral [ARX95, DAL90, ISM^{+}98, IKU^{+}96, Kak98, PV93, SOAK93, SYAK95, SSS^{+}99, YKI95]. ordalii [ASR^{+}95, GCO91, PLLT94]. Oreochromis [CYA^{+}98, LLLS95, Mai92, MOBF98, SR99]. Oreochronuis [CTC^{+}94]. organ [SK90]. organism [BLP96, CTC^{+}94, CNS91, JBRM98, LMTW95, MAR99, RD93]. organisms [KDL95]. organophosphate [JSW92]. organs [PBP94, PMS98, Söv92]. origin [VCC94]. ornamental [LSL91]. Oroochromis [CP96]. Oryzias [HHKE90, HIH^{+}96]. oscar [MHM95, PGLH96]. Osmerus [AAHY90, MSM^{+}90]. Osteichthyes [ED91]. osteochondritis [OMF97]. other [IR91, LR97, LLN94]. otherwise [AS92b]. outbreak [AS92a]. outbreaks [AR95, AS91, BLSC93, CNN^{+}94, NHWE95, SFB^{+}91a]. outer [BLHA92, Ber90, BA91, Dav91b, Dav91a]. ovarian [KRL96]. oxolinic [SE99]. oxolonic [LH90]. oxyrinchus [FDS92]. Oxytetracycline [CPLR95, KN90b, TNAF99]. oyster [AL95, ABC^{+}98]. oysters [AAL95].

Pacific [FJS90, Jon90]. Pagothenia [FMDC93]. Pagrus [DHM97, RQW92, WR98]. Pallas [BHB90]. Panaque [KDL95]. pancreas [BL96, CFHR98, MRD98, MRD^{+}92, RDMW98]. pancreatic [BB96a, BB96b, CND90, CSP98, DLP^{+}94, DHS94, GSS98, HEH^{+}90, HLHW98, HCW93, JGOB95, JS95, LAB92, LLGJS94, MB97, MO95, MC94, MF95, RM95b, SJMPP91, SOF^{+}92, SOS94, SDS90, SDS95, SA94, SY94, SBD^{+}92, SHT96, TRSD98, TT99, THBL96, WHC^{+}98]. panophthalmitis [BFPW90]. papilloma [MS90]. paracolo [GMHK93]. paradoxum [KT93a]. Paralichthys [UUNY91]. Paramoeba [FHMG95]. Paramyxea [AAL95]. parasite [AMM95, Dav98, HB96, HB98, KT93a, KT93b, LTF92, LPO^{+}99, LDCA92, MBM96, MSW94, PM91, SBJ^{+}98, SA90, Tar90, Tho90, VW96, VC93]. parasites [APAFSP95, CM92, ERSA98]. parasitic [BBW98, CSBCV98, CSBP^{+}99, UUNY91]. parasitica [WR92]. parasitism [SQTF90]. parasitizing [CMMM95, JK90]. parastacida [Lan91a, LT92]. Parastacidae [Lan91a, Lan91b]. parauberis [DDGP^{+}96]. parr [IRV^{+}91, TRR96]. parrotfish [YI90]. Partial [AEJ95, SJH^{+}93, CND90, HYHA98]. particles [AAHY90, GHdB90, LDCA92]. particular [DIB97]. particulate [ARX95]. passive [AMW96]. Pasteurella [BAR95, BAR97a, BAR97b, YIM^{+}97]. pasteurellosis [BAR97b, ZBM^{+}99]. pathogen [AAAR98, AB90, BAR97a, BTL^{+}92, CSM93, EA94, GSC^{+}99, KBMGL90, LLLS95, NLS^{+}90, OJW^{+}96, PLLT94, RGA^{+}98, SPC90, WR94, WRC95]. pathogenesis [SRRN94]. Pathogenic [SMG96, AHA92, Ber95, BQLC97, Bro90, CGA93, GRR92, GWFE97, IR91, KBE^{+}99, KBW95, NE91, SBW94, SGOH94, ZFW97, ZW98]. Pathogenicity [AMN^{+}93, LR97, AKK96, Che92b, SY94, SK90, VRK92]. pathogens [ABS92, BSH^{+}90, BA91, BSHJ91, CH92, CSBP^{+}99, KWJ90, OSG95, SRL^{+}96]. Pathological [GACHF91, FWM94, NDBL94, RV91]. Pathology [BHB90, BB94, LNFT95, NDQT95, SFB^{+}91b, SFB^{+}91a, HSP97, LE91a, MRD^{+}92, PSB^{+}92, RTEC94, SMBM95]. pattern [GRR92]. patterns [Ber90, NDBL94]. PCR [KKS^{+}98, LLGJS94, TIM^{+}96]. pearl [RVA98]. pearl-spot [RVA98]. Pecten [HOF96]. peeli [PI91]. pejerrey [HLTS93]. pen [SAK91]. pen-cultured [SAK91]. penaeid [AAAR98, KKS^{+}98, LTBL94, LLN94, LTL96, LTC96]. Penaeus [CPLR95, LTKC93, LTBL94, LTKC95, LTC96, MA99, RRJ^{+}96, RK91, TIM^{+}96, VPSL93, WTKC97]. peptides [KWJ90, KBE^{+}99]. peptidoglycan [IKU^{+}96]. peranal [DBIS96]. perch [BP91, FTN97]. Percichthyidae [PI91]. performance [KBW95]. period [RMP^{+}94]. peripheral [MS90]. Peritoneal [JLR93, AOES99]. peritonitis [SEMB^{+}99]. Perkinsus [ABC^{+}98]. peroral [DBIS96]. Peroxidase [AOES99, JHL^{+}91]. peroxidase-antiperoxidase [JHL^{+}91]. peroxide [DOF99, MSWB98]. persistence [OJW^{+}98]. Persistent [VCC94, MF95]. persistently [FAC99]. Peters [SR99]. pfeifferi [AS93a]. pH [Col95]. phagocytes [LLLS95]. Phagocytic [PR90]. Pharmacokinetic [SLJ97, ACH93, MHB94, SE99]. Pharmacological [TSL^{+}94]. phase [EGTH92]. phases [CC91]. pholis [AEL^{+}99, ED91]. phosphate [WFS^{+}95]. phospholipase [CDPS90, HCM92]. phospholipids [Col95]. Photobacterium [ZBM^{+}99]. phylogeny [HYHA98]. Physalopteridae [HB98]. physiological [NBWO91]. PI [YDOW97]. Pickford [FHBN91]. Picorna [GHdB90]. Picorna-like [GHdB90]. picornavirus [AAHY90]. picornavirus-like [AAHY90]. pigment [FH91]. pigment-producing [FH91]. pigmented [AS91]. pike [BRB^{+}94, Wik90]. Pisces [Jon90, RMRR96, RM95a, Rou94b, Rou94a]. piscicida [BAR97a, BAR97b, NG99, YIM^{+}97, ZBM^{+}99]. piscicola [TNB^{+}93]. piscidda [BAR95]. piscine [ZW98]. Piscirickettsia [AFJ^{+}97, GCS95, LF94a]. PJ [TIM^{+}96]. PKD [BFCH91, MAR99]. PK‘ [ARD92]. PKX [HK98, MAR99, dMMMK96]. Planococcus [AB90, AS92a]. plaque [SOS94]. Plasma [KRL96, SGNG93, SNA^{+}92, SEMB^{+}99, SKQ^{+}99, YA95]. Plasmid [IYBF93, NBWO91]. Plasmid-mediated [IYBF93]. Plasmids [AHTY92]. Plecoglossus [KEY91]. plecostomus [KDL95]. Plectropomus [BB94]. Plehn [RHL^{+}94, SI91]. Pleistophora [FTB90, HLTK96, HLTK97b]. pleistophorid [Lan91b]. Pleistophoridae [Lan91a]. plumbeus [BGMC96]. plymuthica [AS92b, NLS^{+}90]. Poey [CF91]. pollution [BSM^{+}99]. polyacrylamide [LWHD97]. polyamine [DTMD99]. polyclonal [LAB92]. Polymerase [SD97a, TIM^{+}96, GWG99, KKS^{+}98, LLGJS94, LTKC93, MHM95]. polymorphic [MEJ^{+}94]. Polysaccharide [YMM91]. Polysaccharide-induced [YMM91]. polystyrene [DLB95]. Pomacanthus [LHS94]. Pomacentridae [NG99]. ponds [AA91]. population [SJMPP91, SCE94, SA95]. populations [AOES99]. portal [OJW^{+}96]. Portugal [CD98, DEA94, ED91, SRL^{+}96]. possible [BSB91, CRG^{+}97, SRRN94, VC93]. post [JGOB95, Lan91c, MA99, SBD^{+}92, SMBM95, SHT96, TRSD98]. post-larval [MA99]. post-mortem [Lan91c]. post-smolt [JGOB95]. post-smolts [SBD^{+}92, SMBM95, SHT96, TRSD98]. posterocolpa [Tho90]. postponement [PRCC96]. postulate [CNS91]. potassium [MLB95]. potential [AFW94, BTL^{+}92, BA91, CSP98, SGNG93, Tat91]. potentiated [SLJ97]. potentiates [BB99, CA92]. pout [PM91]. practices [RRS93]. prawn [ASN90, RRJ^{+}96, TWC99]. prawns [LTC98, RVSK99, VPSL93]. praziquantel [Tho90]. pre [JBWT92]. pre-adult [JBWT92]. predictive [SBD99]. prednisolone [SKQ^{+}99]. preferred [CCD95]. preliminary [GACHF91, PMD90, SbFSAP92]. Prescreening [OCC^{+}95]. Presence [NHWE95, NK93, OCC^{+}95]. presporogonic [CC91]. presumptive [AS92a, AS92b]. Prevalence [MD90, NC99, SY94, Vir92]. prevent [SDS95]. preventing [SKA99]. previously [DM90]. primarily [MF91]. primary [BSBD99, BWC92, BBCT^{+}94, GHB91, MB95b]. primers [SD97a]. principal [LSAB99]. Prionace [FDS92]. probable [RSTE93]. probe [LTKC93]. probes [CMMM95, DHS94, MAR99]. probiotic [ASR^{+}95]. problems [RV91]. procedure [BENJ92, KKS^{+}98]. process [GSC^{+}99, SNA^{+}92, SI91]. produce [Dav91a]. produced [Bro90, GL91, MB92, MBM96, MC94, NDBL94, NE91, PLLT94, SH92]. produces [Dav91a, RSTE93]. producing [AB90, FH91]. product [Tat91, ZH96]. Production [BAR97b, NS96, RFN93, EA91, FW98, HLTK97a, JOW^{+}97, ZH96, ZW98]. products [CATR97, CAR97, CYA^{+}98, FNTF95, HCM92, KON90, MA99, SNØ91]. profile [KRL96]. profiles [NBWO91]. progression [ABC^{+}98, DRL94, TOK^{+}94]. Progressive [BFPW90, ASR^{+}93]. prokaryotic [JBRM98]. Proleptus [HB98]. proliferation [OJW^{+}98]. proliferative [ARD92, BFCH91, MAR99, SD97a]. Propagation [SM96]. properties [BAR95, GMHK93]. Prophylactic [DMH^{+}98]. propidium [YDOW97]. proposal [BSM^{+}99]. protease [ABJE93, BWC92, CDPS90, EBL^{+}94, PBC^{+}90, SNA^{+}92, SNØ^{+}93, WDCM92, ZW98]. Proteases [CT91, EMM90, NE91, SH92, ZFW97]. Protection [DWV94, SBN^{+}91, WCS90, BSBD99, BL96, GG97, IRM^{+}96, LSR90, SBW94, YMM91, YKI95]. Protective [FMY94, GMHK93, SK97, KEY91]. protein [Ber90, BENJ92, CCD95, Dav91b, FAAE^{+}99, Møy93, NDBL94, SGNG93, SKQ^{+}99, WHC^{+}98, SABòJ99]. protein-types [Dav91b]. proteinase [WHC^{+}98]. proteins [BLHA92, BWC92, BA91, Dav91a, LF94b, SBB^{+}93, SEMB^{+}99]. Protista [PI91]. protocol [BSM^{+}99]. Protozoa [ED91, FTB90, SKK90]. Protozoan [APAFSP95, Dav98]. protozoans [FHBN91]. pseudoalcaligenes [AS92b]. Pseudoalteromonas [NG99]. Pseudocaranx [AMN^{+}93]. pseudomonad [SD93]. Pseudomonas [AS92b, BQLC97, CSP98, DFGG^{+}99, FH91, GRR92, SAAH90]. pseudophyllidean [HN94]. psychrophilus [LOF96, OMF97]. Pterophyllum [RKMF97]. pulsed [LSAB99]. punctatus [AM92, Ain93, ARX95, BLSC93, BCD90, CT91, CP96, CA92, DGK99, GWG99, KLW^{+}98, LYD94, PV93, SK97, WBA91]. Puntazzo [APR99, DAKD98]. purification [AEJ95, EHH91, SJH^{+}93]. purified [ABJE93, CGøL^{+}95, SABòJ99]. putative [CSM93]. putrefaciens [SAAH90].

quadricarinatus [JR98, JBRM98]. quality [DHTC99]. quantification [DES92, DHB^{+}95]. quinolone [DHB^{+}95]. quinqueradiata [IKU^{+}96]. Quoy [SA90].

rabbit [FW98]. rabbitfish [SAAH90]. rabbits [CATR97]. radiation [BR92]. radicals [NVBK93]. Rafinesque [AM92, Ain93, ARX95, BLSC93, BCD90, CT91, CA92, DGK99, FDS92, GWG99, KLW^{+}98, LYD94, PV93, SK97, WBA91]. rainbow [AOES99, ACH93, Ath92, AS91, AGS^{+}92, AS92a, AS92b, BSBD99, BBCT^{+}94, BS98, BLW95, BMEC97, CGA93, CH92, DES91, DES92, DT96, DOF99, DT93, FOM^{+}95, FWM94, GCS95, IGH92, ISM^{+}98, JMK^{+}95, KR95, KBW95, LW97, LB98, LOF96, MB92, MSGA95, NLS^{+}90, NKKH97, OLM^{+}97, PR90, PMD90, RMR95, SJMPP91, SOAK93, SYAK95, SSJ97, SBN^{+}91, SB94, SEMB^{+}99, SBW94, SAMP97, SBJ^{+}98, SDM^{+}98, SK90, TW90a, TW90b, TW92, TSL^{+}94, TNB^{+}93, VW96, WFS^{+}95, WVG^{+}98, WPRM94, WCS90, WST96, YA95]. raised [AFJ^{+}97, DGK99]. Raja [FDS92]. ranarum [AO94]. range [RVSK99]. rapid [BAHL91, BR99, HAC^{+}92, TT99, VPSL93]. rate [SDM^{+}98, TOK^{+}94]. rather [DT96]. Ratinesque [CP96]. ray [Tho90]. reacting [FH91]. reaction [GWG99, KKS^{+}98, LLGJS94, LTKC93, MHM95, SD97a, TIM^{+}96]. reactive [HIMK93]. reactivity [ERSA98, PRL95]. reared [BHS98, Bru92, CLK^{+}97, KLW^{+}98, LGSO97, OMF97, OBL^{+}99, PI91, YI90]. rearrangements [LGL^{+}98]. recessive [Mai92]. Recombinant [ZH96]. record [BAE^{+}91, HHS93]. records [LTF92]. recovered [AB90, MA99, WS90]. Recovery [AS91, AGS^{+}92, AS92a, AS92b, Aus93]. recrudescence [IRM^{+}96]. red [CCC93, CD95, FCC92, RSTE93, Vir92]. redclaw [JR98]. Redescription [LPO^{+}99]. redmouth [AS92b, DGK99]. Reduced [JSW92]. reducing [ASR^{+}95]. Reduction [Kak98]. reef [SA90, HB96]. Reeve [HOF96]. reference [AO94, DIB97, MCRF94, MC94]. Regan [LDTC93, MCL97]. regions [HSK96]. Regression [MS90, BK96]. regulated [Dav91a]. Reichenow [Dav90]. reinfection [FHMG95]. related [BA91, CSP98, DT96]. relatedness [HSK94]. relation [CT91]. Relationship [TOK^{+}94, BSB91, LLN94, MB97]. relationships [DTS^{+}92]. Relevance [SA95]. reliable [Bri95]. renal [PGLH96, YA95]. Renibacterium [BHB90, BHG91, DGR97, GSC^{+}99, GL91, GHB91, GBH93, JHL^{+}91, OCC^{+}95, SAK91, SA95]. renicola [MEMSB99]. repair [BR92]. repeated [GACHF91]. Replication [NKKH97, DT93]. report [BFPW90, CNH^{+}97, FOM^{+}95, MOBF98, PMD90, SbFSAP92]. reproductive [DT96]. required [FAAEC99]. requirements [Bro90]. requires [NHOJ93]. rerio [RMT^{+}92]. RES [DIB97]. resembling [MS91]. Reservoir [LTF92, AFW94]. Residues [MSGA95, KN90b]. Resistance [IMR93, AHTY92, BHA94, FHMG95, FM98, HWGV91, IFMR91, IYBF93, IKU^{+}96, LYLS95, NDBL94, RRER90, SOAK93, SYAK95, SGNG93, SB93, SBJ^{+}98, WGSH94, WVG^{+}98, WDV93, WS90]. resistant [BRW92, HLJ^{+}91]. respect [LF94b]. response [AHA92, CYA^{+}98, DFK93, GJS^{+}97, HN94, HM90, HLTK96, HLTK97b, JLR93, Kak98, KRK91, LES^{+}93, LJHE91, RHL^{+}94, RMR95, SSJ97, SLL96, TNAF99, TW90a, YA95]. responses [ARX95, ABJE93, DES91, ERJ91, HAH92, HN97, JMK^{+}95, TW90b]. responsiveness [RT94]. results [GBH93, MF91]. Retention [BWC92, PERL96, PRL95]. reticuloendothelial [DIB97]. Retinal [CCC93]. reverse [LLGJS94]. review [Ano95a, Ano95b, Ano98, RRS93, Ano90b]. Reviews [Ano91b, Ano91c, Ano94b, Ano90c, Ano92a, Ano92b, Ano93, Ano94a, Ano95c, Ano96]. rhabdovirus [Col95, LLN94]. rhabdoviruses [LF94b, LLN94]. Rhinoptera [Tho90]. Rhizocephala [KDC95]. rhodei [AS93a]. Rhodococcus [SBMB95]. rhodurus [TMMA96b, TMMA96a]. Ribeiroia [HN97]. ribosomal [HYHA98]. ribosomes [CDPS90]. Richardson [NLS^{+}90, PR90, PMD90, SJMPP91, SBN^{+}91, SK90, WCS90]. rickettsia [CTC^{+}94, CNS91, RD93, KDL95]. rickettsia-like [CTC^{+}94, CNS91, RD93, KDL95]. rickettsial [ASRT^{+}94]. Risk [JGOB95, VD98a, VD98b, VD99]. Risso [MB95b]. River [Vir92, LMH^{+}97]. rivulatus [SAAH90]. Rivulus [CF91]. RNA [HYHA98, LHS94, SOF^{+}92, SOS94]. RNA-RNA [LHS94]. roach [AS93a, AS93b]. rock [AL95, AAL95]. rodlet [IHKS90, IKU98]. Rohde [RQW92]. Role [NVBK93, LE91a, WCS90]. rosenbergii [ASN90, LTC98, TWC99]. rot [TRR96]. rotavirus [MSM^{+}90]. rotaviruses [DTS^{+}92]. Roubal [RQW92]. Roughley [AL95]. roughleyi [AL95]. route [DRL94]. routes [CMAM96, PERL96]. RPS [LLN94]. rRNA [MHM95]. RT [LLGJS94]. ruckeri [DGK99, RMP^{+}94, Dav91b, Dav91a, F.92, JHL^{+}91, Rod91, VRK92]. rupestris [BBS96, CA95, GSS98]. rutili [Tar90]. Rutilus [AS93a, AS93b]. RV [TIM^{+}96]. RV-PJ [TIM^{+}96].

S.E [SQTF90]. sablefish [BHB90]. sac [BHT92, BB96a, BB96b, SD97b]. Saccharomyces [RRER90]. Saccostrea [AL95, AAL95]. saimon [SM96]. salar [AFJ^{+}97, ABJE93, BK96, BHG91, BSHJ91, BARH90, BSWR93, Bru92, BR92, BMO^{+}98, CGøL^{+}95, CFHR98, CNH^{+}97, CN99, CSP98, DIHS94, DS95, DLB95, DBIS96, DFK93, DF94, DFS94, DRP92, EGTH92, FHMG95, FM98, GG99, GHB91, GG97, GJS^{+}97, HEH^{+}90, HNWE94, HCM92, IFMR91, IRV^{+}91, IYBF93, JS95, JSB90, JBWT92, JLR93, KSPO99, LTGC99, LSR90, LLP92, LJHE91, MHB94, MRD98, MF95, Møy93, MLB95, NLTN92, NVSB94, NR97, NRV^{+}98, NC99, NHWE95, OCC^{+}95, OGFL92, PSB^{+}92, PMS98, PRL95, PERL96, RT94, Rit97, RRER90, RD93, RR98, RDMW98, SN91, SNØ^{+}93, SRRN94, SNó95, SBD^{+}92, SM96, SBMB95, SAMP97, SHT96, SABòJ99, SHB98, SCBB93, SSS^{+}99, SOP96, SDB99, TRSD98, TT99, Tat91, TL94, TRR96, VD98a, VD98b, VD99, WGSH94, WBS^{+}96, vdIOK96]. salaris [SHB98]. salinity [PFO^{+}98, SPC90]. Salmo [AFJ^{+}97, ABJE93, BK96, BHG91, BSHJ91, BARH90, BSWR93, Bru92, BR92, BMEC97, BMO^{+}98, CGøL^{+}95, CFHR98, CNH^{+}97, CN99, CSP98, DIHS94, DIB^{+}95, DS95, DLB95, DBIS96, DFK93, DF94, DFS94, DBCA97, DRP92, EGTH92, FHMG95, FM98, GG99, GHB91, GG97, GJS^{+}97, HEH^{+}90, HNWE94, HCM92, IFMR91, IRV^{+}91, IYBF93, JGOB95, JS95, JSB90, JBWT92, JLR93, KSPO99, LTGC99, LSR90, LLP92, LJHE91, MHB94, MRD98, MSW94, MF95, Møy93, MLB95, NLS^{+}90, NLTN92, NVSB94, NR97, NRV^{+}98, NC99, NHWE95, OCC^{+}95, OGFL92, PSB^{+}92, PR90, PMS98, PRL95, PERL96, RT94, Rit97, RRER90, RD93, RTEC94, RR98, RGA^{+}98, RDMW98, SN91, SNØ^{+}93, SRRN94, SNó95, SBD^{+}92, SMBM95, SM96, SBMB95, SAMP97, SK90, SHT96, SABòJ99, SHB98, SCBB93, SSS^{+}99, SOP96, SDB99, TRSD98]. Salmo [TT99, Tat91, TL94, TRR96, VD98a, VD98b, VD99, WGSH94, WBS^{+}96, ZPVH99, vdIOK96]. salmoides [Che92b]. salmon [AFJ^{+}97, ABJE93, BFPW90, BK96, BHG91, BSHJ91, BDM91, BARH90, BSWR93, Bru92, BR92, BMO^{+}98, CGøL^{+}95, CFHR98, CNH^{+}97, CN99, CSP98, DIHS94, DIB^{+}95, DS95, DLB95, DBIS96, DFK93, DF94, DFS94, DRP92, EGTH92, FHMG95, FM98, GCS95, GG99, GHB91, GG97, GJS^{+}97, HEH^{+}90, HK98, HNWE94, HCM92, IFMR91, IRV^{+}91, IYBF93, JGOB95, JS95, JSB90, JSW92, JBWT92, JLR93, KSPO99, LTGC99, LSR90, LLP92, LJHE91, MHB94, MSWB98, MRD98, MSW94, MF95, Møy93, MLB95, MRD^{+}92, NLTN92, NVSB94, NR97, NRV^{+}98, NC99, NHWE95, NKKH97, OCC^{+}95, OGFL92, PSB^{+}92, PMS98, PRL95, PERL96, PLLT94, RT94, Rit97, RRER90, RD93, RR98, RDMW98, SAK91, SN91, SNØ^{+}93, SRRN94, SNó95, SKA99, SBD^{+}92, SMBM95, SM96, SBMB95, SAMP97, SHT96, SABòJ99, SHB98, SCBB93, SSS^{+}99]. salmon [SOP96, SDB99, TOK^{+}94, TRSD98, TT99, Tat91, TMMA96b, TMMA96a, TL94, TRR96, VD98a, VD98b, VD99, WGSH94, WBS^{+}96, vdIOK96]. salmonae [BSBD99, SKA99, SBJ^{+}98, SDM^{+}98, SBD99]. salmonae-associated [SDM^{+}98]. salmonicida [AEJ95, AE99, ASR^{+}95, BLHA92, DWV94, Ell91, EGTH92, GJS^{+}97, NHOJ93, Ner96, NDBL94, REM90, SKQ^{+}99, ABJE93, AEJ95, AE99, Aus93, BSH^{+}90, BAHL91, BLHA92, BHA94, BWC92, Ber90, Ber91, BENJ92, BSHJ91, BHP^{+}94, BBS96, CDPS90, CCD95, DMH^{+}98, DSA^{+}98, DES92, DPPE96, DPPE97, DB96, EA94, EBL^{+}94, Ell91, FH91, HEH^{+}90, HMH96, HHE94, HWGV91, HCM92, HLJ^{+}91, IFMR91, IR91, IYBF93, IMR93, IRM^{+}96, KSS91, KON90, LWHD97, LTGC99, LE91a, LH90, LSR90, LJHE91, MGE90, MF91, Møy93, MLB95, NHOJ93, Ner96, NDBL94, NRV^{+}98, PRCC96, PBC^{+}90, RT94, RV91, REM90, RSTE93, SNØ91, SNA^{+}92, SNØ^{+}93, SH92, SD93, SABòJ99, SDB99, Tat91, TGK94, TL94, WLHD97, WDCM92, WDV93, Wik90]. salmonid [AFW94, BSB91, CMMM95, DGR97, EC91, GBH93, HIMK93, LE91b, MO95, RSTE93, SRL^{+}96, SBMB95, WR92]. salmonids [Aus93, BHB90, BHS98, CNS91, FJS90, KON90, OMLF95, OMF97, OBL^{+}99, PSL^{+}93, SD97a, SKQ^{+}99, Tar90, dMMMK96]. salmoninarum [BHB90, BHG91, DGR97, GSC^{+}99, GL91, GHB91, GBH93, JHL^{+}91, OCC^{+}95, SAK91, SA95]. salmonis [AFJ^{+}97, BSWR93, EMM90, GCS95, JSB90, JSW92, JBWT92, LF94a, MSWB98, Rit97, SSS^{+}99]. salmositica [AFW94, FW96, FW98, KBW95, SBW94, TW90a, TW90b, TW92, VW96, WT92, ZFW97, ZW98]. Salvelinus [AFW94, CC91, KRL96, MO95, RM92, SDS90, SDS95, SA94, SY94, SEMB^{+}99, SA95, SHB98]. samples [GBH93, TT99]. sand [LPO^{+}99]. sand-eel [LPO^{+}99]. sandwich [GBH93]. Sanguinicola [RHL^{+}94, SI91]. Sanguinicolidae [RHL^{+}94, SI91]. Saprolegnia [BQLC97, RM92, WR92]. Saprolegniaceae [RM92]. saprolegniosis [BLSC93]. sarafloxacin [MHB94]. Sarcomastigophora [TW90b, TW92, WT92]. Sauvage [LPO^{+}99]. saxatilis [TNB^{+}93]. scalare [RKMF97]. scallops [HOF96]. scanning [SHB95, TRR96]. Schlegel [CLK^{+}97, GMHK93, HHKE90, HIH^{+}96, HCW93, HLTK96, HLTK97a, HLTK97b, IKU^{+}96, KEY91, LKM99, YI90]. Schneider [AMN^{+}93, BBW98, DHM97, ERSA98, PBP94, RQW92, WR98]. Sciaenops [CCC93, CD95]. scleritis [OMF97]. Scomber [Dav98, MD90]. Scombridae [Jon90]. scombrus [Dav98, MD90]. Scophthalmidae [BRAP99]. Scophthalmus [AHA92, ALVS94, ASR^{+}93, BRAP99, DAL90, DHB^{+}95, FNTF95, Gat91, LNFT95, LES^{+}93, LPO^{+}98, NDQT95, OJW^{+}96, OJW^{+}98, PFO^{+}98, SBN^{+}91, TNAF99]. Scopthalmus [DDGP^{+}96]. Scotia [CC91, VCC94]. Scotland [AGS^{+}92, IFMR91, IYBF93, LTGC99, MSW94, RTEC94, RDMW98, SMBM95]. Scottish [ASR^{+}93, BHA94]. screening [BR99]. Scylla [KDC95]. SDS [LWHD97]. SDS-polyacrylamide [LWHD97]. sea [APAFSP95, Ath92, AB90, BR99, BSWR93, CSBCV98, CNH^{+}97, DAL90, DFGG^{+}99, DAKD98, FTB90, LGSO97, LW97, REM90, RRS93, SE99, SBAP90, SbFSAP92, SSS^{+}99, TAD99, Vir92, DM90, DAKD98]. sea-caged [CNH^{+}97]. search [SMHS93]. seasonality [ED91]. seawater [BHA95]. secondary [HLHW98, MF91]. secreted [PBC^{+}90]. secreting [BARH90, DES92, WBA91]. secretion [CDPS90, ZW98]. sediments [HLJ^{+}91]. segment [HSK96, SOF^{+}92, SOS94]. Sehlegel [UUNY91]. sekoke [KN90a]. selected [EM93, KWJ90, SSM95]. selection [BRW92]. selective [F.92, GBH93]. selective-differential [F.92]. SEMBV [TIM^{+}96]. semi [CGøL^{+}95, DMTM95]. semi-defined [DMTM95]. semi-purified [CGøL^{+}95]. semicirculatus [LHS94]. Senegal [FTB90]. senegalensis [FTB90, ZBM^{+}99]. sensitive [KN90b, PI93]. sensitivities [BAR95]. sensitivity [CPLR95, JSW92]. separate [NHWE95]. septicaemia [Col95, DT93, SBB^{+}93, SK97]. septum [PMS98]. sequence [LTKC93, Ner96]. sequences [HYHA98, MHM95]. sequencing [MHM95]. sequential [DRL94, MRD^{+}92, SFB^{+}91a]. Serine [SNA^{+}92, ABJE93, BWC92, EBL^{+}94, SNØ^{+}93, WDCM92, WHC^{+}98]. Seriola [IKU^{+}96]. seriolicida [AKK96, IKU^{+}96]. Serological [LLN94, RM95b, MRD98]. serology [LF94b, WPRM94]. serotype [SBD^{+}92]. serotypes [Dav91b, LYD94]. serotyping [CND90, LAB92]. serovars [LPD94]. Serranidae [BB94]. serrata [KDC95]. Serratia [AB90, AS92b, BTL^{+}92, NLS^{+}90]. Serum [LYLS95, BB97, BR99, CMAM96, FW98, HEH^{+}90, Møy93, SAK91, SB94, WB94, WST96]. serum-free [WB94]. several [Che92a]. sex [DT96]. Sh [SMBM95]. shallow [BHS98]. shanny [AEL^{+}99]. shark [BGMC96, HB98, SA90]. sharks [BBW98, HB96]. sharpeyi [MBM96]. shasta [IGH92]. shellfish [BA91]. Shetland [SMBM95]. Short [DPPE97, DAKD98, ERSA98, Góm98, ISM^{+}98, VHCCGMOS98, EM93, PS93]. short-term [EM93, PS93]. shrimp [CPLR95, KKS^{+}98, LTBL94, LLN94, LTL96, LTC96, RVSK99, RK91, TLMM94, WTKC97]. shrimps [AAAR98, DVB^{+}98]. Siamese [LDTC93]. Siddall [CD98]. Siderophore [EA91, GRR92, PLLT94]. Siderophores [BA91, MB92, Dav91a]. Siganus [SAAH90]. significance [AS93b]. silicone [WFS^{+}95]. silver [FTN97]. similar [LR97, PGLH96]. similarity [MMJ^{+}97]. simplified [DB96]. Simultaneous [HLJ^{+}91]. single [SE99]. single-dose [SE99]. site [BWC92]. sites [HLJ^{+}91]. situ [KSPO99, BB96a, BB96b, LTKC95, MAR99, SSSH92, SMHS93, SDM^{+}98]. six [MC94]. size [DT96, SOS94, SDS90]. skeletal [HB96, PM91, SA90]. skin [AMNH99, AS92b, BHS98, BGMC96, BR92, CSM93, FOM^{+}95, LMTW95, MB98, PS93, SRRN94, Tho90, VD98b]. skull [AA91]. Sleeping [BL96]. Sleepy [CNN^{+}94]. small [HR95]. smear [VPSL93]. smelt [AAHY90, MSM^{+}90]. Smith [Lan91a, Lan91b, LT92]. smolt [JGOB95]. smolting [WBS^{+}96]. smolts [SBD^{+}92, SMBM95, SHT96, TRSD98]. snakehead [Che92b, CLC90, CRWP95, Sub93]. snapper [SW90]. snook [KBMGL90]. sobria [NE91]. sockeye [HK98]. sodium [FMY94]. solar [DIB^{+}95, JGOB95, MSW94, SMBM95]. sole [ZBM^{+}99]. Solea [ZBM^{+}99]. soluble [ARX95, GL91, vdIOK96]. somatic [MMB93]. Some [BAR95, CD98, SKK90, DSA^{+}98]. Somniosus [BBW98]. sonicates [CATR97]. source [CCD95]. sources [WHHL95]. South [HOF96, Vir92, BDM91, CGA93, DHB91, Jon90, LYLS95]. South-East [LYLS95]. southern [CTC^{+}94]. soy [Ber90]. soybean [BK96]. soybeans [vdIOK96]. sp [AS92a, BLH^{+}90, Ber95, Bro90, Bru92, BP91, CP96, DBCA97, FTB90, FHMG95, HHS93, JOW^{+}97, Lan91a, LDTC93, MBM96, NDQT95, RVA98, SBAP90, SBMB95, Sub93, UUNY91, WRC95, SBD^{+}92, SMBM95]. spadefish [BLP96]. Spain [APAFSP95, AEL^{+}99, RMP^{+}94]. Spanish [SBAP90]. Sparidae [FTB90, Rou94b, Rou94a, WR98]. Sparus [APAFSP95, Ath92, DFGG^{+}99, FTB90, SbFSAP92]. spathaceum [WCS90]. spawners [BR99]. SPDV [MRD98, RDMW98]. special [IHKS90, MCRF94]. species [CPLR95, CLK^{+}97, CMMM95, DSA^{+}98, FLT^{+}98, FDS92, FCC92, HSP97, HAC^{+}92, KT93b, LMBS92, LSAB99, LTBL94, MEJ^{+}94, MCRF94, PGLH96]. Specific [HEH^{+}90, BR99, CYA^{+}98, CRWP95, DIB97, DES92, RRER90, WLHD97, WVG^{+}98, WCS90, YKI95]. specificities [BSHJ91]. specificity [PBC^{+}90]. Sphaerospora [DBCA97, MSW94, MEMSB99, SBAP90, SbFSAP92, dMMMK96]. Sphaerosporidae [MEMSB99, SBAP90]. spinal [VD98a]. Spindle [MS91]. Spindle-cell [MS91]. spined [JK90]. spinosa [AMM95]. Spirurida [HB98]. spleen [DIB^{+}95, DISS96, PRL95, SABòJ99]. splendens [LDTC93]. Splenic [CD95]. spontaneous [PGLH96]. spore [BP91]. spores [YDOW97]. sporogonic [IGH92, MSW94]. sporogony [AMM95]. spot [FCC92, KKS^{+}98, RVA98, RVSK99, Vir92, WTKC97]. spotted [CNN^{+}94]. spp [AA91, BHS98, BA91, CATR97, CAR97, CYA^{+}98, EA91, GRR92, HHS93, KSPO99, dMMMK96]. spring [RPT^{+}93]. Squalus [FDS92]. squamous [FCS91, MB94]. ssp [AE99, Ner96, AEJ95, Ell91, GJS^{+}97, NHOJ93, ZBM^{+}99]. St [LMH^{+}97]. stage [GCS95, MEMSB99]. stages [BB94, Ber95, Dav90, DM90, IGH92, JSB90, JBWT92, MSWB98, MSW94, Tho90]. staining [YDOW97, dMMMK96]. starved [DPPE96]. state [MF91, SDS95]. static‘green [ASN90]. status [CGøL^{+}95, DT96, MCRF94, SRL^{+}96, WBS^{+}96]. steatosis [STP^{+}98]. Steindachner [LLLS95]. Steinsfjorden [SQTF90]. sterile [SEMB^{+}99]. steroid [KRL96]. stickleback [JK90, LS93]. Stimpson [LTBL94]. stimulated [EC91]. stimulation [SJGB98, ZW98]. Stirling [Bro91]. stonefish [PBP94]. strain [ASR^{+}95, FH91, Gat91, HCW93, JOW^{+}97, MC94, NG99, PLLT94]. strains [AKK96, BAR95, Ber90, CCH^{+}94, DHTC99, ERJ91, KSS91, KRSE93, LAB92, LYLS95, LSR90, LSAB99, MEJ^{+}94, MC94, MF91, NBWO91, RV91, RFN93, RM95b, WDCM92, WT92]. strategic [WR92]. Streptococcosis [DDGP^{+}96, AMW96]. Streptococcus [BLH^{+}90, CP96, DMS99, DDGP^{+}96]. stress [AMNH99, CSP98, Kak98, SD93, TRSD98]. stress-inducible [SD93]. striatus [CLC90, CRWP95, Sub93]. striped [AMN^{+}93, TNB^{+}93]. structural [LF94b, SBB^{+}93]. structure [KR95]. structures [AO94]. Studies [GWFE97, TMMA96a, BBS96, DLB95, Dav90, DBCA97, ERJ91, F.92, FM98, HN94, JK90, LR97, LLN94, MHB94, SLJ97, SSM95, SHB98, WT92, dMMMK96]. Study [Che92b, MF95, ACH93, AKK96, AS93a, BS98, CRG^{+}97, CFS93, CW91, DMOR97, DRL94, FAAE^{+}99, GACHF91, HIH^{+}96, RR98, SE99, SQTF90, SK90, TWC99, VFCV99]. Stuhlmann [AO94]. stylirostris [LTBL94]. sub [SMBM95]. sub-type [SMBM95]. subcutaneous [DWV94]. sublethal [DWV94, HAH92]. subpopulations [RT94]. subsp [EGTH92, BLHA92, DWV94, REM90]. subsp. [NDBL94]. substances [JOW^{+}97]. substrate [BWC92]. Successful [SCBB93, PSM^{+}96]. such [SH92]. suecica [ABS92]. suffering [KRSE93]. sulphachlorpyridazine [OSG95]. sulphadiazine [OSG95]. sulphate [MLB95, Tho90]. sulphated [DS95]. sulphonamides [SLJ97]. Sumatra [TLMM94]. sunfish [FCS91]. superoxide [KSS91]. supplementation [CGøL^{+}95]. supply [FMDC93]. suppression [SDM^{+}98]. suratensis [RVA98]. surface [BSH^{+}90, DPPE97, DSB90, FW96, SABòJ99]. surfaces [LS93]. survey [MRD98, TLMM94]. Survival [EA94, GJS^{+}97, OJW^{+}98, DGR97, LF94a, PFO^{+}98, REM90]. susceptibilities [DT93]. Susceptibility [KSS91, LE91b, Che92a, IR91, LB98, Söv92]. susceptible [CHL99]. suspected [SPC90]. suttoni [KDL95]. SVCV [RPT^{+}93]. swamp [RM95a]. Swedish [HAH92]. swim [VC93]. swim-bladder [VC93]. swimbladder [NVBK93]. swimming [KBW95]. Sydney [AAL95]. sydneyi [AAL95]. Symposium [Bro91]. Synanceja [PBP94]. Synbranchidae [RM95a]. Synbranchus [RM95a]. Syndinidae [HHS93]. syndrome [AR95, AS91, AS92a, BMO^{+}98, CCH^{+}94, CRWP95, CNH^{+}97, KKS^{+}98, LF94b, LR97, Mai92, RVSK99, RWC93, Sub93, TOK^{+}94, WR94, WRC95, YIA92]. syns [CD98]. system [BEGA93, DIB97, LSL91, SNØ91]. Systematic [CTC^{+}94]. Systemic [KKP92, MOBF98, RKMF97, TIM^{+}96]. systems [APAFSP95, ASN90, APR99, PSB^{+}92].

T [SSJ97]. Taiwan [CTC^{+}94, LTC98, TWC99, WTKC97]. tapeworm [HN94]. tarda [GMHK93]. Tasmania [CNH^{+}97, CN99, HSP97, NC99]. Taurulus [Dav90]. tauvina [CNN^{+}94]. technique [DLP^{+}94, JHL^{+}91, MB94, YDOW97]. techniques [dMMMK96]. tegumental [LS93]. teleost [DHM97, FMDC93, FTB90]. teleostean [CA95]. Teleostei [BRAP99, Dav90, LMTW95, LPO^{+}99]. teleosts [STP^{+}98]. Temminck [CLK^{+}97, GMHK93, HHKE90, HIH^{+}96, HCW93, HLTK96, HLTK97a, HLTK97b, IKU^{+}96, KEY91, LKM99, YI90]. temmincki [WS90]. Temminek [UUNY91]. temperature [BSBD99, BHS98, CSBCV98, DHTC99, PFO^{+}98, SKQ^{+}99, SBD99]. temperatures [CRWP95]. ten [LMBS92]. tench [MB95a]. tension [KN90a]. tenuimanus [Lan91a, Lan91b, LT92]. term [EM93, PS93, YKI95]. test [AL95, ASRT^{+}94, TT99]. testicularis [SBAP90]. testing [LLP92, OCC^{+}95]. tetrachloride [AM92]. Tetramicra [LES^{+}93]. Tetraselmis [ABS92]. Thailand [MMJ^{+}97, PCKS94]. their [BSB91, GCS95, LF94b]. Thelohan [MSW94, Dav90]. thelohanellosis [YLSW99]. Thelohanellus [YLSW99]. Thelohania [SQTF90]. therapy [SCBB93]. thermal [AMNH99, HAH92, SBD99]. thermophilum [IKU^{+}96]. these [ISM^{+}98]. thick [MB95b]. thick-lipped [MB95b]. three [GHB91, HIMK93, JK90, KT93b, LLN94, NHWE95, WHHL95]. three-spined [JK90]. thrombosis [SNØ91]. thrombus [SNA^{+}92]. throughout [MMB93]. thymalli [DL99]. Thymallus [DL99]. thymectomized [FW98]. thymocyte [FW98]. thyroid [HIH^{+}96]. thyrsites [Lan91c, LTF92]. tiger [CPLR95, RRJ^{+}96, RK91, WTKC97]. tilapia [CP96, CTC^{+}94, CYA^{+}98, LLLS95, Mai92, MOBF98, SR99]. time [BHG91, PFO^{+}98]. Tinca [MB95a]. tip [SA90]. Tissue [DSA^{+}98, SDB99, WHHL95, DLB95, DHV^{+}99, DT93, KSPO99, PMD90, RPT^{+}93, SSJ97, TT99]. tissues [DFS94, JHL^{+}91, KN90b, NHWE95, YA95]. TNP [HK98]. TNP-470 [HK98]. toadfish [KGW90]. tolerance [IRV^{+}91, SPC90]. tomcod [LMH^{+}97, LMH^{+}97]. Toxicity [KON90, NLTN92]. toxin [Bro90, RFN93]. toxins [CMAM96, Ell91]. tract [JOW^{+}97]. traits [SGNG93]. transcription [LLGJS94]. transfer [Rit97]. transforming [KGW90]. transition [MB94]. transmission [AFJ^{+}97, DEA94, LT92, SKA99, SBMB95, Sub93]. transversum [PMS98]. Trapezia [HHS93]. treatment [AA91, DHB91, KGW90, LB98, MSWB98, MSGA95, SSJ97, SKA99, SAMP97, SSS^{+}99, TNAF99]. treatments [TSL^{+}94]. Trematoda [RHL^{+}94]. trematodes [DMOR97]. trial [DHB^{+}95]. triamcinolone [HM90]. tributaries [BLH^{+}90]. tributyltin [ABC^{+}98]. trichlorfon [Tho90]. Trichodina [Sub93]. trimethoprim [OSG95]. Trisopterus [PM91]. tropical [WRC95]. trout [AOES99, ACH93, Ath92, AS91, AGS^{+}92, AS92a, AS92b, BSBD99, BB94, BBCT^{+}94, BS98, BLW95, BMEC97, CGA93, CH92, CC91, DES91, DES92, DT96, DBCA97, DOF99, DT93, EC91, FOM^{+}95, FAAEC99, FWM94, GCS95, IGH92, ISM^{+}98, JMK^{+}95, KR95, KBW95, LW97, LB98, LOF96, MB92, MO95, MSGA95, MF91, NLS^{+}90, NKKH97, OLM^{+}97, PR90, PMD90, RTEC94, RGA^{+}98, RMR95, SJMPP91, SOAK93, SYAK95, SSJ97, SBN^{+}91, SB94, SDS90, SDS95, SA94, SY94, SEMB^{+}99, SBW94, SAMP97, SBJ^{+}98, SDM^{+}98, SK90, SA95, SHB98, TW90a, TW90b, TW92, TSL^{+}94, TNB^{+}93, VW96, WFS^{+}95, WVG^{+}98, WPRM94, WCS90, WST96, YA95, ZPVH99]. trutta [ZPVH99, BMEC97, DBCA97, RTEC94, RGA^{+}98]. Trypanoplasma [SKK90]. Trypanosoma [BB97, BB99, DMTM95, DTMD99, WB94, ZPVH97]. trypanosome [DMTM95, DTMD99]. trypsin [BARH90, NK93]. tryptone [Ber90]. tryptone-soy-Coomassieagar [Ber90]. tshawytscha [BFPW90, SKA99]. tuberculosis [Góm98]. tumour [MS91, SW90]. Tumours [MW94]. tuna [Jon90]. turbot [AHA92, ALVS94, ASR^{+}93, BRAP99, DAL90, DDGP^{+}96, DHB^{+}95, FNTF95, Gat91, LNFT95, LES^{+}93, LPO^{+}98, NDQT95, OJW^{+}96, OJW^{+}98, PFO^{+}98, SBN^{+}91, TNAF99]. twitch [KN90a]. Two [TAD99, AE99, CLK^{+}97, ERSA98, LYD94, LTBL94, RMP^{+}94, SBN^{+}91, SRL^{+}96, WLHD97]. type [LTKC93, LTKC95, SMBM95, VPSL93]. types [Dav91b]. typical [GG97, RV91, WDCM92]. typing [DMS99]. tyrannus [SPC90].

ulcer [BGMC96, GG99]. ulcerated [Aus93]. ulcerations [MB98]. ulcerative [AR95, CCH^{+}94, CRWP95, FCC92, LF94b, LR97, RWC93, Sub93, Vir92, WR94, WRC95, YIA92]. ulcers [LMH^{+}97, SRRN94, Wik90]. ultrasound [PMS98]. Ultrastructural [AM92, AMM95, Dav90, HN94, RHL^{+}94, CA95, DM90, IKU98, VFCV99]. Ultrastructure [DTMD99, FTB90, GWFE97, MMB93, SFB^{+}91b]. ultraviolet [BR92]. ultraviolet-B [BR92]. unassociated [SS90]. undecimalis [KBMGL90]. underyearlings [HK98]. undescribed [DM90]. unidentified [Dav98]. unit [SBD99]. University [Bro91]. Uptake [PMD90, DSA^{+}98]. urophthalmus [SASG94, SMG96, CR91]. uruguayanae [RMRR96]. usage [Rod91]. Use [DHS94, WBA91, GCO91, MC94, OGFL92, VW96, WR92]. used [Che92a, LTGC99, MB94, MLB95, TL94]. using [CCH^{+}94, CMMM95, FAAE^{+}99, GWG99, GBH93, KN90b, LWHD97, LHS94, MCRF94, PMS98, SBN^{+}91, SCBB93, dMMMK96]. Utilization [HHE94].

V [NK93]. V. [GCO91]. vaccinated [GJS^{+}97]. Vaccination [ERJ91, LW97, PV93, ARX95, DAL90, FAAE^{+}99, IRM^{+}96, Rod91, SEMB^{+}99]. Vaccine [GG99, CSP98, KRK91, KEY91, LW97, PERL96, SBW94]. vaccines [BCD90, GG97, LLP92, NR97, PRL95, SH92, SBN^{+}91, Tat91, TGK94]. Valenciennes [Cai92, HLTS93, WS90]. validating [SSM95]. vannamei [LTBL94, MA99]. variabilis [Dav90]. variant [AKK96]. variation [DGR97]. various [CYA^{+}98, Tho90]. vascular [BB94, IKU98]. Vavraia [Lan91a, LT92]. Venezuela [AAAR98]. verrucosa [PBP94]. versus [BCD90]. vetoquinol [SE99]. VHSV [FAAEC99, FAC99]. via [BMEC97, DHB^{+}95]. viability [YDOW97]. Vibrio [AAAR98, ASR^{+}93, ASR^{+}95, BSH^{+}90, BHS98, Ber95, BGMC96, BAE^{+}91, BA91, BEGA93, BSHJ91, Bro90, CPLR95, CH92, DMH^{+}98, DHB91, ERJ91, FLT^{+}98, FNTF95, GRR92, GG99, GCO91, HEH^{+}90, HAC^{+}92, HLJ^{+}91, KEY91, KRSE93, KSPO99, KBMGL90, LPD94, LSR90, MB92, MEJ^{+}94, MA99, NRV^{+}98, OJW^{+}96, OJW^{+}98, PLLT94, RHC^{+}93, RK91, SJH^{+}93]. vibriosis [DAL90, KRSE93, LPD94, NRV^{+}98, SYAK95, SBN^{+}91]. viraemia [RPT^{+}93]. Viral [DT93, LGSO97, YI90, AMN^{+}93, CLK^{+}97, CFS93, GHdB90, LNFT95, SSSH92, SMHS93, SBB^{+}93, SRL^{+}96]. virginica [ABC^{+}98]. Virulence [OMLF95, SOF^{+}92, Ber90, CT91, DGR97, DHTC99, LYLS95, MO95, SOS94]. virulent [CCH^{+}94, DWV94, DSB90, RM95b, WT92]. Virus [LDCA92, BB96a, BB96b, CHL99, CND90, CFHR98, CSP98, DLP^{+}94, DHS94, DT93, DRL94, EHH91, GSS98, GWG99, HEH^{+}90, HLHW98, HNWE94, HCW93, HEHR91, JS95, LAB92, LLGJS94, LTBL94, LTL96, MB97, MRD98, MO95, MC94, MF95, NS96, NHWE95, NKKH97, PSL^{+}93, PS98, RVSK99, RPT^{+}93, RM95b, RDMW98, SOF^{+}92, SOS94, SBB^{+}93, SDS90, SDS95, SA94, SY94, SS90, SBD^{+}92, SM96, TIM^{+}96, TRSD98, TT99, THBL96, TAD99, TWC99, WPRM94, WHC^{+}98, ZH96]. Virus-like [LDCA92]. viruses [EC91, GWFE97]. viscosus [GG99]. vital [YDOW97]. vitamin [CR91, CCC93, FWM94, TMMA96b, TMMA96a]. vitamin-C-deficient [CCC93]. vitamin-D-free [TMMA96b]. vitamins [FWM94, WVG^{+}98]. vitro [FLT^{+}98, Che92a, CW93, DS95, DF94, FW98, IR91, LYGS97, LH90, MB92, NK93, OSG95, PFO^{+}98, RT94, SNA^{+}92, TNAF99, TW90b, WT92, YIM^{+}97, ZFW97, ZW98]. vivo [FW98, NDBL94, RMT^{+}92, TNAF99, TW90b]. VL4355 [PLLT94]. VNN [CLK^{+}97, LGSO97]. VP2 [HSK96]. VP2/NS [HSK96]. VR [HCW93]. VR-299 [HCW93]. vulnificus [BAE^{+}91, BEGA93].

Walbaum [AOES99, ACH93, Ath92, AS91, AGS^{+}92, AS92a, AS92b, BFPW90, BSBD99, BBCT^{+}94, BDM91, BS98, BLW95, BMEC97, CGA93, CH92, DES91, DES92, DT96, DOF99, DT93, DRL94, FW98, FOM^{+}95, FWM94, GCS95, HK98, IGH92, ISM^{+}98, JMK^{+}95, KR95, KBW95, LMH^{+}97, LW97, LB98, LOF96, MB92, MSGA95, NKKH97, OLM^{+}97, RMR95, SAK91, SOAK93, SYAK95, SSJ97, SB94, SA94, SY94, SKA99, SEMB^{+}99, SBW94, SAMP97, SBJ^{+}98, SDM^{+}98, TOK^{+}94, TW90a, TW90b, TW92, TSL^{+}94, TNB^{+}93, VW96, WFS^{+}95, WVG^{+}98, WPRM94, WST96, YA95]. Wales [Vir92, HOF96]. walls [RRER90]. water [ASN90, AB90, DHTC99, LW97, MGE90, NRV^{+}98, NBWO91, REM90, SE99, SKQ^{+}99, SBD99]. waters [BRB^{+}94]. Watson [SASG94]. west [AEL^{+}99]. western [ZPVH97, ZPVH99, OGFL92]. which [Aus93, BWC92, BA91]. white [BP91, KKS^{+}98, RVSK99, WTKC97]. whole [CATR97]. whole-cell [CATR97]. wild [BFCH91, SBAP90]. Winter [SRRN94, AL95, BLSC93, DFGG^{+}99, GG99]. without [CGøL^{+}95]. wolffish [HMH96]. work [SNA^{+}92]. wound [AO94, BR92]. wound-inflicting [AO94]. wrasse [BBS96, GSS98, LTGC99, TL94]. WSBV [KKS^{+}98]. WSSV [RVSK99].

X [FMDC93, ARD92, DM90, SNØ91]. X-cell [FMDC93, DM90]. Xanthidae [HHS93]. xenobiotic [KGW90]. xenoma [BSBD99, SDM^{+}98]. xenoma-expression [BSBD99]. yabby [LT92]. Yamaguti [HN94]. YBV [LTBL94]. year [RMP^{+}94, SQTF90]. yeast [JLR93]. yeasts [LTC98]. yellow [AS91, LTBL94, LTL96]. yellow-head [LTL96]. yellow-pigmented [AS91]. yellowtail [IKU^{+}96, NS96]. Yersinia [DGK99, Dav91b, Dav91a, F.92, JHL^{+}91, Rod91, RMP^{+}94, VRK92]. yolk [BHT92, BB96a, BB96b, GMHK93, SD97b]. yolk-sac [BB96a, BB96b].

zarnowskii [JK90]. zebrafish [RMT^{+}92]. zinc [RMT^{+}92, SEMB^{+}99]. Zitnan [DL99]. Zn [ISM^{+}98].

References

Awa:1991:OTC

[AA91] J. N. Awa and W. O. Alegbeleye. Occurrence and treatment of cracked-skull disease affecting Clarias spp. in homestead ponds in Nigeria. Journal of Fish Diseases, 14(4):467–473, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alvarez:1998:VHP

[AAAR98] Alvarez, Austin, Alvarez, and Reyes. Vibrio harveyi: a pathogen of penaeid shrimps and fish in Venezuela. Journal of Fish Diseases, 21(4):313–316, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1990:IPL

[AAHY90] W. Ahne, K. Anders, M. Halder, and M. Yoshimizu. Isolation of picornavirus-like particles from the European smelt, Osmerus eperlanus (L.). Journal of Fish Diseases, 13(2):167– 168, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:1995:MDM

[AAL95] T. J. Anderson, R. D. Adlard, and R. J. G. Lester. Molecular diagnosis of Marteilia sydneyi (Paramyxea) in Sydney rock oysters, Saccostrea commercialis (Angas). Journal of Fish Diseases, 18(6):507–510, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1990:IFP

[AB90] B. Austin and A.-C. Billaud. Inhibition of the fish pathogen, Serratia liquefaciens, by an antibiotic-producing isolate of Planococcus recovered from sea water. Journal of Fish Diseases, 13(6):553–556, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:1998:ETH

[ABC^{+}98] R. S. Anderson, L. L. Brubacher, L. Ragone Calvo, M. A. Unger, and E. M. Burreson. Effects of tributyltin and hypoxia on the progression of Perkinsus marinus infections and host defence mechanisms in oyster, Crassostrea virginica (Gmelin). Journal of Fish Diseases, 21(5):371–380, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arnesen:1993:IRA

[ABJE93] J. A. Arnesen, R. Bjørnsdottir, T. Ø. Jørgensen, and G. Eggset. Immunological responses in Atlantic salmon, Salmo salar L., against purified serine protease and haemolysins from Aeromonas salmonicida. Journal of Fish Diseases, 16(5):409–423, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1992:IBF

[ABS92] B. Austin, E. Baudet, and M. Stobie. Inhibition of bacterial fish pathogens by Tetraselmis suecica. Journal of Fish Diseases, 15(1):55–61, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1993:MGP

[ACH93] D. J. Alderman and R. S. Clifton-Hadley. Malachite green: a pharmacokinetic study in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 16(4):297–311, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arnesen:1999:ICT

[AE99] J. A. Arnesen and G. Eggset. Isolation and characterisation of two extracellular metalloproteases from Aeromonas salmonicida ssp. salmonicida. Journal of Fish Diseases, 22(1):35–43, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arnesen:1995:PPC

[AEJ95] J. A. Arnesen, G. Eggset, and T. Ø. Jørgensen. Partial purification and characterization of extracellular metalloproteases from Aeromonas salmonicida ssp. salmonicida. Journal of Fish Diseases, 18(4):283–295, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arias:1999:DLD

[AEL^{+}99] C. Arias, J. Estévez, B. Lores, M. L. Sanmartín, and J. Leiro. Discovery of Loma dimorpha Loubés, Maurant, Gasc, De Buron & Barral 1984, in a new host, the shanny, Lipophrys pholis (L.), from the north-west coast of Spain. Journal of Fish Diseases, 22(2):155–159, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Almendras:1997:EIH

[AFJ^{+}97] F. E. Almendras, I. C. Fuentealba, S. R. M. Jones, F. Markham, and E. Spangler. Experimental infection and horizontal transmission of Piscirickettsia salmonis in freshwater-raised Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 20(6):409–418, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ardelli:1994:BCS

[AFW94] B. F. Ardelli, G. M. Forward, and P. T. K. Woo. Brook charr, Salvelinus fontinalis (Mitchill), and cryptobiosis: a potential salmonid reservoir host for Cryptobia salmositica Katz, 1951. Journal of Fish Diseases, 17(6):567–577, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1992:RJL

[AGS^{+}92] B. Austin, C. J. Gonzalez, M. Stobie, J. I. Curry, and M. F. Mcloughlin. Recovery of Janthinobacterium lividum from diseased rainbow trout, Oncorhynchus mykiss (Walbaum), in Northern Ireland and Scotland. Journal of Fish Diseases, 15 (4):357–359, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:1991:CCH

[AH91] T. Aoki and I. Hirono. Cloning and characterization of the haemolysin determinants from Aeromonas hydrophila. Journal of Fish Diseases, 14(3):303–312, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Al-Harbi:1992:IRT

[AHA92] A. H. Al-Harbi and B. Austin. The immune response of turbot, Scophthalmus maximus (L.), to lipopolysaccharide from a fish-pathogenic Cytophage-like bacterium. Journal of Fish Diseases, 15(5):449–452, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ansary:1992:PAR

[AHTY92] A. Ansary, R. M. Haneef, J. L. Torres, and M. Yadav. Plasmids and antibiotic resistance in Aeromonas hydrophila isolated in Malaysia from healthy and diseased fish. Journal of Fish Diseases, 15(2):191–196, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ainsworth:1993:CLI

[Ain93] A. J. Ainsworth. Carbohydrate and lectin interactions with Edwardsiella ictaluri and channel catfish, Ictalurus punctatus (Rafinesque), anterior kidney leucocytes and hepatocytes. Journal of Fish Diseases, 16(5):449–459, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alim:1996:CPS

[AKK96] S. R. Alim, K. Kawai, and R. Kusuda. Comparative pathogenicity study on antigenically variant strains of Enterococcus seriolicida. Journal of Fish Diseases, 19(1):39–46, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adlard:1995:DDT

[AL95] R. D. Adlard and R. J. G. Lester. Development of a diagnostic test for Mikrocytos roughleyi, the aetiological agent of Australian winter mortality of the commercial rock oyster, Saccostrea commercialis (Iredale & Roughley). Journal of Fish Diseases, 18(6):609–614, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Angulo:1994:HAM

[ALVS94] L. Angulo, J. E. Lopez, J. A. Vicente, and A. M. Saborido. Haemorrhagic areas in the mouth of farmed turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 17(2):163– 169, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Afifi:1992:UCA

[AM92] S. H. Afifi and J. R. Macmillan. Ultrastructural changes associated with carbon tetrachloride hepatotoxicity in channel catfish, Ictalurus punctatus Rafinesque. Journal of Fish Diseases, 15(2):119–129, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Azevedo:1995:UDS

[AMM95] C. Azevedo, E. Matos, and P. Matos. Ultrastructural data on sporogony of the coccidian parasite Calyptospora spinosa from the liver of the Amazonian fish, Crenicichla lepidota Heckel. Journal of Fish Diseases, 18(6):475–479, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arimoto:1993:PCA

[AMN^{+}93] M. Arimoto, K. Mori, T. Nakai, K. Muroga, and I. Furusawa. Pathogenicity of the causative agent of viral nervous necrosis disease in striped jack, Pseudocaranx dentex (Bloch & Schneider). Journal of Fish Diseases, 16(5):461–469, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aranichi:1999:ETS

[AMNH99] F. Aranichi, N. Mano, M. Nakane, and H. Hirose. Effects of thermal stress on skin defence lysins of European eel, Anguilla anguilla L. Journal of Fish Diseases, 22(3):227–229, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Akhlaghi:1996:CPA

[AMW96] M. Akhlaghi, B. L. Munday, and R. J. Whittington. Comparison of passive and active immunization of fish against streptococcosis (enterococcosis). Journal of Fish Diseases, 19(3):251–258, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1990:A

[Ano90a] Anonymous. ANNOUNCEMENT. Journal of Fish Diseases, 13(6):563, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1990:BRb

[Ano90b] Anonymous. Book Review. Journal of Fish Diseases, 13 (6):561, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1990:BRa

[Ano90c] Anonymous. Book reviews. Journal of Fish Diseases, 13(4): 333, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1991:A

[Ano91a] Anonymous. ANNOUNCEMENT. Journal of Fish Diseases, 14(6):697, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1991:BRa

[Ano91b] Anonymous. Book reviews. Journal of Fish Diseases, 14(1): 117–119, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1991:BRb

[Ano91c] Anonymous. Book reviews. Journal of Fish Diseases, 14(5): 603–604, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1992:BRa

[Ano92a] Anonymous. Book reviews. Journal of Fish Diseases, 15(1): 103, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1992:BRb

[Ano92b] Anonymous. Book reviews. Journal of Fish Diseases, 15 (6):545, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1992:E

[Ano92c] Anonymous. Erratum. Journal of Fish Diseases, 15(4):361, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1993:BR

[Ano93] Anonymous. Book reviews. Journal of Fish Diseases, 16 (6):613, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1994:BRa

[Ano94a] Anonymous. Book reviews. Journal of Fish Diseases, 17(2): 193, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1994:BRb

[Ano94b] Anonymous. Book reviews. Journal of Fish Diseases, 17 (5):555, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1995:BRb

[Ano95a] Anonymous. Book review. Journal of Fish Diseases, 18(4): 381, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1995:BRc

[Ano95b] Anonymous. Book review. Journal of Fish Diseases, 18(5): 473, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1995:BRa

[Ano95c] Anonymous. Book reviews. Journal of Fish Diseases, 18 (3):281–282, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1996:BR

[Ano96] Anonymous. Book reviews. Journal of Fish Diseases, 19 (3):259–260, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1998:BR

[Ano98] Anonymous. Book review. Journal of Fish Diseases, 21(2): 159, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1999:E

[Ano99] Anonymous. Erratum. Journal of Fish Diseases, 22(2):165, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avenant-Oldewage:1994:IDC

[AO94] A. Avenant-Oldewage. Integumental damage caused by Dolops ranarum (Stuhlmann, 1891) (Crustacea: Branchiura) to Clarias gariepinus (Burchell), with reference to normal histology and wound-inflicting structures. Journal of Fish Diseases, 17(6):641–647, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Afonso:1999:PAM

[AOES99] A. Afonso, C. Oliveira, A. E. Ellis, and M. T. Silva. Peroxidase activity as a measure of neutrophil populations in inflammatory peritoneal exudates of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 22(2):133–142, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alvarez-Pellitero:1995:PPG

[APAFSP95] P. Alvarez-Pellitero, A. Sitjà-Bob Adilla, A. Franco-Sierra, and O. Palenzuela. Protozoan parasites of gilthead sea bream, Sparus aurata L., from different culture systems in Spain. Journal of Fish Diseases, 18(2):105–115, March 1995. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou:1999:DPP

[APR99] F. Athanassopoulou, Th Prapas, and H. Rodger. Diseases of Puntazzo puntazzo Cuvier in marine aquaculture systems in Greece. Journal of Fish Diseases, 22(3):215–218, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:1995:FAO

[AR95] M. Ahmed and M. A. Rab. Factors affecting outbreaks of epizootic ulcerative syndrome in farmed fish in Bangladesh. Journal of Fish Diseases, 18(3):263–272, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:1992:DMA

[ARD92] A. Adams, R. H. Richards, and M. Marin De Mateo. Development of monoclonal antibodies to PK‘ X’, the causative agent of proliferative kidney disease. Journal of Fish Diseases, 15 (6):515–521, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ainsworth:1995:IRC

[ARX95] A. J. Ainsworth, C. D. Rice, and L. Xue. Immune responses of channel catfish, Ictalurus punctatus (Rafinesque), after oral or intraperitoneal vaccination with particulate or soluble Edwardsiella ictaluri antigen. Journal of Fish Diseases, 18(5): 397–409, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1991:RYP

[AS91] B. Austin and M. Stobie. Recovery of yellow-pigmented bacteria from dead and moribund fish during outbreaks of rainbow trout, Oncorhynchus mykiss (Walbaum), fry syndrome in England. Journal of Fish Diseases, 14(6):677–682, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1992:RML

[AS92a] B. Austin and M. Stobie. Recovery of Micrococcus luteus and presumptive Planococcus sp. from moribund fish during an outbreak of rainbow trout, Oncorhynchus mykiss (Walbaum), fry syndrome in England. Journal of Fish Diseases, 15(2): 203–206, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1992:RSP

[AS92b] B. Austin and M. Stobie. Recovery of Serratia plymuthica and presumptive Pseudomonas pseudoalcaligenes from skin lesions in rainbow trout, Oncorhynchus mykiss (Walbaum), otherwise infected with enteric redmouth. Journal of Fish Diseases, 15(6):541–543, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou:1993:CSM

[AS93a] F. Athanassopoulou and C. Sommerville. A comparative study of the myxosporeans Myxidium rhodei Léger, 1905 and Myxidium pfeifferi Auerbach, 1908 in roach, Rutilus rutilus L. Journal of Fish Diseases, 16(1):27–38, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou:1993:SMI

[AS93b] F. Athanassopoulou and C. Sommerville. The significance of myxosporean infections in roach, Rutilus rutilus L., in different habitats. Journal of Fish Diseases, 16(1):39–51, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:1990:MLM

[ASN90] I. G. Anderson, M. Shariff, and G. Nash. Mass larval mortalities in giant freshwater prawn, Macrobrachium rosenbergii de Man, cultured in Malaysian modified static‘green water’ systems. Journal of Fish Diseases, 13(2):127–134, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1993:VAC

[ASR^{+}93] B. Austin, M. Stobie, P. A. W. Robertson, H. G. Glass, J. R. Stark, and M. Mudarris. Vibrio alginolyticus: the cause of gill disease leading to progressive low-level mortalities among juvenile turbot, Scophthalmus maximus L., in a Scottish aquarium. Journal of Fish Diseases, 16(3):277–280, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1995:PSV

[ASR^{+}95] B. Austin, L. F. Stuckey, P. A. W. Robertson, I. Effendi, and D. R. W. Griffith. A probiotic strain of Vibrio alginolyticus effective in reducing diseases caused by Aeromonas salmonicida, Vibrio anguillarum and Vibrio ordalii. Journal of Fish Diseases, 18(1):93–96, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alday-Sanz:1994:IDT

[ASRT^{+}94] V. Alday-Sanz, H. Rodger, T. Turnbull, A. Adams, and R. H. Richards. An immunohistochemical diagnostic test for rickettsial disease. Journal of Fish Diseases, 17(2):189–191, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou:1992:ISB

[Ath92] F. Athanassopoulou. Ichthyophoniasis in sea bream, Sparus aurata (L.), and rainbow trout, Oncorhynchus mykiss (Walbaum), from Greece. Journal of Fish Diseases, 15(5):437–441, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1993:RAI

[Aus93] B. Austin. Recovery of ‘atypical’ isolates of Aeromonas salmonicida, which grow at 37^{o}C, from ulcerated nonsalmonids in England. Journal of Fish Diseases, 16(2):165– 168, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biosca:1991:SRO

[BA91] E. G. Biosca and C. Amaro. Siderophores and related outer membrane proteins in Vibrio spp. which are potential pathogens of fish and shellfish. Journal of Fish Diseases, 14 (3):249–263, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biosca:1991:FRV

[BAE^{+}91] E. G. Biosca, C. Amaro, C. Esteve, E. Alcaide, and E. Garay. First record of Vibrio vulnificus biotype 2 from diseased European eel, Anguilla anguilla L. Journal of Fish Diseases, 14(1):103–109, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barnes:1991:FDR

[BAHL91] A. C. Barnes, S. G. B. Amyes, T. S. Hastings, and C. S. Lewin. Fluoroquinolones display rapid bactericidal activity and low mutation frequencies against Aeromonas salmonicida. Journal of Fish Diseases, 14(6):661–667, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:1995:SBP

[BAR95] V. Bakopoulos, A. Adams, and R. H. Richards. Some biochemical properties and antibiotic sensitivities of Pasteurella piscidda isolated in Greece and comparison with strains from Japan, France and Italy. Journal of Fish Diseases, 18(1):1– 7, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:1997:EIL

[BAR97a] V. Bakopoulos, A. Adams, and R. H. Richards. The effect of iron limitation growth conditions on the cell and extracellular components of the fish pathogen Pasteurella piscicida. Journal of Fish Diseases, 20(4):297–305, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:1997:PCM

[BAR97b] V. Bakopoulos, A. Adams, and R. H. Richards. Production and characterization of monoclonal antibodies against Pasteurella piscicida, the causative agent of fish pasteurellosis. Journal of Fish Diseases, 20(4):307–315, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braun:1990:ILT

[BARH90] R. Braun, J. A. Arnesen, A. Rinne, and K. Hjelmeland. Immunohistological localization of trypsin in mucus-secreting cell layers of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 13(3):233–238, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bennett:1994:PAV

[BB94] S. M. Bennett and M. B. Bennett. Pathology of attachment and vascular damage associated with larval stages of Dissonus manteri Kabata, 1966 (Copepoda: Dissonidae) on the gills of coral trout, Plectropomus leopardus (Lacépède) (Serranidae). Journal of Fish Diseases, 17(5):447–459, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biering:1996:EIAa

[BB96a] E. Biering and Ø. Bergh. Experimental infection of Atlantic halibut, Hippoglossus hippoglossus L., yolk-sac larvae with infectious pancreatic necrosis virus: detection of virus by immunohistochemistry and in situ hybridization. Journal of Fish Diseases, 19(4):261–269, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biering:1996:EIAb

[BB96b] E. Biering and Ø. Bergh. Experimental infection of Atlantic halibut, Hippoglossus hippoglossus L., yolk-sac larvae with infectious pancreatic necrosis virus: detection of virus by immunohistochemistry and in situ hybridization. Journal of Fish Diseases, 19(6):405–413, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bienek:1997:CAG

[BB97] D. R. Bienek and M. Belosevic. Comparative assessment of growth of Trypanosoma danilewskyi (Laveran & Mesnil) in medium containing fish or mammalian serum. Journal of Fish Diseases, 20(3):217–221, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bienek:1999:MFC

[BB99] D. R. Bienek and M. Belosevic. Macrophage or fibroblastconditioned medium potentiates Growth of Trypanosoma danilewskyi Laveran & Mesnil 1904. Journal of Fish Diseases, 22(5):359–367, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bols:1994:DCL

[BBCT^{+}94] N. C. Bols, A. Barlian, M. Chirino-Trejo, S. J. Caldwell, P. Goegan, and L. E. J. Lee. Development of a cell line from primary cultures of rainbow trout, Oncorhynchus mykiss (Walbaum), gills. Journal of Fish Diseases, 17(6):601–611, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bricknell:1996:ASI

[BBS96] I. R. Bricknell, D. W. Bruno, and J. Stone. Aeromonas salmonicida infectivity studies in goldsinny wrasse, Ctenolabrus rupestris (L.). Journal of Fish Diseases, 19(6):469–474, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:1998:OLA

[BBW98] J. D. Borucinska, G. W. Benz, and H. E. Whiteley. Ocular lesions associated with attachment of the parasitic copepod Ommatokoita elongata (Grant) to corneas of Greenland sharks, Somniosus microcephalus (Bloch & Schneider). Journal of Fish Diseases, 21(6):415–422, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burkart:1990:ICC

[BCD90] M. A. Burkart, T. G. Clark, and H. W. Dickerson. Immunization of channel catfish, Ictalurus punctatus Rafinesque, against Ichthyophthirius multifiliis (Fouquet): killed versus live vaccines. Journal of Fish Diseases, 13(5):401–410, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Branson:1991:DND

[BDM91] E. J. Branson and D. Nieto Diaz-Munoz. Description of a new disease condition occurring in farmed Coho salmon, Oncorhynchus kisutch (Walbaum), in South America. Journal of Fish Diseases, 14(2):147–156, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biosca:1993:EAS

[BEGA93] E. G. Biosca, C. Esteve, E. Garay, and C. Amaro. Evaluation of the API 20E system for identification and discrimination of Vibrio vulnificus biotypes 1 and 2. Journal of Fish Diseases, 16(1):79–82, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjornsdottir:1992:LPA

[BENJ92] R. Bjørnsdottir, G. Eggset, R. Nilsen, and T. Ø. Jørgensen. The A-layer protein of Aeromonas salmonicida: further characterization and a new isolation procedure. Journal of Fish Diseases, 15(2):105–118, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bernoth:1990:AGT

[Ber90] E.-M. Bernoth. Autoagglutination, growth on tryptone-soy-Coomassieagar, outer membrane protein patterns and virulence of Aeromonas salmonicida strains. Journal of Fish Diseases, 13(2):145–155, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bernoth:1991:ICA

[Ber91] E.-M. Bernoth. Identification of cultured Aeromonas salmonicida by an indirect dot blot immunoassay. Journal of Fish Diseases, 14(3):419–422, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergh:1995:BAE

[Ber95] Ø. Bergh. Bacteria associated with early life stages of halibut, Hippoglossus hippoglossus L., inhibit growth of a pathogenic Vibrio sp. Journal of Fish Diseases, 18(1):31–40, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1993:HCL

[BF93] D. Bucke and S. W. Feist. Histopathological changes in the livers of dab, Limanda limanda (L.). Journal of Fish Diseases, 16(4):281–296, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1991:OPK

[BFCH91] D. Bucke, S. W. Feist, and R. S. Clifton-Hadley. The occurrence of proliferative kidney disease (PKD) in cultured and wild fish: further investigations. Journal of Fish Diseases, 14 (5):583–588, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Backman:1990:PPC

[BFPW90] S. Backman, H. W. Ferguson, J. F. Prescott, and B. P. Wilcock. Progressive panophthalmitis in Chinook salmon, Oncorhynchus tshawytscha (Walbaum): a case report. Journal of Fish Diseases, 13(5):345–353, September 1990. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bertone:1996:VCA

[BGMC96] S. Bertone, C. Gili, A. Moizo, and L. Calegari. Vibrio carchariae associated with a chronic skin ulcer on a shark, Carcharhinus plumbeus (Nardo). Journal of Fish Diseases, 19(6): 429–434, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barnes:1994:ARS

[BHA94] A. C. Barnes, T. S. Hastings, and S. G. B. Amyes. Amoxycillin resistance in Scottish isolates of Aeromonas salmonicida. Journal of Fish Diseases, 17(4):357–363, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barnes:1995:AAA

[BHA95] A. C. Barnes, T. S. Hastings, and S. G. B. Amyes. Aquaculture antibacterials are antagonized by seawater cations. Journal of Fish Diseases, 18(5):463–465, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bell:1990:PEI

[BHB90] G. R. Bell, R. W. Hoffmann, and L. L. Brown. Pathology of experimental infections of the sablefish, Anoplopoma fimbria (Pallas), with Renibacterium salmoninarum, the agent of bacterial kidney disease in salmonids. Journal of Fish Diseases, 13(5):355–367, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Benediktsdottir:1991:ITC

[BHG91] E. Benediktsdóttir, S. Helgason, and S. Gudmundsdóttir. Incubation time for the cultivation of Renibacterium salmoninarum from Atlantic salmon, Salmo salar L., broodfish. Journal of Fish Diseases, 14(1):97–102, January 1991. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boyd:1994:AGD

[BHP^{+}94] E. F. Boyd, M. P. Hiney, J. F. Pfdfn, P. R. Smith, and D. A. Caugant. Assessment of genetic diversity among Aeromonas salmonicida isolates by multilocus enzyme electrophoresis. Journal of Fish Diseases, 17(1):97–98, January 1994. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Benediktsdottir:1998:VSI

[BHS98] Benediktsdóttir, Helgason, and Sigurjónsdóttir. Vibrio spp. isolated from salmonids with shallow skin lesions and reared at low temperature. Journal of Fish Diseases, 21(1):19–28, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bergh:1992:EIE

[BHT92] Ø. Bergh, G. H. Hansen, and R. E. Taxt. Experimental infection of eggs and yolk sac larvae of halibut, Hippoglossus hippoglossus L. Journal of Fish Diseases, 15(5):379–391, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baeverfjord:1996:DRS

[BK96] G. Baeverfjord and A. Krogdahl. Development and regression of soybean meal induced enteritis in Atlantic salmon, Salmo salar L., distal intestine: a comparison with the intestines of fasted fish. Journal of Fish Diseases, 19(5):375–387, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boucher:1996:SDP

[BL96] P. Boucher and F. Baudin Laurencin. Sleeping disease and pancreas disease: comparative histopathology and acquired cross-protection. Journal of Fish Diseases, 19(4):303–310, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baya:1990:ASS

[BLH^{+}90] A. M. Baya, B. Lupiani, F. M. Hetrick, B. S. Roberson, R. Lukacovic, E. May, and C. Poukish. Association of Streptococcus sp. with fish mortalities in the Chesapeake Bay and its tributaries. Journal of Fish Diseases, 13(3):251–253, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barnes:1992:AOM

[BLHA92] A. C. Barnes, C. S. Lewin, T. S. Hastings, and S. G. B. Amyes. Alterations in outer membrane proteins identified in a clinical isolate of Aeromonas salmonicida subsp. salmonicida. Journal of Fish Diseases, 15(3):279–282, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beck:1996:OIL

[BLP96] K. Beck, G. Lewbart, and G. Piner. The occurrence of an Ichthyobodo-like organism on captive Atlantic spadefish, Chaetodipterus faber (Broussonet). Journal of Fish Diseases, 19(1):111–112, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bly:1993:EFA

[BLSC93] J. E. Bly, L. A. Lawson, A. J. Szalai, and L. W. Clem. Environmental factors affecting outbreaks of winter saprolegniosis in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 16(6):541–549, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bryant:1995:IAA

[BLW95] M. S. Bryant, R. J. G. Lester, and R. J. Whittington. Immunogenicity of amoebic antigens in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 18 (1):9–19, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bury:1997:LDB

[BMEC97] N. R. Bury, J. C. McGeer, F. B. Eddy, and G. A. Codd. Liver damage in brown trout, Salmo trutta L., and rainbow trout, Oncorhynchus mykiss (Walbaum), following administration of the cyanobacterial hepatotoxin microcystin-LR via the dorsal aorta. Journal of Fish Diseases, 20(3):209–215, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Byrne:1998:HKS

[BMO^{+}98] Byrne, MacPhee, Ostland, Johnson, and Ferguson. Haemorrhagic kidney syndrome of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21(2):81–91, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bunton:1991:KSM

[BP91] T. E. Bunton and S. L. Poynton. Kudoa sp. (Myxosporea, Multivalvulida) infection in juvenile white perch, Morone americana (Gmelin): histopathology and spore morphology. Journal of Fish Diseases, 14(5):589–594, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bly:1997:ISP

[BQLC97] J. E. Bly, S. M.-A. Quiniou, L. A. Lawson, and L. W. Clem. Inhibition of Saprolegnia pathogenic for fish by Pseudomonas fluorescens. Journal of Fish Diseases, 20(1):35–40, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1992:IUB

[BR92] A. M. Bullock and R. J. Roberts. The influence of ultraviolet-B radiation on the mechanism of wound repair in the skin of the Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(2):143–152, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Breuil:1999:REM

[BR99] G. Breuil and B. Romestand. A rapid ELISA method for detecting specific antibody level against nodavirus in the serum of the sea bass, Dicentrarchus labrax (L.): application to the screening of spawners in a sea bass hatchery. Journal of Fish Diseases, 22(1):45–52, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Branson:1999:MIC

[BRAP99] E. Branson, A. Riaza, and P. Alvarez-Pellitero. Myxosporean infection causing intestinal disease in farmed turbot, Scophthalmus maximus (L.), (Teleostei: Scophthalmidae). Journal of Fish Diseases, 22(5):395–399, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bogovski:1994:ICM

[BRB^{+}94] S. Bogovski, L. Rossi, V. Bocchini, S. Lastraioli, C. Aiello, and L. Santi. Immunohistochemical characterization of malignant lymphoma in northern pike, Esox lucius L., from Estonian Baltic waters. Journal of Fish Diseases, 17(6):557–566, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bricknell:1995:RMI

[Bri95] I. R. Bricknell. A reliable method for the induction of experimental furunculosis. Journal of Fish Diseases, 18(2):127–133, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brown:1990:NRE

[Bro90] C. Brown. Nutritional requirements for extracellular accumulation of toxin produced by a pathogenic Vibrio sp. Journal of Fish Diseases, 13(3):239–243, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bromage:1991:BDF

[Bro91] Niall Bromage. ‘Bacterial Diseases of Fish’: An International Symposium held at the University of Stirling, 26–29 June 1990. Journal of Fish Diseases, 14(3):247, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1992:ISF

[Bru92] D. W. Bruno. Ichthyobodo sp. on farmed Atlantic salmon, Salmo salar L., reared in the marine environment. Journal of Fish Diseases, 15(4):349–351, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:1992:ESM

[BRW92] K. Buchmann, A. Roepstorff, and P. J. Waller. Experimental selection of mebendazole-resistant gill monogeneans from the European eel, Anguilla anguilla L. Journal of Fish Diseases, 15(5):393–400, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1998:LFR

[BS98] D. W. Bruno and D. A. Smail. Lymphoma in farmed rainbow trout, Oncorhynchus mykiss (Walbaum): a case study. Journal of Fish Diseases, 21(5):381–386, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barker:1991:CBT

[BSB91] G. A. Barker, S. N. Smith, and N. R. Bromage. Commensal bacteria and their possible relationship to the mortahty of incubating salmonid eggs. Journal of Fish Diseases, 14(2): 199–210, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beaman:1999:EPA

[BSBD99] H. J. Beaman, D. J. Speare, M. Brimacombe, and J. Daley. Evaluating protection against Loma salmonae generated from primary exposure of rainbow trout, Oncorhynchus mykiss (Walbaum), outside of the xenoma-expression temperature boundaries. Journal of Fish Diseases, 22(6):445–450, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

BOGwald:1990:EIA

[BSH^{+}90] J. Bøgwald, K. Stensvåg, J. Hoffman, S. Espelid, K. O. Holm, and T. Jørgensen. Electrophoretic and immunochemical analysis of surface antigens of the fish pathogens Vibrio salmonicida and Vibrio anguillarum. Journal of Fish Diseases, 13 (4):293–301, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

BoGwald:1991:ASA

[BSHJ91] J. BøGwald, K. Stensvag, J. Hoffman, and T. Jørgensen. Antibody specificities in Atlantic salmon, Salmo salar L., against the fish pathogens Vibrio salmonicida and Vibrio anguillarum. Journal of Fish Diseases, 14(1):79–87, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bernet:1999:HFP

[BSM^{+}99] D. Bernet, H. Schmidt, W. Meier, P. Burkhardt-Holm, and T. Wahli. Histopathology in fish: proposal for a protocol to assess aquatic pollution. Journal of Fish Diseases, 22(1):25– 34, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bron:1993:FMA

[BSWR93] J. E. Bron, C. Sommerville, R. Wootten, and G. H. Rae. Fallowing of marine Atlantic salmon, Salmo salar L., farms as a method for the control of sea lice, Lepeophtheirus salmonis (Kroyer, 1837). Journal of Fish Diseases, 16(5):487–493, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baya:1992:SMP

[BTL^{+}92] A. M. Baya, A. E. Toranzo, B. Lupiani, Y. Santos, and F. M. Hetrick. Serratia marcescens: a potential pathogen for fish. Journal of Fish Diseases, 15(1):15–26, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bennett:1992:RAF

[BWC92] A. J. Bennett, P. W. Whitby, and G. Coleman. Retention of antigenicity by a fragment of Aeromonas salmonicida 70-kDa serine protease which includes the primary substrate binding site expressed as β-galactosidase hybrid proteins. Journal of Fish Diseases, 15(6):473–484, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1992:GAP

[CA92] D. Chen and A. J. Ainsworth. Glucan administration potentiates immune defence mechanisms of channel catfish, Ictalurus punctatus Rafinesque. Journal of Fish Diseases, 15(4):295– 304, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Casal:1995:NUD

[CA95] G. Casal and C. Azevedo. New ultrastructural data on the microsporidian Ichthyosporidium giganteum infecting the marine teleostean fish Ctenolabrus rupestris (L.). Journal of Fish Diseases, 18(2):191–194, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cai:1992:ICG

[Cai92] W. Cai. Isoenzymatic changes in grass carp, Ctenopharyngodon idellus Cuvier & Valenciennes, affected with haemorrhagic disease. Journal of Fish Diseases, 15(4):305–313, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1997:EPM

[CAR97] S.-C. Chen, A. Adams, and R. H. Richards. Extracellular products from Mycobacterium spp. in fish. Journal of Fish Diseases, 20(1):19–25, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1997:CAE

[CATR97] S.-C. Chen, A. Adams, K. D. Thompson, and R. H. Richards. A comparison of the antigenicity of the extracellular products and whole-cell sonicates from Mycobacterium spp. in rabbits, mice and fish by immunoblotting and enzyme-linked immunosorbent assay. Journal of Fish Diseases, 20(6):427–442, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cone:1991:DMP

[CC91] D. K. Cone and R. Cusack. Description of myxosporean presporogonic phases in the kidney of brook trout, Salvelinus fontinalis (Mitchill), in Nova Scotia, Canada. Journal of Fish Diseases, 14(4):503–505, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Collins:1993:RLL

[CCC93] B. K. Collins, L. L. Collier, and J. S. Collins. Retinal and lenticular lesions in vitamin-C-deficient juvenile red drum, Sciaenops ocellatus (L.). Journal of Fish Diseases, 16(3):229– 237, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Coleman:1995:GPE

[CCD95] G. Coleman, R. J. Collighan, and S. J. Dodsworth. Glycogen, the preferred energy source for extracellular protein formation and growth of Aeromonas salmonicida. Journal of Fish Diseases, 18(3):221–228, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cartwright:1994:IVS

[CCH^{+}94] G. A. Cartwright, D. Chen, P. J. Hanna, N. Gudkovs, and K. Tajima. Immunodiagnosis of virulent strains of Aeromonas hydrophila associated with epizootic ulcerative syndrome (EUS) using a monoclonal antibody. Journal of Fish Diseases, 17(2):123–133, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Colorni:1995:SCL

[CD95] A. Colorni and A. Diamant. Splenic and cardiac lymphocystis in the red drum, Sciaenops ocellatus (L.). Journal of Fish Diseases, 18(5):467–471, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cruz:1998:SOB

[CD98] C. Cruz and A. J. Davies. Some observations on Babesiosoma bettencourti (França, 1908) n. comb. (syns. Haemogregarina bettencourti França, 1908; Desseria bettencourti Siddall, 1995) from eels, Anguilla anguilla L., in Portugal. Journal of Fish Diseases, 21(6):443–448, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Constantino:1999:DFG

[CDGM99] F. Constantino, A. A. De Ocampo, and L. J. García-Márquez. Dermal fibroma in goldfish, Carassius auratus (L.). Journal of Fish Diseases, 22(3):223–226, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campbell:1990:SAP

[CDPS90] C. M. Campbell, D. Duncan, N. C. Price, and L. Stevens. The secretion of amylase, phospholipase and protease from Aeromonas salmonicida, and the correlation with membraneassociated ribosomes. Journal of Fish Diseases, 13(6):463– 474, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Courtney:1991:OCR

[CF91] L. A. Courtney and J. W. Fournie. Ocular chondrosarcomas in Rivulus marmoratus Poey. Journal of Fish Diseases, 14(1): 111–116, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Christie:1998:IPD

[CFHR98] K. E. Christie, K. Fyrand, L. Holtet, and H. M. Rowley. Isolation of pancreas disease virus from farmed Atlantic salmon, Salmo salar L., in Norway. Journal of Fish Diseases, 21(5): 391–394, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chou:1993:AMA

[CFS93] H.-Y. Chou, H. Fukuda, and T. Sano. Application of a monoclonal antibody against viral nucleoprotein to an aetiological study of infectious haematopoietic necrosis. Journal of Fish Diseases, 16(2):149–153, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carson:1993:CEL

[CGA93] J. Carson, N. Gudkovs, and B. Austin. Characteristics of an Enterococcus-like bacterium from Australia and South Africa, pathogenic for rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 16(4):381–388, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Christiansen:1995:ASI

[CGøL^{+}95] R. Christiansen, J. Glette, ø. Lie, O. J. Torrissen, and R. Waagbø. Antioxidant status and immunity in Atlantic salmon, Salmo salar L., fed semi-purified diets with and without astaxanthin supplementation. Journal of Fish Diseases, 18(4):317–328, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1992:AVP

[CH92] D. Chen and P. J. Hanna. Attachment of Vibrio pathogens to cells of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 15(4):331–337, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1992:VSN

[Che92a] S.-C. Chen. In vitro susceptibility of Nocardia asteroides to several commonly used disinfectants. Journal of Fish Diseases, 15(4):345–348, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1992:SPN

[Che92b] S.-C. Chen. Study on the pathogenicity of Nocardia asteroides to the Formosa snakehead, Channa maculata (Lacépède), and largemouth bass, Micropterus salmoides (Lacépède). Journal of Fish Diseases, 15(1):47–53, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chi:1999:ECC

[CHL99] S. C. Chi, W. W. Hu, and B. J. Lo. Establishment and characterization of a continuous cell line (GF-1) derived from grouper, Epinephelus coioides (Hamilton): a cell line susceptible to grouper nervous necrosis virus (GNNV). Journal of Fish Diseases, 22(3):173–182, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chinabut:1990:MSC

[CLC90] S. Chinabut, C. Limsuwan, and P. Chanratchakool. Mycobacteriosis in the snakehead, Channa striatus (Fowler). Journal of Fish Diseases, 13(6):531–535, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chi:1997:MMA

[CLK^{+}97] S. C. Chi, C. F. Lo, G. H. Kou, P. S. Chang, S. E. Peng, and S. N. Chen. Mass mortalities associated with viral nervous necrosis (VNN) disease in two species of hatchery-reared grouper, Epinephelus fuscogutatus and Epinephelus akaara (Temminck & Schlegel). Journal of Fish Diseases, 20(3):185– 193, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cross:1992:ICC

[CM92] M. L. Cross and R. A. Matthews. Ichthyophthiriasis in carp, Cyprinus carpio L.: fate of parasites in immunized fish. Journal of Fish Diseases, 15(6):497–505, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carbis:1996:EMS

[CMAM96] C. R. Carbis, G. F. Mitchell, J. W. Anderson, and I. Mc-Cauley. The effects of microcystins on the serum biochemistry of carp, Cyprinus carpio L., when the toxins are administered by gavage, immersion and intraperitoneal routes. Journal of Fish Diseases, 19(2):151–159, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cunningham:1995:IGM

[CMMM95] C. O. Cunningham, D. M. Mcgillivray, K. Mackenzie, and W. T. Melvin. Identification of Gyrodactylus (Monogenea) species parasitizing salmonid fish using DNA probes. Journal of Fish Diseases, 18(6):539–544, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clark:1999:FIA

[CN99] A. Clark and B. F. Nowak. Field investigations of amoebic gill disease in Atlantic salmon, Salmo salar L., in Tasmania. Journal of Fish Diseases, 22(6):433–443, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Christie:1990:IPN

[CND90] K. E. Christie, S. Ness, and H. O. Djupvik. Infectious pancreatic necrosis virus in Norway: partial serotyping by monoclonal antibodies. Journal of Fish Diseases, 13(4):323–327, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clark:1997:CNG

[CNH^{+}97] A. Clark, B. Nowak, J. Handlinger, B. L. Munday, and S. Percival. Clubbing and necrosis gill (CNG) syndrome in seacaged Atlantic salmon, Salmo salar L., in Tasmania: an initial report. Journal of Fish Diseases, 20(1):59–68, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chua:1994:ION

[CNN^{+}94] F. H. C. Chua, M. L. Ng, K. L. Ng, J. J. Loo, and J. Y. Wee. Investigation of outbreaks of a novel disease, ‘Sleepy Grouper Disease’, affecting the brown-spotted grouper, Epinephelus tauvina Forsskål. Journal of Fish Diseases, 17(4):417–427, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cvitanich:1991:IRL

[CNS91] J. D. Cvitanich, O. Garate N., and C. E. Smith. The isolation of a rickettsia-like organism causing disease and mortality in Chilean salmonids and its confirmation by Koch’s postulate. Journal of Fish Diseases, 14(2):121–145, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Coll:1995:LPI

[Col95] J. M. Coll. Low-pH increases the binding of haemorrhagic septicaemia rhabdovirus to membrane phospholipids. Journal of Fish Diseases, 18(6):519–527, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:1996:HES

[CP96] P. H. Chang and J. A. Plumb. Histopathology of experimental Streptococcus sp. infection in tilapia, Oroochromis niloticus (L.), and channel catfish, Ictafurus punctatus (Ratinesque). Journal of Fish Diseases, 19(3):235–241, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chanratchakool:1995:OSV

[CPLR95] P. Chanratchakool, M. Pearson, C. Limsuwan, and R. J. Roberts. Oxytetracycline sensitivity of Vibrio species isolated from diseased black tiger shrimp, Penaeus monodon Fabricius. Journal of Fish Diseases, 18(1):79–82, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

ChavezDeMartinez:1991:HVC

[CR91] M. C. Chávez De Martínez and R. H. Richards. Histopathology of vitamin C deficiency in a cichlid, Cichlasoma urophthalmus (Günther). Journal of Fish Diseases, 14(5):507–519, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carbis:1997:SFC

[CRG^{+}97] C. R. Carbis, G. T. Rawlin, P. Grant, G. F. Mitchell, J. W. Anderson, and I. McCauley. A study of feral carp, Cyprinus carpio L., exposed to Microcystis aeruginosa at Lake Mokoan, Australia, and possible implications for fish health. Journal of Fish Diseases, 20(2):81–91, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carbis:1996:HCC

[CRM^{+}96] C. R. Carbis, G. T. Rawlin, G. F. Mitchell, J. W. Anderson, and I. McCauley. The histopathology of carp, Cyprinus carpio L., exposed to microcystins by gavage, immersion and intraperitoneal administration. Journal of Fish Diseases, 19 (3):199–207, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chinabut:1995:HSC

[CRWP95] S. Chinabut, R. J. Roberts, G. R. Willoughby, and M. D. Pearson. Histopathology of snakehead, Channa striatus (Bloch), experimentally infected with the specific Aphanomyces fungus associated with epizootic ulcerative syndrome (EUS) at different temperatures. Journal of Fish Diseases, 18(1):41–47, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cecchini:1998:ITL

[CSBCV98] Cecchini, Saroglia, Berni, and Cognetti-Varriale. Influence of temperature on the life cycle of Diplectanum aequans (Monogenea, Diplectanidae), parasitic on sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 21(1):73–75, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Company:1999:BPP

[CSBP^{+}99] R. Company, A. Sitjà-Bobadilla, M. J. Pujalte, E. Garay, P. Alvarez-Pellitero, and J. Perez-Sanchez. Bacterial and parasitic pathogens in cultured common dentex, Dentex dentex L. Journal of Fish Diseases, 22(4):299–309, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carson:1993:CJP

[CSM93] J. Carson, L. M. Schmidtke, and B. L. Munday. Cytophaga johnsonae: a putative skin pathogen of juvenile farmed barramundi, Lates calcarifer Bloch. Journal of Fish Diseases, 16 (3):209–218, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Colquhoun:1998:PFI

[CSP98] D. J. Colquhoun, E. Skjerve, and T. T. Poppe. Pseudomonas fluorescens, infectious pancreatic necrosis virus and environmental stress as potential factors in the development of vaccine related adhesions in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21(5):355–364, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chabot:1991:PAH

[CT91] D. J. Chabot and R. L. Thune. Proteases of the Aeromonas hydrophila complex: identification, characterization and relation to virulence in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 14(2):171–183, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1994:SGC

[CTC^{+}94] S.-C. Chen, M.-C. Tung, S.-P. Chen, J.-F. Tsai, P.-C. Wang, R.-S. Chen, S.-C. Lin, and A. Adams. Systematic granulomas caused by a rickettsia-like organism in Nile tilapia, Oreochronuis niloticus (L.), from southern Taiwan. Journal of Fish Diseases, 17(6):591–599, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chowdhury:1991:SFC

[CW91] M. B. R. Chowdhury and H. Wakabayashi. A study on Flexibacter columnaris infection in loach, Misgurnus anguillicaudatus (Bleeker, Günther). Journal of Fish Diseases, 14 (3):389–394, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1993:VAA

[CW93] S.-C. Chen and P.-C. Wang. In vitro activity of antimicrobial agents against Nocardia asteroides. Journal of Fish Diseases, 16(3):269–272, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1998:NSI

[CYA^{+}98] Chen, Yoshida, Adams, Thompson, and Richards. Nonspecific immune response of Nile tilapia, Oreochromis nilotica, to the extracellular products of Mycobacterium spp. and to various adjuvants. Journal of Fish Diseases, 21(1):39–46, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Doukas:1998:SCA

[DAKD98] Doukas, Athanassopoulou, Karagouni, and Dotsika. Short communication Aeromonas hydrophila infection in cultured sea bass, Dicentrarchus labrax L., and Puntazzo puntazzo Cuvier from the Aegean Sea. Journal of Fish Diseases, 21 (4):317–320, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dec:1990:EOV

[DAL90] C. Dec, P. Angelidis, and F. Baudin Laurencin. Effects of oral vaccination against vibriosis in turbot, Scophthalmus maximus (L.), and sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 13(5):369–376, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1990:USE

[Dav90] A. J. Davies. Ultrastructural studies on the endogenous stages of Eimeria variabilis (Thélohan, 1893) Reichenow, 1921, from Coitus (Taurulus) bubalis Euphrasen (Teleostei: Cottidae). Journal of Fish Diseases, 13(6):447–461, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1991:YRP

[Dav91a] R. L. Davies. Yersinia ruckeri produces four iron-regulated outer membrane proteins but does not produce detectable siderophores. Journal of Fish Diseases, 14(5):563–570, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1991:CAY

[Dav91b] R. L. Davies. Clonal analysis of Yersinia ruckeri based on biotypes, serotypes and outer membrane protein-types. Journal of Fish Diseases, 14(2):221–228, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1998:UPP

[Dav98] Davies. An unidentified protozoan parasite from the blood of Atlantic mackerel, Scomber scombrus L. Journal of Fish Diseases, 21(2):145–147, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Djebara:1996:SAA

[DB96] M. Djebara and T. H. Birkbeck. A simplified assay for the activity of Aeromonas salmonicida glycerophospholipids cholesterol acyltransferase. Journal of Fish Diseases, 19(5):395– 397, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeMateo:1997:LHS

[DBCA97] M. De Mateo, G. Bovo, M. Comuzzi, and A. Adams. Lectin histochemical studies on Sphaerospora sp. (Myxosporea) from Italian brown trout, Salmo trutta L. Journal of Fish Diseases, 20(1):51–58, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1996:IEL

[DBIS96] R. A. Dalmo, J. Bogwald, K. Ingebrigtsen, and R. Seljelid. The immunomodulatory effect of laminaran [β (1,3)-D-glucan] on Atlantic salmon, Salmo salar L., anterior kidney leucocytes after intraperitoneal, peroral and peranal administration. Journal of Fish Diseases, 19(6):449–457, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Domeenech:1996:SCT

[DDGP^{+}96] A. Domeénech, J. F. Derenaáandez-Garayzábal, C. Pascual, J. A. Garcia, M. T. Cutuli, M. A. Moreno, M. D. Collins, and L. Dominguez. Streptococcosis in cultured turbot, Scopthalmus maximus (L.), associated with Streptococcus parauberis. Journal of Fish Diseases, 19(1):33–38, January 1996. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1994:ITH

[DEA94] A. J. Davies, J. C. Eiras, and R. T. E. Austin. Investigations into the transmission of Haemogregarina bigemina Laveran & Mesnil, 1901 (Apicomplexa: Adeleorina) between intertidal fishes in Portugal. Journal of Fish Diseases, 17(3):283–289, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davidson:1991:CRL

[DES91] G. A. Davidson, A. E. Ellis, and C. J. Secombes. Cellular responses of leucocytes isolated from the gut of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 14(6):651–659, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davidson:1992:EAQ

[DES92] G. A. Davidson, A. E. Ellis, and C. J. Secombes. An ELISPOT assay for the quantification of specific antibodysecreting cells to Aeromonas salmonicida in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 15(1):85–89, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dannevig:1994:ASS

[DF94] B. H. Dannevig and K. Falk. Atlantic salmon, Salmo salar L., develop infectious salmon anaemia (ISA) after inoculation with in vitro infected leucocytes. Journal of Fish Diseases, 17 (2):183–187, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Domenech:1999:APA

[DFGG^{+}99] A. Doménech, J. F. Fernández-Garayzábal, J. A. García, M. T. Cutuli, M. Blanco, A. Gibello, M. A. Moreno, and L. Domínguez. Association of Pseudomonas anguilliseptica infection with ‘winter disease’ in sea bream, Sparus aurata L. Journal of Fish Diseases, 22(1):69–71, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dannevig:1993:LAS

[DFK93] B. H. Dannevig, K. Falk, and J. Krogsrud. Leucocytes from Atlantic salmon, Salmo salar L., experimentally infected with infectious salmon anaemia (ISA) exhibit an impaired response to mitogens. Journal of Fish Diseases, 16(4):351–359, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dannevig:1994:IIT

[DFS94] B. H. Dannevig, K. Falk, and E. Skjerve. Infectivity of internal tissues of Atlantic salmon, Salmo salar L., experimentally infected with the aetiological agent of infectious salmon anaemia (ISA). Journal of Fish Diseases, 17(6):613–622, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Danley:1999:EFR

[DGK99] M. L. Danley, A. E. Goodwin, and H. S. Killian. Epizootics in farm-raised channel catfish, Ictalurus punctatus (Rafinesque), caused by the enteric redmouth bacterium Yersinia ruckeri. Journal of Fish Diseases, 22(6):451–456, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dale:1997:EVV

[DGR97] O. B. Dale, S. K. Gutenberger, and J. S. Rohovec. Estimation of variation of virulence of Renibacterium salmoninarum by survival analysis of experimental infection of salmonid fish. Journal of Fish Diseases, 20(3):177–183, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:1991:ITC

[DHB91] M. G. Dixon, T. Hecht, and C. R. Brandt. Identification and treatment of a Clostridium and Vibrio infection in South African abalone, Haliotis midae L. Journal of Fish Diseases, 14(6):693–695, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duis:1995:DQA

[DHB^{+}95] K. Duis, C. Hammer, M. C. M. Beveridge, V. Inglis, and E. Braum. Delivery of quinolone antibacterials to turbot, Scophthalmus maximus (L.), via bioencapsulation: quantification and efficacy trial. Journal of Fish Diseases, 18(3):229– 238, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dyer:1997:NBL

[DHM97] A. R. Dyer, P. De La M. Hall, and G. I. Matsumoto. Needle biopsy of liver in a marine teleost, Pagrus auratus (Bloch & Schneider). Journal of Fish Diseases, 20(6):463–466, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dopazo:1994:UCC

[DHS94] C. P. Dopazo, F. M. Hetrick, and S. K. Samal. Use of cloned cDNA probes for diagnosis of infectious pancreatic necrosis virus infections. Journal of Fish Diseases, 17(1):1–16, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Decostere:1999:IWQ

[DHTC99] A. Decostere, F. Haesebrouck, J. F. Turnbull, and G. Charlier. Influence of water quality and temperature on adhesion of high and low virulence Flavobacterium columnare strains to isolated gill arches. Journal of Fish Diseases, 22(1):1–11, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Decostere:1999:CAF

[DHV^{+}99] A. Decostere, F. Haesebrouck, E. Van Driessche, G. Charlier, and R. Ducatelle. Characterization of the adhesion of Flavobacterium columnare (Flexibacter columnaris) to gill tissue. Journal of Fish Diseases, 22(6):465–474, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1995:AIL

[DIB^{+}95] R. A. Dalmo, K. Ingebrigtsen, J. BóGwald, T. E. Horsberg, and R. Seljelid. Accumulation of immunomodulatory laminaran [β (1,3)-D-glucan] in the spleen and kidney of Atlantic salmon, Salmo solar L. Journal of Fish Diseases, 18(6): 545–553, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1997:NSD

[DIB97] R. A. Dalmo, K. Ingebrigtsen, and J. Bøgwald. Non-specific defence mechanisms in fish, with particular reference to the reticuloendothelial system (RES). Journal of Fish Diseases, 20(4):241–273, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1994:IAI

[DIHS94] R. A. Dalmo, K. Ingebrigtsen, T. E. Horsberg, and R. Seljelid. Intestinal absorption of immunomodulatory laminaran and derivatives in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 17(6):579–589, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1996:AIL

[DISS96] R. A. Dalmo, K. Ingebrigtsen, B. Sveinbjørnsson, and R. Seljelid. Accumulation of immunomodulatory laminaran (β (1,3)-D-glucan) in the heart, spleen and kidney of Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 19(2):129– 136, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Denham:1999:OGT

[DL99] K. L. Denham and J. Long. Occurrence of Gyrodactylus thymalli Zitnan, 1960 on grayling, Thymallus thymallus (L.), in England. Journal of Fish Diseases, 22(3):247–252, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1995:MAC

[DLB95] R. A. Dalmo, R. Mcqueen Leifson, and J. Bøgwald. Microspheres as antigen carriers: studies on intestinal absorption and tissue localization of polystyrene microspheres in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(1):87– 91, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davis:1994:DIP

[DLP^{+}94] P. J. Davis, L. A. Laidler, P. W. Perry, D. Rossington, and R. Alcock. The detection of infectious pancreatic necrosis virus in asymptomatic carrier fish by an integrated cellculture and ELISA technique. Journal of Fish Diseases, 17 (2):99–110, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diamant:1990:DXC

[DM90] A. Diamant and A. H. Mcvicar. Distribution of X-cell disease in common dab, Limanda limanda L., in the North Sea, and ultrastructural observations of previously undescribed developmental stages. Journal of Fish Diseases, 13(1):25–37, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1998:PED

[DMH^{+}98] R. A. Dalmo, B. Martinsen, T. E. Horsberg, A. Ramstad, C. Syvertsen, R. Seljelid, and K. Ingebrigtsen. Prophylactic effect of β (1,3)-D-glucan (laminaran) against experimental Aeromonas salmonicida and Vibrio salmonicida infections. Journal of Fish Diseases, 21(6):459–462, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

deMateo:1996:CSP

[dMMMK96] M. Marin de Mateo, J. McGeorge, D. Morris, and M. L. Kent. Comparative studies of PKX and Sphaerospora spp. from salmonids using lectin and monoclonal antibody staining techniques. Journal of Fish Diseases, 19(1):55–63, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:1997:HAC

[DMOR97] B. S. Dezfuli, M. Manera, S. Onestini, and R. Rossi. Histopathology of the alimentary canal of Anguilla anguilla L. associated with digenetic trematodes: a light and electron microscopic study. Journal of Fish Diseases, 20(4):317–322, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dodson:1999:BMT

[DMS99] S. V. Dodson, J. J. Maurer, and E. B. Shotts. Biochemical and molecular typing of Streptococcus iniae isolated from fish and human cases. Journal of Fish Diseases, 22(5):331–336, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1995:EGE

[DMTM95] A. J. Davies, C. Mastri, D. E. Thorborn, and D. Mackintosh. Experiments with growth of the eel trypanosome, Trypanosoma granulosum Laveran & Mesnil, 1902, in semidefined and defined media. Journal of Fish Diseases, 18(6): 599–608, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Derksen:1999:EHP

[DOF99] J. A. Derksen, V. E. Ostland, and H. W. Ferguson. Effects of hydrogen peroxide on clearance of formalin-killed Flavobacterium branchiophilum from the gills of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 22 (1):59–67, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deere:1996:DAS

[DPPE96] D. Deere, J. Porter, R. Pickup, and C. Edwards. Direct analysis of starved Aeromonas salmonicida. Journal of Fish Diseases, 19(6):459–467, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deere:1997:SCC

[DPPE97] D. Deere, J. Porter, R. W. Pickup, and C. Edwards. Short communicationMeasurement of cell surface hydrophobicity of Aeromonas salmonicida. Journal of Fish Diseases, 20(5):393– 397, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Drolet:1994:REP

[DRL94] B. S. Drolet, J. S. Rohovec, and J. C. Leong. The route of entry and progression of infectious haematopoietic necrosis virus in Oncorhynchus mykiss (Walbaum): a sequential immunohistochemical study. Journal of Fish Diseases, 17 (4):337–347, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Drinan:1992:MIA

[DRP92] E. M. Drinan, H. D. Rodger, and R. Palmer. Microsporidian infection in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(2):211–214, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1995:IEL

[DS95] R. A. Dalmo and R. Seljelid. The immunomodulatory effect of LPS, laminaran and sulphated laminaran [β (l,3)-D-glucan] on Atlantic salmon, Salmo salar L., macrophages in vitro. Journal of Fish Diseases, 18(2):175–185, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalmo:1998:TDC

[DSA^{+}98] R. A. Dalmo, T. Seternes, S. M. Arnesen, T. ò. Jørgensen, and J. Bøgwald. Tissue distribution and cellular uptake of Aeromonas salmonicida lipopolysaccharide (LPS) in some marine fish species. Journal of Fish Diseases, 21(5):321–334, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DelCorral:1990:AHC

[DSB90] F. Del Corral, E. B. Shotts, Jr., and J. Brown. Adherence, haemagglutination and cell surface characteristics of motile aeromonads virulent for fish. Journal of Fish Diseases, 13 (4):255–268, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dorson:1993:VHS

[DT93] M. Dorson and C. Torhy. Viral haemorrhagic septicaemia virus replication in external tissue excised from rainbow trout, Oncorhynchus mykiss (Walbaum), and hybrids of different susceptibilities. Journal of Fish Diseases, 16(4):403–408, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davie:1996:CAR

[DT96] P. S. Davie and H. Thorarensen. Coronary arteriosclerosis in rainbow trout, Oncorhynchus mykiss (Walbaum), is related to heart size rather than sex or reproductive status. Journal of Fish Diseases, 19(4):283–288, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1999:UCF

[DTMD99] A. J. Davies, D. E. Thorborn, C. Mastrl, and P. Daszak. Ultrastructure of culture forms of the eel trypanosome, Trypanosoma granulosum Laveran & Mesnil, 1902, exposed to polyamine biosynthesis inhibitors. Journal of Fish Diseases, 22(4):285–298, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dopazo:1992:ARA

[DTS^{+}92] C. P. Dopazo, A. E. Toranzo, S. K. Samal, B. S. Roberson, A. Baya, and F. M. Hetrick. Antigenic relationships among rotaviruses isolated from fish. Journal of Fish Diseases, 15 (1):27–36, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dierckens:1998:AHC

[DVB^{+}98] Dierckens, Vandenberghe, Beladjal, Huys, Mertens, and Swings. Aeromonas hydrophila causes ‘black disease’ in fairy shrimps (Anostraca; Crustacea). Journal of Fish Diseases, 21 (2):113–119, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Daly:1994:PAC

[DWV94] J. G. Daly, G. F. Wiegertjes, and W. B. Van Muiswinkel. Protection against carp erythrodermatitis following bath or subcutaneous exposure to sublethal numbers of virulent Aeromonas salmonicida subsp. nova. Journal of Fish Diseases, 17(1):67–75, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Esteve:1991:SPA

[EA91] C. Esteve and C. Amato. Siderophore production in Aeromona spp. Isolated from European eel, Anguilla anfuilla L. Journal of Fish Diseases, 14(3):423–427, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Effendi:1994:SFP

[EA94] I. Effendi and B. Austin. Survival of the fish pathogen Aeromonas salmonicida in the marine environment. Journal of Fish Diseases, 17(4):375–385, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eggset:1994:EGC

[EBL^{+}94] G. Eggset, R. Bjørnsdottir, R. Mcqueen Leifson, J. A. Arnesen, D. H. Coucheron, and T. Ø. Jørgensen. Extracellular glycerophospholipid:cholesterol acyltransferase from Aeromonas salmonicida: activation by serine protease. Journal of Fish Diseases, 17(1):17–29, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Estepa:1991:IMS

[EC91] A. Estepa and J. M. Coll. Infection of mitogen-stimulated trout leucocytes with salmonid viruses. Journal of Fish Diseases, 14(5):555–562, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eiras:1991:HBL

[ED91] J. C. Eiras and A. J. Davies. Haemogregarina bigemina Laveran & Mesnil, 1901 (Protozoa, Apicomplexa) from Blennius pholis L. (Osteichthyes, Blenniidae): an investigation of seasonality in Portugal. Journal of Fish Diseases, 14(6):683–687, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Enger:1992:ILA

[EGTH92] Ø. Enger, B. Gunnlaugsdottir, B. K. Thorsen, and B. Hjeltnes. Infectious load of Aeromonas salmonicida subsp. salmonicida during the initial phase of a cohabitant infection experiment with Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(5):425–430, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eaton:1991:EHN

[EHH91] B. T. Eaton, A. D. Hyatt, and S. Hengstberger. Epizootic haematopoietic necrosis virus: purification and classification. Journal of Fish Diseases, 14(2):157–169, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1991:AET

[Ell91] A. E. Ellis. An appraisal of the extracellular toxins of Aeromonas salmonicida ssp. salmonicida. Journal of Fish Diseases, 14(3):265–277, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ekless:1993:IMA

[EM93] L. M. Ekless and R. A. Matthews. Ichthyophthirius multifiliis: axenic isolation and short-term maintenance in selected monophasic media. Journal of Fish Diseases, 16(5):437–447, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1990:CPE

[EMM90] A. E. Ellis, N. Masson, and A. L. S. Munro. A comparison of proteases extracted from Caligus elongatus (Nordmann, 1832) and Lepeophtheirus salmonis (Kroyer, 1838). Journal of Fish Diseases, 13(2):163–165, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Espelid:1991:VES

[ERJ91] S. Espelid, O. M. Rødseth, and T. Ø. Jørgensen. Vaccination experiments and studies of the humoral immune responses in cod, Gadus morhua L., to four strains of monoclonal-defined Vibrio anguillarum. Journal of Fish Diseases, 14(2):185–197, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Estevez:1998:SCC

[ERSA98] Estévez, Rodríguez, Soto, and Arias. Short communication Cross-reactivity between Aggregata octopiana Schneider 1875 and Aggregata eberthi Labbé 1895 (Apicomplexa, Aggregatidae): two coccidian parasites of cephalopods. Journal of Fish Diseases, 21(4):309–312, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

F:1992:DSD

[F.92] C. J. Rodgers M. A. F. F. Development of a selectivedifferential medium for the isolation of Yersinia ruckeri and its application in epidemiological studies. Journal of Fish Diseases, 15(3):243–254, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Alonso:1999:MSF

[FAAE^{+}99] M. Fernandez-Alonso, F. Alvarez, A. Estepa, R. Blasco, and J. M. Coll. A model to study fish DNA immersion vaccination by using the green fluorescent protein. Journal of Fish Diseases, 22(3):237–241, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Alonso:1999:GDB

[FAAEC99] M. Fernandez-Alonso, F. Alvarez, A. Estepa, and J. M. Coll. G disulphide bond native conformation is required to elicit trout neutralizing antibodies against VHSV. Journal of Fish Diseases, 22(3):219–222, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fernandez-Alonso:1999:IFV

[FAC99] M. Fernandez-Alonso and J. M. Coll. Induced fusion of VHSV persistently infected fish cells. Journal of Fish Diseases, 22 (5):401–406, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fraser:1992:ASA

[FCC92] G. C. Fraser, R. B. Callinan, and L. M. Calder. Aphanomyces species associated with red spot disease: an ulcerative disease of estuarine fish from eastern Australia. Journal of Fish Diseases, 15(2):173–181, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fitzgerald:1991:MSC

[FCS91] S. D. Fitzgerald, W. W. Carlton, and G. Sandusky. Metastatic squamous cell carcinoma in a hybrid sunfish. Journal of Fish Diseases, 14(4):481–487, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farrell:1992:ACA

[FDS92] A. P. Farrell, P. S. Davie, and R. Sparksman. The absence of coronary arterial lesions in five species of elasmobranchs, Raja nasuta, Squalus acanthias L., Isurus oxyrinchus Rafinesque, Prionace glauca (L.) and Lamna nasus (Bonnaterre). Journal of Fish Diseases, 15(6):537–540, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frerichs:1991:IBP

[FH91] G. N. Frerichs and A. Holliman. Isolation of a brown pigmentproducing strain of Pseudomonas fluorescens cross-reacting with Aeromonas salmonicida diagnostic antisera. Journal of Fish Diseases, 14(5):599–601, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Forsythe:1991:OBP

[FHBN91] J. W. Forsythe, R. T. Hanlon, R. A. Bullis, and E. J. Noga. Octopus bimaculoides (Pickford & McConnaughey, 1949): a marine invertebrate host for ectoparasitic protozoans. Journal of Fish Diseases, 14(4):431–442, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Findlay:1995:DRR

[FHMG95] V. L. Findlay, M. Helders, B. L. Munday, and R. Gurney. Demonstration of resistance to reinfection with Paramoeba sp. by Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(6):639–642, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farrell:1990:CLP

[FJS90] A. P. Farrell, J. A. Johansen, and R. L. Saunders. Coronary lesions in Pacific salmonids. Journal of Fish Diseases, 13(1): 97–100, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fang:1998:VIE

[FLT^{+}98] Fang, Ling, Tan, Ge, and Sin. In vitro inhibition of epithelial cell invasion by Aeromonashydrophila Vibrio species by fish Aeromonas hydrophila major adhesin. Journal of Fish Diseases, 21(4):273–280, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Findlay:1998:FSA

[FM98] Findlay and Munday. Further studies on acquired resistance to amoebic gill disease (AGD) in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21(2):121–125, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Franklin:1993:XCG

[FMDC93] C. E. Franklin, J. C. Mckenzie, W. Davison, and P. W. Carey. X-cell gill disease obliterates the lamellar blood supply in the Antarctic teleost, Pagothenia borchgreviniki (Boulenger 1902). Journal of Fish Diseases, 16(3):249–254, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fujiki:1994:PES

[FMY94] K. Fujiki, H. Matsuyama, and T. Yano. Protective effect of sodium alginates against bacterial infection in common carp, Cyprinus carpio L. Journal of Fish Diseases, 17(4):349–355, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fouz:1995:HLC

[FNTF95] B. Fouz, B. Novoa, A. E. Toranzo, and A. Figueras. Histopathological lesions caused by Vibrio damsela in cultured turbot, Scophthalmus maximus (L.): inoculations with live cells and extracellular products. Journal of Fish Diseases, 18(4):357–364, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1995:NMS

[FOM^{+}95] H. W. Ferguson, V. E. Ostland, D. D. Macphee, J. S. Lumsden, P. J. Byrne, and D. J. Speare. ‘No-mucus skin disease’ of rainbow trout, Oncorhynchus mykiss (Walbaum): a case report. Journal of Fish Diseases, 18(1):49–57, January 1995.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frerichs:1996:IEM

[Fre96] G. N. Frerichs. Identification and elimination of mycoplasmas in fish cell line cultures. Journal of Fish Diseases, 19(6): 435–439, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Faye:1990:UDP

[FTB90] N. Faye, B. S. Toguebaye, and G. Bouix. Ultrastructure and development of Pleistophora senegalensis sp. nov. (Protozoa, Microspora) from the gilt-head sea bream, Sparus aurata L. (Teleost, Sparidae) from the coast of Senegal. Journal of Fish Diseases, 13(3):179–192, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frances:1997:ESP

[FTN97] J. Frances, R. Tennent, and B. F. Nowak. Epitheliocystis in silver perch, Bidyanus bidyanus (Mitchell). Journal of Fish Diseases, 20(6):453–457, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:1996:BCM

[FW96] S. Feng and P. T. K. Woo. Biological characterization of a monoclonal antibody against a surface membrane antigen on Cryptobia salmositica Katz, 1951. Journal of Fish Diseases, 19(2):137–143, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:1998:VVE

[FW98] Feng and Woo. In vitro and in vivo effects of rabbit anti-thymocyte serum on circulating leucocytes and production of complement fixing antibodies in thymectomized Oncorhynchus mykiss (Walbaum) infected with Cryptobia salmositica Katz 1951. Journal of Fish Diseases, 21(4):241– 248, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frischknecht:1994:CPC

[FWM94] R. Frischknecht, T. Wahli, and W. Meier. Comparison of pathological changes due to deficiency of vitamin C, vitamin E and combinations of vitamins C and E in rainbow trout, Oncorhynchus mykiss Walbaum. Journal of Fish Diseases, 17 (1):31–45, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gerundo:1991:PER

[GACHF91] N. Gerundo, D. J. Alderman, R. S. Clifton-Hadley, and S. W. Feist. Pathological effects of repeated doses of malachite green: a preliminary study. Journal of Fish Diseases, 14(5): 521–532, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gatesoupe:1991:EIT

[Gat91] F.-J. Gatesoupe. Experimental infection of turbot, Scophthalmus maximus (L.), larvae with a strain of Aeromonas hydrophila. Journal of Fish Diseases, 14(4):495–498, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:1993:FMM

[GBG^{+}93] S. Gómez, A. Bernabé, M. A. Gómez, J. A. Navarro, and J. Sánchez. Fish mycobacteriosis: morphopathological and immunocytochemical aspects. Journal of Fish Diseases, 16 (2):137–141, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:1993:DRS

[GBH93] S. Gudmundsdottir, E. Benediktsdottir, and S. Helgason. Detection of Renibacterium salmoninarum in salmonid kidney samples: a comparison of results using double-sandwich ELISA and isolation on selective medium. Journal of Fish Diseases, 16(3):185–195, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grisez:1991:UAI

[GCO91] L. Grisez, R. Ceusters, and F. Ollevier. The use of API 20E for the identification of Vibrio anguillarum and V. ordalii. Journal of Fish Diseases, 14(3):359–365, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gaggero:1995:FIP

[GCS95] A. Gaggero, H. Castro, and A. M. Sandino. First isolation of Piscirickettsia salmonis from Coho salmon, Oncorhynchus kisutch (Walbaum), and rainbow trout, Oncorhynchus mykiss (Walbaum), during the freshwater stage of their life cycle. Journal of Fish Diseases, 18(3):277–280, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:1997:ECP

[GG97] B. K. Gudmundsdóttir and S. Gudmundsdóttir. Evaluation of cross protection by vaccines against atypical and typical furunculosis in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 20(5):343–350, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Greger:1999:VDW

[GG99] E. Greger and T. Goodrich. Vaccine development for winter ulcer disease, Vibrio viscosus, in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 22(3):193–199, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

GudmundsdoTtir:1991:CET

[GHB91] S. GudmundsdóTtir, S. Helgason, and E. Benediktsdottir. Comparison of the effectiveness of three different growth media for primary isolation of Renibacterium salmoninarum from Atlantic salmon, Salmo salar L., broodfish. Journal of Fish Diseases, 14(1):89–96, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Glazebrook:1990:PLV

[GHdB90] J. S. Glazebrook, M. P. Heasman, and S. W. de Beer. Picornalike viral particles associated with mass mortalities in larval barramundi, Lates calcarifer Bloch. Journal of Fish Diseases, 13(3):245–249, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:1997:SHA

[GJS^{+}97] B. K. Gudmundsdóttir, H. Jónsdóttir, V. Steinthórsdóttir, B. Magnadóttir, and S. Gudmundsdóttir. Survival and humoral antibody response of Atlantic salmon, Salmo salar L., vaccinated against Aeromonas salmonicida ssp. achromogenes. Journal of Fish Diseases, 20(5):351–360, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Griffiths:1991:IMS

[GL91] S. G. Griffiths and W. H. Lynch. Instability of the major soluble antigen produced by Renibacterium salmoninarum. Journal of Fish Diseases, 14(1):55–66, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gutierrez:1993:PPE

[GMHK93] M. A. Gutierrez, T. Miyazaki, H. Hatta, and M. Kim. Protective properties of egg yolk IgY containing anti- Edwardsiella tarda antibody against paracolo disease in the Japanese eel, Anguilla japonica Temminck & Schlegel. Journal of Fish Diseases, 16(2):113–122, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:1998:SCM

[Góm98] S. Gómez. Short communicationUnusual morphopathological features in a case of fish tuberculosis. Journal of Fish Diseases, 21(3):237–239, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gierer:1992:SPF

[GRR92] W. Gierer, W. Rabsch, and R. Reissbrodt. Siderophore pattern of fish-pathogenic Vibrio anguiliarum, Aeromonas spp. and Pseudomonas spp. from the German Baltic coast. Journal of Fish Diseases, 15(5):417–423, September 1992. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gonzalez:1999:EIP

[GSC^{+}99] M. González, F. Sánchez, M. I. Concha, J. Figueroa, M. I. Montecinos, and G. León. Evaluation of the internalization process of the fish pathogen Renibacterium salmoninarum in cultured fish cells. Journal of Fish Diseases, 22(3):231–235, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson:1998:IPN

[GSS98] D. R. Gibson, D. A. Smail, and C. Sommerville. Infectious pancreatic necrosis virus: experimental infection of goldsinny wrasse, Ctenolabrus rupestris L. (Labridae). Journal of Fish Diseases, 21(6):399–406, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Granzow:1997:SUM

[GWFE97] H. Granzow, F. Weiland, D. Fichtner, and P. J. Enzmann. Studies of the ultrastructure and morphogenesis of fish pathogenic viruses grown in cell culture. Journal of Fish Diseases, 20(1):1–10, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gray:1999:DCC

[GWG99] W. L. Gray, R. J. Williams, and B. R. Griffin. Detection of channel catfish virus DNA in acutely infected channel catfish, Ictalurus punctatus (Rafinesque), using the polymerase chain reaction. Journal of Fish Diseases, 22(2):111–116, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hanna:1992:DMA

[HAC^{+}92] P. J. Hanna, K. Altmann, D. Chen, A. Smith, S. Cosic, P. Moon, and L. S. Hammond. Development of monoclonal antibodies for the rapid identification of epizootic Vibrio species. Journal of Fish Diseases, 15(1):63–69, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoglund:1992:HRE

[HAH92] J. Höglund, J. Andersson, and J. Härdig. Haematological responses in the European eel, Anguilla anguilla L., to sublethal infestation by Anguillicola crassus in a thermal effluent of the Swedish Baltic. Journal of Fish Diseases, 15(6): 507–514, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heupel:1996:MPM

[HB96] M. R. Heupel and M. B. Bennett. A myxosporean parasite (Myxosporea: Multivalvulida) in the skeletal muscle of epaulette sharks, Hemiscyllium ocellatum (Bonnaterre), from the Great Barrier Reef. Journal of Fish Diseases, 19(2):189– 191, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heupel:1998:IES

[HB98] M. R. Heupel and M. B. Bennett. Infection of the epaulette shark, Hemiscyllium ocellatum (Bonnaterre), by the nematode parasite Proleptus australis Bayliss (Spirurida: Physalopteridae). Journal of Fish Diseases, 21(6):407–414, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huntly:1992:NLE

[HCM92] P. J. Huntly, G. Coleman, and A. L. S. Munro. The nature of the lethal effect on Atlantic salmon, Salmo salar L., of a lipopolysaccharide-free phospholipase activity isolated from the extracellular products of Aeromonas salmonicida. Journal of Fish Diseases, 15(1):99–102, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hsu:1993:DIP

[HCW93] Y.-L. Hsu, B.-S. Chen, and J.-L. Wu. Demonstration of infectious pancreatic necrosis virus strain VR-299 in Japanese eel, Anguilla japonica Temminck & Schlegel. Journal of Fish Diseases, 16(2):123–129, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haavarstein:1990:SIS

[HEH^{+}90] L. S. Håvarstein, C. Endresen, B. Hjeltnes, K. E. Christie, and J. Glette. Specific immunoglobulins in serum from Atlantic salmon, Salmo salar L., immunized with Vibrio salmonicida and infectious pancreatic necrosis virus. Journal of Fish Diseases, 13(2):101–111, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hyatt:1991:EHN

[HEHR91] A. D. Hyatt, B. T. Eaton, S. Hengstberger, and G. Russel. Epizootic haematopoietic necrosis virus: detection by ELISA, immunohistochemistry and immunoelectronmicroscopy. Journal of Fish Diseases, 14(6):605–617, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hirst:1994:UHC

[HHE94] I. D. Hirst, T. S. Hastings, and A. E. Ellis. Utilization of haem compounds by Aeromonas salmonicida. Journal of Fish Diseases, 17(4):365–373, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harada:1990:LLM

[HHKE90] T. Harada, J. Hatanaka, S. S. Kubota, and M. Enomoto. Lymphoblastic lymphoma in medaka, Oryzias latipes (Temminck et Schlegel). Journal of Fish Diseases, 13(2):169–173, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hudson:1993:ITS

[HHS93] D. A. Hudson, N. B. Hudson, and J. D. Shields. Infection of Trapezia spp. (Decapoda: Xanthidae) by Hematodinium sp. (Duboscquodinida: Syndinidae): a new family record of infection. Journal of Fish Diseases, 16(3):273–276, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harada:1996:MST

[HIH^{+}96] T. Harada, H. Itoh, J. Hatanaka, S. Kamiya, and M. Enomoto. A morphological study off a thyroid carcinoma in a medaka, Oryzias latipes (Temminck & Schlegel). Journal of Fish Diseases, 19(4):271–277, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayashi:1993:MAC

[HIMK93] Y. Hayashi, H. Izawa, T. Mikami, and H. Kodama. A monoclonal antibody cross-reactive with three salmonid herpesviruses. Journal of Fish Diseases, 16(5):479–486, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Higgins:1998:TAF

[HK98] M. J. Higgins and M. L. Kent. TNP-470, the analogue of fumagillin-DCH, controls PKX in naturally infected sockeye salmon, Oncorhynchus nerka (Walbaum), underyearlings. Journal of Fish Diseases, 21(6):455–458, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hong:1998:IAS

[HLHW98] J-R. Hong, T-L. Lin, Y-L. Hsu, and J-L. Wu. Induction of apoptosis and secondary necrosis by infectious pancreatic necrosis virus in fish embryonic cells. Journal of Fish Diseases, 21(6):463–468, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Husevaag:1991:SOV

[HLJ^{+}91] B. Husevåg, B. T. Lunestad, P. J. Johannessen, Ø. Enger, and O. B. Samuelsen. Simultaneous occurrence of Vibrio salmonicida and antibiotic-resistant bacteria in sediments at abandoned aquaculture sites. Journal of Fish Diseases, 14(6): 631–640, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hung:1996:HIR

[HLTK96] H-W. Hung, C-F. Lo, C-C. Tseng, and G-H. Kou. Humoral immune response of Japanese eel, Anguilla japonica Temminck & Schlegel, to Pleistophora anguillarum Hoshina, 1951 (Microspora). Journal of Fish Diseases, 19(3):243–250, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hung:1997:APJ

[HLTK97a] H-W. Hung, C-F. Lo, C-C. Tseng, and G-H. Kou. Antibody production in Japanese eels, Anguilla japonica Temminck & Schlegel. Journal of Fish Diseases, 20(3):195–200, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hung:1997:ARG

[HLTK97b] H-W. Hung, C-F. Lo, C-C. Tseng, and G-H. Kou. Antibody response of glass eels, Anguilla japonica Temminck & Schlegel, to Pleistophora anguillarum Hoshina (Microspora) infection. Journal of Fish Diseases, 20(3):237–239, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hatai:1993:MIP

[HLTS93] K. Hatai, O. Lawhavinit, K. Toda, and Y. Sugou. Mycobacterium infection in pejerrey, Odonthestes bonariensis Cuvier & Valenciennes. Journal of Fish Diseases, 16(4):397–402, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Houghton:1990:IJC

[HM90] G. Houghton and R. A. Matthews. Immunosuppression in juvenile carp, Cyprinus carpio L.: the effects of the corticosteroids triamcinolone acetonide and hydrocortisone 21hemisuccinate (cortisol) on acquired immunity and the humoral antibody response to Ichthyophthirius multifiliis Fouquet. Journal of Fish Diseases, 13(4):269–280, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hellberg:1996:IAA

[HMH96] H. Hellberg, E. Moksness, and S. Høie. Infection with atypical Aeromonas salmonicida in farmed common wolffish, Anarhichas lupus L. Journal of Fish Diseases, 19(4):329– 332, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoole:1994:USI

[HN94] D. Hoole and H. Nisan. Ultrastructural studies on intestinal response of carp, Cyprinus carpio L., to the pseudophyllidean tapeworm, Bothriocephalus acheilognathi Yamaguti, 1934. Journal of Fish Diseases, 17(6):623–629, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huizinga:1997:CRG

[HN97] H. W. Huizinga and M. J. Nadakavukaren. Cellular responses of goldfish, Carassius auratus (L.), to metacercariae of Ribeiroia marini (Faust & Hoffman, 1934). Journal of Fish Diseases, 20(6):401–408, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hovland:1994:OIS

[HNWE94] T. Hovland, A. Nylund, K. Watanabe, and C. Endresen. Observation of infectious salmon anaemia virus in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 17(3):291– 296, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heasman:1996:DBI

[HOF96] M. P. Heasman, W. A. O’Connor, and A. W. J. Frazer. Digenean (Bucephalidae) infections in commercial scallops, Pecten fumatus Reeve, and doughboy scallops, Chlamys (Mimachlamys) asperrima (Lamarck), in Jervis Bay, New South Wales. Journal of Fish Diseases, 19(5):333–339, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hallett:1995:EID

[HR95] S. L. Hallett and F. R. Roubal. Experiments on the infection dynamics of Caligus epidemicus (Copepoda: Caligidae) on the small marine fish, Ambassis marianus (Günther). Journal of Fish Diseases, 18(1):59–66, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hansen:1990:IEA

[HRO90] G. H. Hansen, J. Raa, and J. A. Olafsen. Isolation of Enterobacter agglomerans from dolphin fish, Coryphaena hippurus L. Journal of Fish Diseases, 13(1):93–96, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hosono:1994:ERJ

[HSK94] N. Hosono, S. Suzuki, and R. Kusuda. Evidence for relatedness of Japanese isolates of birnaviruses from marine fish to IPNV. Journal of Fish Diseases, 17(4):433–437, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hosono:1996:GBI

[HSK96] N. Hosono, S. Suzuki, and R. Kusuda. Genogrouping of birnaviruses isolated from marine fish: a comparison of VP2/ NS junction regions on genome segment A. Journal of Fish Diseases, 19(4):295–302, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Handlinger:1997:PFM

[HSP97] J. Handlinger, M. Soltani, and S. Percival. The pathology of Flexibacter maritimus in aquaculture species in Tasmania, Australia. Journal of Fish Diseases, 20(3):159–168, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Houghton:1991:DRC

[HWGV91] G. Houghton, G. F. Wiergertjes, A. Groeneveld, and W. B. Van Muiswinkel. Differences in resistance of carp, Cyprinus carpio L., to atypical Aeromonas salmonicida. Journal of Fish Diseases, 14(3):333–341, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hirose:1998:PAC

[HYHA98] Hirose, Yabu, Hirono, and Aoki. The phylogeny of Anguillicola crassus and A. globiceps based on partial 18S ribosomal RNA sequences. Journal of Fish Diseases, 21(4):265–271, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1991:ARA

[IFMR91] V. Inglis, G. N. Frerichs, S. D. Millar, and R. H. Richards. Antibiotic resistance of Aeromonas salmonicida isolated from Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Diseases, 14(3):353–358, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ibarra:1992:EIR

[IGH92] A. M. Ibarra, G. A. E. Gall, and R. P. Hedrick. Experimental infection of rainbow trout, Oncorhynchus mykiss (Walbaum), with cryopreserved developmental and sporogonic stages of the myxosporean Ceratomyxa shasta (Noble). Journal of Fish Diseases, 15(4):353–355, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Imagawa:1990:LHS

[IHKS90] T. Imagawa, Y. Hashimoto, Y. Kon, and M. Sugimura. Lectin histochemistry as special markers for rodlet cells in carp, Cyprinus carpio L. Journal of Fish Diseases, 13(6):537–540, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Itami:1996:ERA

[IKU^{+}96] T. Itami, M. Kondo, M. Uozu, A. Suganuma, T. Abe, A. Nakagawa, N. Suzuki, and Y. Takahashi. Enhancement of resistance against Enterococcus seriolicida infection in yellowtail, Seriola quinqueradiata (Temminck & Schlegel), by oral administration of peptidoglycan derived from Bifidobacterium thermophilum. Journal of Fish Diseases, 19(2):185–187, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Imagawa:1998:ABR

[IKU98] Imagawa, Kitagawa, and Uehara. An association between rodlet cells and the vascular endothelial cells in the head kidney of carp, Cyprinus carpio L.: ultrastructural observations. Journal of Fish Diseases, 21(2):153–157, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1993:RAS

[IMR93] V. Inglis, S. D. Millar, and R. H. Richards. Resistance of Aeromonas salmonicida to amoxicillin. Journal of Fish Diseases, 16(4):389–395, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1991:VSA

[IR91] V. Inglis and R. H. Richards. The in vitro susceptibility of Aeromonas salmonicida and other fish-pathogenic bacteria to 29 antimicrobial agents. Journal of Fish Diseases, 14(6): 641–650, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1996:APA

[IRM^{+}96] V. Inglis, D. Robertson, K. Miller, K. D. Thompson, and R. H. Richards. Antibiotic protection against recrudescence of latent Aeromonas salmonicida during furunculosis vaccination. Journal of Fish Diseases, 19(5):341–348, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1991:FAS

[IRV^{+}91] V. Inglis, R. H. Richards, K. J. Varma, I. H. Sutherland, and E. S. Brokken. Florfenicol in Atlantic salmon, Salmo salar L., parr: tolerance and assessment of efficacy against furunculosis. Journal of Fish Diseases, 14(3):343–351, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inoue:1998:SCD

[ISM^{+}98] M. Inoue, S. Satoh, M. Maita, V. Kiron, and N. Okamoto. Short communicationRecovery from derangement of natural killer-like activity of leucocytes due to Zn or Mn deficiency in rainbow trout, Oncorhynchus mykiss (Walbaum), by the oral administration of these elements. Journal of Fish Diseases, 21(3):233–236, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Inglis:1993:PMA

[IYBF93] V. Inglis, E. Yimer, E. J. Bacon, and S. Ferguson. Plasmidmediated antibiotic resistance in Aeromonas salmonicida isolated from Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Diseases, 16(6):593–599, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jimenez:1998:PIO

[JBRM98] R. Jiménez, R. Barniol, X. Romero, and M. Machuca. A prokaryotic intracellular organism in the cuticular epithelium of cultured crayfish, Cherax quadricarinatus (von Martens), in Ecuador. Journal of Fish Diseases, 21(5):387–390, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonsdottir:1992:HAP

[JBWT92] H. Jónsdóttir, J. E. Bron, R. Wootten, and J. F. Turnbull. The histopathology associated with the pre-adult and adult stages of Lepeophtheirus salmonis on the Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(6):521–527, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jarp:1995:RFF

[JGOB95] J. Jarp, A. G. Gjevre, A. B. Olsen, and T. Bruheim. Risk factors for furunculosis, infectious pancreatic necrosis and mortality in post-smolt of Atlantic salmon, Salmo solar L. Journal of Fish Diseases, 18(1):67–78, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansson:1991:DRS

[JHL^{+}91] E. Jansson, T. Hongslo, R. Lindberg, O. Ljungberg, and B.- M. Svensson. Detection of Renibacterium salmoninarum and Yersinia ruckeri by the peroxidase-antiperoxidase immunohistochemical technique in melanin-containing cells of fish tissues. Journal of Fish Diseases, 14(6):689–692, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jastrzebski:1990:LEM

[JK90] M. Jastrzebski and Z. Komorowski. Light and electron microscopic studies on Goussia zarnowskii (Jastrzebski, 1982): an intestinal coccidium parasitizing the three-spined stickleback, Gasterosteus aculeatus (L.). Journal of Fish Diseases, 13(1):1–24, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:1993:PHK

[JLR93] J. B. Jørgensen, H. Lunde, and B. Robertsen. Peritoneal and head kidney cell response to intraperitoneally injected yeast glucan in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 16(4):313–325, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung:1995:HHN

[JM95] S. J. Jung and T. Miyazaki. Herpesviral haematopoietic necrosis of goldfish, Carassius auratus (L.). Journal of Fish Diseases, 18(3):211–220, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jang:1995:EGR

[JMK^{+}95] S. I. Jang, M. J. Marsden, Y. G. Kim, M. S. Choi, and C. J. Secombes. The effect of glycyrrhizin on rainbow trout, Oncorhynchus mykiss (Walbaum), leucocyte responses. Journal of Fish Diseases, 18(4):307–315, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:1990:GAD

[Jon90] J. B. Jones. Goussia auxidis (Dogiel, 1948) (Apicomplexa: Calyptosporidae) from tuna (Pisces: Scombridae) in the South Pacific. Journal of Fish Diseases, 13(3):215–223, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Joborn:1997:CFI

[JOW^{+}97] A. Jöborn, J. C. Olsson, A. Westerdahl, P. L. Conway, and S. Kjelleberg. Colonization in the fish intestinal tract and production of inhibitory substances in intestinal mucus and faecal extracts by Carnobacterium sp. strain K1. Journal of Fish Diseases, 20(5):383–392, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jimenez:1998:IDC

[JR98] R. Jiménez and X. Romero. Iron deposits on the cuticle and in the hepatopancreas of the Australian redclaw crayfish, Cherax quadricarinatus (Von Martens). Journal of Fish Diseases, 21 (5):395–398, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jelmert:1995:NAM

[JRV95] A. Jelmert, O. M. Rødseth, and T. Van Der Meeren. Nephrocalcinosis associated with mass mortality in cultured Atlantic halibut, Hippoglossus hippoglossus L., larvae. Journal of Fish Diseases, 18(4):365–369, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johansen:1995:MIP

[JS95] L.-H. Johansen and A.-I. Sommer. Multiplication of infectious pancreatic necrosis virus (IPNV) in head kidney and blood leucocytes isolated from Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(2):147–156, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:1990:HAJ

[JSB90] M. W. Jones, C. Sommerville, and J. Bron. The histopathology associated with the juvenile stages of Lepeophtheirus salmonis on the Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 13(4):303–310, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:1992:RSS

[JSW92] M. W. Jones, C. Sommerville, and R. Wootten. Reduced sensitivity of the salmon louse, Lepeophtheirus salmonis, to the organophosphate dichlorvos. Journal of Fish Diseases, 15 (2):197–202, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kakuta:1998:RSR

[Kak98] I. Kakuta. Reduction of stress response in carp, Cyprinus carpio L., held under deteriorating environmental conditions, by oral administration of bovine lactoferrin. Journal of Fish Diseases, 21(3):161–167, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kjuul:1999:ECP

[KBE^{+}99] A. K. Kjuul, E. E. Büllesbach, S. Espelid, R. Dunham, T. Ò. Jørgensen, G. W. Warr, and O. B. Styrvold. Effects of cecropin peptides on bacteria pathogenic to fish. Journal of Fish Diseases, 22(5):387–394, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kraxberger-Beatty:1990:VHO

[KBMGL90] T. Kraxberger-Beatty, D. J. Mcgarey, H. J. Grier, and D. V. Lim. Vibrio harveyi, an opportunistic pathogen of common snook, Centropomus undecimalis (Bloch), held in captivity. Journal of Fish Diseases, 13(6):557–560, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumaraguru:1995:IPH

[KBW95] A. K. Kumaraguru, F. W. H. Beamish, and P. T. K. Woo. Impact of a pathogenic haemoflagellate, Cryptobia salmositica Katz, on the metabolism and swimming performance of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 18(4):297–305, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knuckey:1995:LIB

[KDC95] I. A. Knuckey, P. J. F. Davie, and L. R. G. Cannon. Loxothylacus ihlei Boschma, (Rhizocephala) and its effects on the mud crab, Scylla serrata (Forskål), in northern Australia. Journal of Fish Diseases, 18(5):389–395, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khoo:1995:RLO

[KDL95] L. Khoo, P. M. Dennis, and G. A. Lewbart. Rickettsialike organisms in the blue-eyed plecostomus, Panaque suttoni (Eigenmann & Eigenmann). Journal of Fish Diseases, 18(2): 157–164, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kitao:1991:APM

[KEY91] T. Kitao, T. Eshima, and T. Yoshida. Analysis of protective mechanisms in cultured ayu, Plecoglossus altivelis Temminck and Schlegel, administered Vibrio vaccine by the immersion method. Journal of Fish Diseases, 14(3):375–381, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kennedy:1990:ENT

[KGW90] C. J. Kennedy, K. A. Gill, and P. J. Walsh. The effects of nifurpirinol treatment on the activities of hepatic xenobiotic transforming enzymes in the gulf toadfish, Opsanus beta (Goode and Bean). Journal of Fish Diseases, 13(6):525–529, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karunasagar:1992:SCF

[KKP92] I. Karunasagar, I. Karunasagar, and R. Pai. Systemic Citrobacter Freundii infection in common carp, Cyprinus carpio L., fingerlings. Journal of Fish Diseases, 15(1):95–98, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:1998:DPC

[KKS^{+}98] Kim, Kim, Sohn, Sim, Park, Heo, Lee, Lee, Jun, and Jang. Development of a polymerase chain reaction (PCR) procedure for the detection of baculovirus associated with white spot syndrome (WSBV) in penaeid shrimp. Journal of Fish Diseases, 21(1):11–17, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khoo:1998:BIF

[KLW^{+}98] L. Khoo, A. T. Leard, P. R. Waterstrat, S. W. Jack, and K. L. Camp. Branchiomyces infection in farm-reared channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 21(6):423–432, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kakuta:1990:ITT

[KN90a] I. Kakuta and K. Namba. Isometric twitch tension of lateral muscle in carp, Cyprinus carpio L., with sekoke disease. Journal of Fish Diseases, 13(2):135–144, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kusser:1990:DOR

[KN90b] W. C. Kusser and S. G. Newman. Detection of oxytetracycline residues in fish tissues using a sensitive bioassay. Journal of Fish Diseases, 13(6):545–548, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kawahara:1990:TIA

[KON90] E. Kawahara, S. Oshima, and S. Nomura. Toxicity and immunogenicity of Aeromonas salmonicida extracellular products to salmonids. Journal of Fish Diseases, 13(6):495–503, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kantham:1995:EBG

[KR95] K. P. L. Kantham and R. H. Richards. Effect of buffers on the gill structure of common carp, Cyprinus carpio L., and rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 18(5):411–423, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karunasagar:1991:IRI

[KRK91] I. Karunasagar, G. Rosalind, and I. Karunasagar. Immunological response of the Indian major carps to Aeromonas hydrophila vaccine. Journal of Fish Diseases, 14(3):413–417, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khan:1996:PSH

[KRL96] M. N. Khan, R. Renaud, and J. F. Leatherland. Plasma steroid hormone profile in an anovulatory Arctic chair, Salvelinus alpinus (L.), exhibiting ovarian cysts. Journal of Fish Diseases, 19(5):389–394, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knappskog:1993:IAV

[KRSE93] D. H. Knappskog, O. M. Rødseth, E. Slinde, and C. Endresen. Immunochemical analyses of Vibrio anguillarum strains isolated from cod, Gadus morhua L., suffering from vibriosis. Journal of Fish Diseases, 16(4):327–338, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knudsen:1999:SAV

[KSPO99] G. Knudsen, H. Sørum, C. Mc.L. Press, and J. A. Olafsen. In situ adherence of Vibrio spp. to cryosections of Atlantic salmon, Salmo salar L., tissue. Journal of Fish Diseases, 22 (6):409–418, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karczewski:1991:SSA

[KSS91] J. M. Karczewski, G. J. E. Sharp, and C. J. Secombes. Susceptibility of strains of Aeromonas salmonicida to killing by cell-free generated superoxide anion. Journal of Fish Diseases, 14(3):367–373, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kagel:1993:HPI

[KT93a] M. Kagel and H. Taraschewski. Host-parasite interface of Diplozoon paradoxum (Monogenea) in naturally infected bream, Abramis brama (L.). Journal of Fish Diseases, 16(5): 501–506, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karanis:1993:HPI

[KT93b] P. Karanis and H. Taraschewski. Host-parasite interface of Caryophyllaeus laticeps (Eucestoda: Caryophyllidae) in three species of fish. Journal of Fish Diseases, 16(4):371–379, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kelly:1990:ELP

[KWJ90] D. Kelly, W. R. Wolters, and J. M. Jaynes. Effect of lytic peptides on selected fish bacterial pathogens. Journal of Fish Diseases, 13(4):317–321, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lecomte:1992:CPM

[LAB92] J. Lecomte, M. Arella, and L. Berthiaume. Comparison of polyclonal and monoclonal antibodies for serotyping infectious pancreatic necrosis virus (IPNV) strains isolated in eastern Canada. Journal of Fish Diseases, 15(5):431–436, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1991:DVP

[Lan91a] J. S. Langdon. Description of Vavraia parastacida sp. nov. (Microspora: Pleistophoridae) from marron, Cherax tenuimanus (Smith), (Decapoda: Parastacidae). Journal of Fish Diseases, 14(6):619–629, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1991:MDP

[Lan91b] J. S. Langdon. Microsporidiosis due to a pleistophorid in marron, Cherax tenuimanus (Smith), (Decapoda: Parastacidae). Journal of Fish Diseases, 14(1):33–44, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1991:MPM

[Lan91c] J. S. Langdon. Myoliquefaction post-mortem (‘milky flesh’) due to Kudoa thyrsites (Gilchrist) (Myxosporea: Multivalvulida) in mahi mahi, Coryphaena hippurus L. Journal of Fish Diseases, 14(1):45–54, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lindenstrom:1998:DTI

[LB98] Lindenstrøm and Buchmann. Dexamethasone treatment increases susceptibility of rainbow trout, Oncorhynchus mykiss (Walbaum), to infections with Gyrodactylus derjavini Mikailov. Journal of Fish Diseases, 21(1):29–38, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lobo-Da-Cunha:1992:VLP

[LDCA92] A. Lobo-Da-Cunha and C. Azevedo. Virus-like particles in the fish parasite Ichthyophthirius multifiliis (Ciliophora). Journal of Fish Diseases, 15(3):273–277, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1993:MID

[LDTC93] J. Lom, I. Dyková, K. Tonguthai, and S. Chinabut. Muscle infection due to Heterosporis sp. in the Siamese fighting fish, Betta splendens Regan. Journal of Fish Diseases, 16(5):513– 516, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:1991:RLE

[LE91a] K.-K. Lee and A. E. Ellis. The role of the lethal extracellular cytolysin of Aeromonas salmonicida in the pathology of furunculosis. Journal of Fish Diseases, 14(4):453–460, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:1991:SSE

[LE91b] K.-K. Lee and A. E. Ellis. Susceptibility of salmonid erythrocytes to lysis by bacterial lipopolysaccharides. Journal of Fish Diseases, 14(4):461–465, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leiro:1993:HIR

[LES^{+}93] J. Leiro, J. Estévez, M. T. Santamarina, M. L. Sanmartín, and F. M. Ubeira. Humoral immune response of turbot, Scophthalmus maximus (L.), to antigens from Tetramicra brevifilum Matthews & Matthews, 1980 (Microspora). Journal of Fish Diseases, 16(6):577–584, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lannan:1994:ESP

[LF94a] C. N. Lannan and J. L. Fryer. Extracellular survival of Piscirickettsia salmonis. Journal of Fish Diseases, 17(5):545–548, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lilley:1994:CRA

[LF94b] J. H. Lilley and G. N. Frerichs. Comparison of rhabdoviruses associated with epizootic ulcerative syndrome (EUS) with respect to their structural proteins, cytopathology and serology. Journal of Fish Diseases, 17(5):513–522, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Low:1998:ARA

[LGL^{+}98] Low, Goh, Lim, Sin, and Leung. Actin rearrangements accompanying Aeromonas hydrophila entry into cultured fish cells. Journal of Fish Diseases, 21(1):55–65, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LeBreton:1997:VNN

[LGSO97] A. Le Breton, L. Grisez, J. Sweetman, and F. Ollevier. Viral nervous necrosis (VNN) associated with mass mortalities in cage-reared sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 20(2):145–151, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewin:1990:VAO

[LH90] C. S. Lewin and T. S. Hastings. In vitro activities of oxolonic acid, ciprofloxacin and norfloxacin against Aeromonas salmonicida. Journal of Fish Diseases, 13(5):377–384, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lupiani:1994:IAP

[LHS94] B. Lupiani, F. M. Hetrick, and S. K. Samal. Identification of the angelfish, Pomacanthus semicirculatus aquareovirus as a member of aquareovirus genogroup A using RNA-RNA blot hybridization. Journal of Fish Diseases, 17(6):667–672, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lund:1991:HIR

[LJHE91] V. Lund, T. Jørgensen, K. O. Holm, and G. Eggset. Humoral immune response in Atlantic salmon, Salmo salar L., to cellular and extracellular antigens of Aeromonas salmonicida. Journal of Fish Diseases, 14(4):443–452, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:1999:GFN

[LKM99] N-S. Lee, J. Kobayashi, and T. Miyazaki. Gill filament necrosis in farmed Japanese eels, Anguilla japonica (Temminck & Schlegel), infected with Herpesvirus anguillae. Journal of Fish Diseases, 22(6):457–463, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Lastra:1994:DMI

[LLGJS94] M. Lopez-Lastra, M. Gonzalez, M. Jashes, and A. M. Sandino. A detection method for infectious pancreatic necrosis virus (IPNV) based on reverse transcription (RT)- polymerase chain reaction (PCR). Journal of Fish Diseases, 17(3):269–282, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leung:1995:IFP

[LLLS95] K. Y. Leung, K. W. Low, T. J. Lam, and Y. M. Sin. Interaction of the fish pathogen Aeromonas hydrophila with tilapia, Oreochromis aureus (Steindachner), phagocytes. Journal of Fish Diseases, 18(5):435–447, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leung:1996:MCC

[LLLS96] K. Y. Leung, T. M. Lim, T. J. Lam, and Y. M. Sin. Morphological changes in carp epithelial cells infected with Aeromonas hydrophila. Journal of Fish Diseases, 19(2):167– 174, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1994:SSR

[LLN94] Y. Lu, P. C. Loh, and E. C. B. Nadala. Serological studies of the rhabdovirus of penaeid shrimp (RPS) and its relationship to three other fish rhabdoviruses. Journal of Fish Diseases, 17(4):303–309, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lillehaug:1992:FTA

[LLP92] A. Lillehaug, T. Lunder, and T. T. Poppe. Field testing of adjuvanted furunculosis vaccines in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(6):485–496, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewis:1992:ETN

[LMBS92] E. J. Lewis, S. M. Mclaughlin, J. E. Bodammer, and T. K. Sawyer. Epitheliocystis in ten new host species of marine fish. Journal of Fish Diseases, 15(3):267–271, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lair:1997:JUA

[LMH^{+}97] S. Lair, Y. Mailhot, R. Higgins, D. B. élanger, L. Berthiaume, Y. De Lafontaine, and D. Martineau. Jaw ulcers in Atlantic tomcod, Microgadus tomcod (Walbaum), from the St Lawrence River. Journal of Fish Diseases, 20(1):11–17, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1995:BOS

[LMTW95] J. S. Langdon, A. Masters, T. Thorne, and S. Wilton. Bizarre organism from the skin of mahi mahi, Coryphaena hippurus L. (Teleostei: Coryphaenidae). Journal of Fish Diseases, 18 (6):481–494, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lamas:1995:PAV

[LNFT95] J. Lamas, M. Noya, A. Figueras, and A. E. Toranzo. Pathology associated with a viral erythrocytic infection in turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 18(5): 425–433, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:1996:NMC

[LOF96] J. S. Lumsden, V. E. Ostland, and H. W. Ferguson. Necrotic myositis in cage cultured rainbow trout, Oncorhynchus mykiss (Walbaum), caused by Flexibacter psychrophilus. Journal of Fish Diseases, 19(2):113–119, March 1996. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Larsen:1994:VAS

[LPD94] J. L. Larsen, K. Pedersen, and I. Dalsgaard. Vibrio anguillarum serovars associated with vibriosis in fish. Journal of Fish Diseases, 17(3):259–267, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leiro:1998:AAG

[LPO^{+}98] Leiro, Paniagua, Ortega, Paramá, Fernández, and Sanmartín. An amoeba associated with gill disease in turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 21(4):281–288, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leiro:1999:RGC

[LPO^{+}99] J. Leiro, A. Paramá, M. Ortega, M. T. Santamarina, and M. L. Sanmartín. Redescription of Glugea caulleryi, a microsporidian parasite of the greater sand-eel, Hyperoplus lanceolatus (Le Sauvage), (Teleostei: Ammodytidae), as Microgemma caulleryi comb. nov. Journal of Fish Diseases, 22 (2):101–110, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lilley:1997:PCS

[LR97] J. H. Lilley and R. J. Roberts. Pathogenicity and culture studies comparing the Aphanomyces involved in epizootic ulcerative syndrome (EUS) with other similar fungi. Journal of Fish Diseases, 20(2):135–144, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lasee:1993:BCT

[LS93] B. A. Lasee and D. R. Sutherland. Bacterial colonization of tegumental surfaces of Culaeatrema inconstans Lasee et al., 1988 (Digenea) from the brook stickleback, Culaea inconstans. Journal of Fish Diseases, 16(1):83–85, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Livesley:1999:CAS

[LSAB99] M. A. Livesley, S. N. Smith, R. A. Armstrong, and G. A. Barker. Characterisation of Aeromonas strains and species by pulsed field gel electrophoresis and principal components analysis. Journal of Fish Diseases, 22(5):369–375, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ling:1991:NAC

[LSL91] K. H. Ling, Y. M. Sin, and T. J. Lam. A new approach to controlling ichthyophthiriasis in a closed culture system of freshwater ornamental fish. Journal of Fish Diseases, 14(5): 595–598, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lillehaug:1990:CPA

[LSR90] A. Lillehaug, R. H. Sørum, and A. Ramstad. Cross-protection after immunization of Atlantic salmon, Salmo salar L., against different strains of Vibrio salmonicida. Journal of Fish Diseases, 13(6):519–523, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1992:ETP

[LT92] J. S. Langdon and T. Thorne. Experimental transmission per os of microsporidiosis due to Vavraia parastacida in the marron, Cherax tenuimanus (Smith), and yabby, Cherax albidus Clark. Journal of Fish Diseases, 15(4):315–322, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1994:IYH

[LTBL94] Y. Lu, L. M. Tapay, J. A. Brock, and P. C. Loh. Infection of the yellow head baculo-like virus (YBV) in two species of penaeid shrimp, Penaeus stylirostris (Stimpson) and Penaeus vannamei (Boone). Journal of Fish Diseases, 17(6):649–656, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1996:MML

[LTC96] C. C. Lu, K. F. J. Tang, and S. N. Chen. Morphogenesis of the membranous labyrinth in penaeid shrimp cells infected with Penaeus monodon bacuiovirus (MBV). Journal of Fish Diseases, 19(5):357–364, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1998:IGC

[LTC98] C. C. Lu, K. F. J. Tang, and S. N. Chen. Identification and genetic characterization of yeasts isolated from freshwater prawns, Macrobrachium rosenbergii de Man, in Taiwan. Journal of Fish Diseases, 21(3):185–192, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1992:RHN

[LTF92] J. S. Langdon, T. Thorne, and W. J. Fletcher. Reservoir hosts and new clupeoid host records for the myoliquefactive myxosporean parasite Kudoa thyrsites (Gilchrist). Journal of Fish Diseases, 15(6):459–471, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laidler:1999:AAS

[LTGC99] L. A. Laidler, J. W. Treasurer, A. N. Grant, and D. I. Cox. Atypical Aeromonas salmonicida infection in wrasse (Labridae) used as cleaner fish of farmed Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Diseases, 22(3):209– 213, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1993:DPM

[LTKC93] C. C. Lu, K. F. J. Tang, G. H. Kou, and S. N. Chen. Development of a Penaeus monodon-type baculovirus (MBV) DNA probe by polymerase chain reaction and sequence analysis. Journal of Fish Diseases, 16(6):551–559, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1995:DPM

[LTKC95] C. C. Lu, K. F. J. Tang, G. H. Kou, and S. N. Chen. Detection of Penaeus monodon-type baculovirus (MBV) infection in Penaeus monodon Fabricius by in situ hybridization. Journal of Fish Diseases, 18(4):337–345, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:1996:DNE

[LTL96] Y. Lu, L. M. Tapay, and P. C. Loh. Development of a nitrocellulose-enzyme immunoassay for the detection of yellow-head virus from penaeid shrimp. Journal of Fish Diseases, 19(1):9–13, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:1997:VRT

[LW97] S. Li and P. T. K. Woo. Vaccination of rainbow trout, Oncorhynchus mykiss (Walbaum), against cryptobiosis: efficacy of the vaccine in fresh and sea water. Journal of Fish Diseases, 20(5):369–374, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lachmann:1997:DEA

[LWHD97] I. Lachmann, U. Wagner, D. Hädge, and K. Drössler. The detection of esterolytically active glycerophospholipid: cholesterol acyltransferase of Aeromonas salmonicida in SDSpolyacrylamide gels using the α-naphthyl acetate–azo dye method. Journal of Fish Diseases, 20(6):459–461, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leff:1994:CIC

[LYD94] A. A. Leff, T. Yoshinaga, and H. W. Dickerson. Cross immunity in channel catfish, Ictalurus punctatus (Rafinesque), against two immobilization serotypes of Ichthyophthirius multifiliis (Fouquet). Journal of Fish Diseases, 17(4):429–432, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:1997:ICF

[LYGS97] S. Y. Lee, Z. Yin, R. Ge, and Y. M. Sin. Isolation and characterization of fish Aeromonas hydrophila adhesins important for in vitro epithelial cell invasion. Journal of Fish Diseases, 20(3):169–175, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leung:1995:SRG

[LYLS95] K. Y. Leung, I. V. Yeap, T. J. Lam, and Y. M. Sin. Serum resistance as a good indicator for virulence in Aeromonas hydrophila strains isolated from diseased fish in South-East Asia. Journal of Fish Diseases, 18(6):511–518, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Montero:1999:CEP

[MA99] A. B. Montero and B. Austin. Characterization of extracellular products from an isolate of Vibrio harveyi recovered from diseased post-larval Penaeus vannamei (Bonne). Journal of Fish Diseases, 22(5):377–386, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mair:1992:CDS

[Mai92] G. C. Mair. Caudal deformity syndrome (CDS): an autosomal recessive lethal mutation in the tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 15(1):71–75, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:1999:SHD

[MAR99] D. J. Morris, A. Adams, and R. H. Richards. In situ hybridization of DNA probes to PKX, the causative organism of proliferative kidney disease (PKD). Journal of Fish Diseases, 22(2):161–163, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mackie:1992:SPV

[MB92] C. Mackie and T. H. Birkbeck. Siderophores produced by Vibrio anguillarum in vitro and in infected rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 15(1):37–45, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Manera:1994:IHT

[MB94] M. Manera and S. Biavati. An immuno-histochemical technique used to demonstrate the transition form of a squamous cell carcinoma in a mirror carp, Cyprinus carpio L. Journal of Fish Diseases, 17(1):93–96, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Manera:1995:FTT

[MB95a] M. Manera and S. Biavati. A fibromyxoma in a tench, Tinca tinca L. Journal of Fish Diseases, 18(6):627–630, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matthews:1995:CIC

[MB95b] R. A. Matthews and P. J. Burgess. Cryptocaryon irritans (Ciliophora): primary infection in thick-lipped mullet, Chelon labrosus (Risso). Journal of Fish Diseases, 18(4):329–335, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McAllister:1997:IPN

[MB97] P. E. McAllister and J. Bebak. Infectious pancreatic necrosis virus in the environment: relationship to effluent from aquaculture facilities. Journal of Fish Diseases, 20(3):201– 207, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mellergaard:1998:FGI

[MB98] S. Mellergaard and O. Bagge. Fishing gear-induced skin ulcerations in Baltic cod, Gadus morhua L. Journal of Fish Diseases, 21(3):205–213, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Masoumian:1996:DMB

[MBM96] M. Masoumian, F. Baska, and K. Malnár. Description of Myxobolus bulbocordis sp. nov. (Myxosporea: Myxobolidae) from the heart of Barbus sharpeyi (Günther) and histopathological changes produced by the parasite. Journal of Fish Diseases, 19(1):15–21, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Melby:1994:AAR

[MC94] H. P. Melby and K. E. Christie. Antigenic analysis of reference strains and Norwegian field strains of aquatic birnaviruses by the use of six monoclonal antibodies produced against the infectious pancreatic necrosis virus N1 strain. Journal of Fish Diseases, 17(4):409–415, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mashima:1997:AJD

[MCL97] T. Y. Mashima, T. E. Cornish, and G. A. Lewbart. Amyloidosis in a jack Dempsey cichild, Cichlasoma biocellatum Regan. Journal of Fish Diseases, 20(1):73–75, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Melby:1994:AAN

[MCRF94] H. P. Melby, P. Caswell-Reno, and K. Falk. Antigenic analysis of Norwegian aquatic birnavirus isolates using monoclonal antibodies — with special reference to fish species, age and health status. Journal of Fish Diseases, 17(1):85–91, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maclean:1990:PDI

[MD90] S. A. Maclean and A. J. Davies. Prevalence and development of intraleucocytic haemogregarines from Northwest and Northeast Atlantic mackerel, Scomber scombrus L. Journal of Fish Diseases, 13(1):59–68, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martinez:1994:FIS

[MEJ^{+}94] I. Martinez, S. Espelid, A. Johansen, J. Welsh, and M. Mcclelland. Fast identification of species and strains of Vibrio by amplification of polymorphic DNA. Journal of Fish Diseases, 17(3):297–302, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1999:EIA

[MEMSB99] K. Molnár, A. El-Mansy, Cs Székely, and F. Baska. Experimental identification of the actinosporean stage of Sphaerospora renicola Dykova & Lom 1982 (Myxosporea: Sphaerosporidae) in oligochaete alternate hosts. Journal of Fish Diseases, 22(2):143–153, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Michel:1991:CSF

[MF91] C. Michel and B. Faivre. Carrier state in furunculosis: secondary infection of trout with different Aeromonas salmonicida strains results in advantage for the primarily harboured one. Journal of Fish Diseases, 14(5):571–576, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Melby:1995:SIB

[MF95] H. P. Melby and K. Falk. Study of the interaction between a persistent infectious pancreatic necrosis virus (IPNV) infection and experimental infectious salmon anaemia (ISA) in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18 (6):579–586, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maheshkumar:1990:CDA

[MGE90] S. Maheshkumar, S. M. Goyal, and P. P. Economon. Concentration and detection of Aeromonas salmonicida from hatchery water. Journal of Fish Diseases, 13(6):513–518, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martinsen:1994:MDP

[MHB94] B. Martinsen, T. E. Horsberg, and M. Burke. Multiple-dose pharmacokinetic and depletion studies of sarafloxacin in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 17 (2):111–121, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccormick:1995:IMC

[MHM95] J. I. Mccormick, M. S. Hughes, and M. F. Mcloughlin. Identification of Mycobacterium chelonae in a cichlid oscar, Astronotus ocellatus Cuvier, by direct cycle sequencing of polymerase chain reaction amplified 16s rRNA gene sequences. Journal of Fish Diseases, 18(5):459–461, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulvey:1995:EPA

[MLB95] B. Mulvey, M. L. Landolt, and R. A. Busch. Effects of potassium aluminium sulphate (alum) used in an Aeromonas salmonicida bacterin on Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(6):495–506, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matthews:1993:CIB

[MMB93] B. F. Matthews, R. A. Matthews, and P. J. Burgess. Cryptocaryon irritans Brown, 1951 (Ichthyophthiriidae): the ultrastructure of the somatic cortex throughout the life cycle. Journal of Fish Diseases, 16(4):339–349, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miyata:1997:GSI

[MMJ^{+}97] M. Miyata, K. Matsuno, S. J. Jung, Y. Danayadol, and T. Miyazaki. Genetic similarity of iridoviruses from Japan and Thailand. Journal of Fish Diseases, 20(2):127–134, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcallister:1995:AVF

[MO95] P. E. Mcallister and W. J. Owens. Assessment of the virulence of fish and mouuscan isolates of infectious pancreatic necrosis virus for salmonid fish by challenge of brook trout, Salvelinus fontinalis (Mitchill). Journal of Fish Diseases, 18(1):97–103, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McGrogan:1998:SDI

[MOBF98] McGrogan, Ostland, Byrne, and Ferguson. Systemic disease involving an iridovirus-like agent in cultured tilapia, Oreochromis niloticus L. — a case report. Journal of Fish Diseases, 21(2):149–152, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moyner:1993:CSP

[Møy93] K. Møyner. Changes in serum protein composition occur in Atlantic salmon, Salmo salar L., during Aeromonas salmonicida infection. Journal of Fish Diseases, 16(6):601–604, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murphy:1992:SPP

[MRD^{+}92] T. M. Murphy, H. D. Rodger, E. M. Drinan, F. Gannon, P. Kruse, and W. Korting. The sequential pathology of pancreas disease in Atlantic salmon farms in Ireland. Journal of Fish Diseases, 15(5):401–408, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McLoughlin:1998:SSS

[MRD98] McLoughlin, Rowley, and Doherty. A serological survey of salmon pancreas disease virus (SPDV) antibodies in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21 (4):305–307, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morita:1990:REC

[MS90] N. Morita and T. Sano. Regression effect of carp, Cyprinus carpio L., peripheral blood lymphocytes on CHV-induced carp papilloma. Journal of Fish Diseases, 13(6):505–511, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morales:1991:SCT

[MS91] P. Morales and R. E. Schmidt. Spindle-cell tumour resembling haemangiopericytoma in a common goldfish, Carassius auratus (L.). Journal of Fish Diseases, 14(4):499–502, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meinertz:1995:RMG

[MSGA95] J. R. Meinertz, G. R. Stehly, W. H. Gingerich, and J. L. Allen. Residues of [^{14}C]-malachite green in eggs and fry of rainbow trout, Oncorhynchus mykiss (Walbaum), after treatment of eggs. Journal of Fish Diseases, 18(3):239–248, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marshall:1990:IRS

[MSM^{+}90] S. H. Marshall, S. K. Samal, T. H. Mcphillips, A. R. Moore, and F. M. Hetrick. Isolation of a rotavirus from smelt, Osmerus mordax (Mitchell). Journal of Fish Diseases, 13(1):87– 91, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcgeorge:1994:LEM

[MSW94] J. Mcgeorge, C. Sommerville, and R. Wootten. Light and electron microscope observations on extrasporogonic and sporogonic stages of a myxosporean parasite of the genus Sphaerospora Thelohan, 1892 from Atlantic salmon, Salmo solar L., in Scotland. Journal of Fish Diseases, 17(3):227– 238, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McAndrew:1998:EHP

[MSWB98] K. J. McAndrew, C. Sommerville, R. Wootten, and J. E. Bron. The effects of hydrogen peroxide treatment on different life-cycle stages of the salmon louse, Lepeophtheirus salmonis (Krøyer 1837). Journal of Fish Diseases, 21(3):221–228, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Majeed:1994:TGB

[MW94] S. K. Majeed and D. S. Wang. Tumours in the gonads of the bagrid catfish, Mystus macropterus (Bleeker). Journal of Fish Diseases, 17(5):527–532, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noterdaeme:1991:BPC

[NBWO91] L. Noterdaeme, S. Bigawa, K. A. Willems, and F. Ollevier. Biochemical and physiological characteristics and plasmid profiles of Aeromonas hydrophila strains, isolated from freshwater fish and from fresh water. Journal of Fish Diseases, 14 (3):313–321, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:1999:PEA

[NC99] B. F. Nowak and A. Clark. Prevalence of epitheliocystis in Atlantic salmon, Salmo salar L., farmed in Tasmania, Australia. Journal of Fish Diseases, 22(1):73–78, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nielsen:1994:ASS

[NDBL94] B. Nielsen, I. Dalsgaard, D. J. Brown, and J. L. Larsen. Aeromonas salmonicida subsp. salmonicida: correlation of protein patterns, antibiotic resistance, exoprotease activity, haemolysis and pathological lesions produced in vivo. Journal of Fish Diseases, 17(4):387–397, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nieto:1995:PES

[NDQT95] J. M. Nieto, S. Devesa, I. Quiroga, and A. E. Toranzo. Pathology of Enterococcus sp. infection in farmed turbot, Scophthalmus maximus L. Journal of Fish Diseases, 18(1):21–30, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nieto:1991:HEP

[NE91] T. P. Nieto and A. E. Ellis. Heterogeneity of extracellular proteases produced by different isolates of Aeromonas hydrophila and A. sobria pathogenic for fish. Journal of Fish Diseases, 14(2):229–235, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nerland:1996:NSG

[Ner96] A. H. Nerland. The nucleotide sequence of the gene encoding GCAT from Aeromonas salmonicida ssp. salmonicida. Journal of Fish Diseases, 19(2):145–150, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nelson:1999:IIP

[NG99] E. J. Nelson and W. C. Ghiorse. Isolation and identification of Pseudoalteromonas piscicida strain Cura-d associated with diseased damselfish (Pomacentridae) eggs. Journal of Fish Diseases, 22(4):253–260, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nerland:1993:ASS

[NHOJ93] A. H. Nerland, B. T. Høgh, A. B. Olsen, and H. B. Jensen. Aeromonas salmonicida ssp. salmonicida requires exogenous arginine and methionine for growth. Journal of Fish Diseases, 16(6):605–608, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nylund:1995:PIS

[NHWE95] A. Nylund, T. Hovland, K. Watanabe, and C. Endresen. Presence of infectious salmon anaemia virus (ISAV) in tissues of Atlantic salmon, Salmo salar L., collected during three separate outbreaks of the disease. Journal of Fish Diseases, 18 (2):135–145, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Neukirch:1993:EVI

[NK93] M. Neukirch and P. Kruse. Enhancement of the in vitro infectivity of V 834 in the presence of trypsin. Journal of Fish Diseases, 16(2):155–159, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nylund:1997:RIS

[NKKH97] A. Nylund, A. M. Kvenseth, B. Krossøy, and K. Hodneland. Replication of the infectious salmon anaemia virus (ISAV) in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 20(4):275–279, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nieto:1990:ISP

[NLS^{+}90] T. P. Nieto, L. R. López, Y. Santos, S. Núñez, and A. E. Toranzo. Isolation of Serratia plymuthica as an opportunistic pathogen in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 13(2):175–177, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nilsem:1992:TII

[NLTN92] H. Nilsem, A. Lillehaug, T. Taksdal, and R. Nordmo. Toxicity of intraperitoneally injected formalin in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15(4):323–329, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nordmo:1997:CDC

[NR97] R. Nordmo and A. Ramstad. Comparison of different challenge methods to evaluate the efficacy of furunculosis vaccines in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 20(2):119–126, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nordmo:1998:EFA

[NRV^{+}98] Nordmo, Holth Riseth, Varma, Sutherland, and Brokken. Evaluation of florfenicol in Atlantic salmon, Salmo salar L.: efficacy against furunculosis due to Aeromonas salmonicida and cold water vibriosis due to Vibrio salmonicida. Journal of Fish Diseases, 21(4):289–297, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nakajima:1996:PMA

[NS96] K. Nakajima and M. Sorimachi. Production of monoclonal antibodies against yellowtail ascites virus. Journal of Fish Diseases, 19(2):161–166, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nemcsok:1993:RFR

[NVBK93] J. Nemcsók, É. Vig, F. Baska, and O. Kufcsák. Role of free radicals in swimbladder inflammation of carp. Journal of Fish Diseases, 16(2):131–136, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nordmo:1994:FAS

[NVSB94] R. Nordmo, K. J. Varma, I. H. Sutherland, and E. S. Brokken. Florfenicol in Atlantic salmon, Salmo salar L.: field evaluation of efficacy against furunculosis in Norway. Journal of Fish Diseases, 17(3):239–244, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostland:1999:ABG

[OBL^{+}99] V. E. Ostland, P. J. Byrne, J. S. Lumsden, D. D. MacPhee, J. A. Derksen, M. Haulena, K. Skar, E. Myhr, and H. W. Ferguson. Atypical bacterial gill disease: a new form of bacterial gill disease affecting intensively reared salmonids. Journal of Fish Diseases, 22(5):351–358, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

OHalloran:1995:PAS

[OCC^{+}95] J. O’Halloran, K. Coombs, E. Carpenter, K. Whitman, and G. Johnson. Prescreening Atlantic salmon, Salmo salar L., broodstock for the presence of Renibacterium salmoninarum by indirect fluorescent antibody testing. Journal of Fish Diseases, 18(1):83–85, January 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olivier:1992:UWB

[OGFL92] G. Olivier, S. G. Griffiths, J. Fildes, and W. H. Lynch. The use of Western blot and electroimmunotransfer blot assays to monitor bacterial kidney disease in experimentally challenged Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 15 (3):229–241, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oisson:1996:TSM

[OJW^{+}96] J. C. Oisson, A. Jöborn, A. Westerdahl, L. Blomberg, S. Kjelleberg, and P. L. Conway. Is the turbot, Scophthalmus maximus (L.), intestine a portal of entry for the fish pathogen Vibrio anguillarum? Journal of Fish Diseases, 19 (3):225–234, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olsson:1998:SPP

[OJW^{+}98] Olsson, Jöborn, Westerdahl, Blomberg, Kjelleberg, and Conway. Survival, persistence and proliferation of Vibrio anguillarum in juvenile turbot, Scophthalmus maximus (L.), intestine and faeces. Journal of Fish Diseases, 21(1):1–9, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostland:1997:IAF

[OLM^{+}97] V. E. Ostland, J. S. Lumsden, D. D. MacPhee, J. A. Derksen, and H. W. Ferguson. Inhibition of the attachment of Flavobacterium branchiophilum to the gills of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 20(2):109–117, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostland:1997:CON

[OMF97] V. E. Ostland, D. G. McGrogan, and H. W. Ferguson. Cephalic osteochondritis and necrotic scleritis in intensively reared salmonids associated with Flexibacter psychrophilus. Journal of Fish Diseases, 20(6):443–451, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostland:1995:VFB

[OMLF95] V. E. Ostland, D. D. Macphee, J. S. Lumsden, and H. W. Ferguson. Virulence of Flavobacterium branchiophilum in experimentally infected salmonids. Journal of Fish Diseases, 18 (3):249–262, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oppegaard:1995:VAA

[OSG95] H. Oppegaard, H. Sørum, and K. Gravningen. In vitro antimicrobial activity of sulphachlorpyridazine and sulphadiazine in combination with trimethoprim against bacterial fish pathogens. Journal of Fish Diseases, 18(4):371–376, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Price:1990:SMK

[PBC^{+}90] N. C. Price, R. Margaret Banks, Carole M. Campbell, Doris Duncan, and L. Stevens. The specificity of the major (70 kDa) protease secreted by Aeromonas salmonicida. Journal of Fish Diseases, 13(1):49–58, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Penrith:1994:HLF

[PBP94] M.-L. Penrith, S. S. Bastianello, and M. J. Penrith. Hepatic lipoidosis and fatty infiltration of organs in a captive African stonefish, Synanceja verrucosa Bloch & Schneider. Journal of Fish Diseases, 17(2):171–176, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pearson:1994:JDF

[PCKS94] M. D. Pearson, S. Chinabut, S. Karnchanakharn, and T. Somsiri. Jaundice disease in the farmed catfish hybrid, Clarias macrocephalus (Gunther) × C. gariepinus (Burchell), in Thailand. Journal of Fish Diseases, 17(4):325–336, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Press:1996:RFV

[PERL96] C. McL. Press, Ø. Evensen, L. J. Reitan, and T. Landsverk. Retention of furunculosis vaccine components in Atlantic salmon, Salmo salar L., following different routes of vaccine administration. Journal of Fish Diseases, 19(3):215–224, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paniagua:1998:ETS

[PFO^{+}98] Paniagua, Fernández, Ortega, Paramá, Sanmartín, and Leiro. Effects of temperature, salinity and incubation time on in vitro survival of an amoeba infecting the gills of turbot, Scophthalmus maximus L. Journal of Fish Diseases, 21(1): 77–80, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Petervary:1996:SNR

[PGLH96] N. Petervary, D. M. Gillette, G. A. Lewbart, and J. C. Harshbarger. A spontaneous neoplasm of the renal collecting ducts in an oscar, Astronotus ocellatus (Cuvier), with comments on similar cases in this species. Journal of Fish Diseases, 19 (4):279–281, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Philbey:1991:CDG

[PI91] A. W. Philbey and B. A. Ingram. Coccidiosis due to Goussia lomi (Protista: Apicomplexa) in aquarium-reared Murray cod, Maccullochella peeli (Mitchell), (Percichthyidae). Journal of Fish Diseases, 14(2):237–242, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pearson:1993:SMD

[PI93] M. Pearson and V. Inglis. A sensitive microbioassay for the detection of antibacterial agents in the aquatic environment. Journal of Fish Diseases, 16(3):255–260, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pybus:1994:GIS

[PLLT94] V. Pybus, M. W. Loutit, I. L. Lamont, and J. R. Tagg. Growth inhibition of the salmon pathogen Vibrio ordalii by a siderophore produced by Vibrio anguillarum strain VL4355. Journal of Fish Diseases, 17(4):311–324, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulsford:1991:MGC

[PM91] A. Pulsford and R. A. Matthews. Macrophages and giant cells associated with a microsporidian parasite causing liquefaction of the skeletal muscle of the Norway pout, Trisopterus esmarkii (Nilsson). Journal of Fish Diseases, 14(1):67–78, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Power:1990:UTD

[PMD90] C. A. Power, R. J. F. Markham, and A. W. Donald. Uptake and tissue distribution of liposomes after intraperitoneal administration to rainbow trout, Oncorhynchus mykiss (Richardson): a preliminary report. Journal of Fish Diseases, 13(4):329–332, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:1998:EAO

[PMS98] Poppe, Midtlyng, and Sande. Examination of abdominal organs and diagnosis of deficient septum transversum in Atlantic salmon, Salmo salar L., using diagnostic ultrasound imaging. Journal of Fish Diseases, 21(1):67–72, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peleteiro:1990:PCE

[PR90] M. C. Peleteiro and R. H. Richards. Phagocytic cells in the epidermis of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 13(3):225–232, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pickup:1996:PNC

[PRCC96] R. W. Pickup, G. Rhodes, R. J. Cobban, and K. J. Clarke. The postponement of non-culturability in Aeromonas salmonicida. Journal of Fish Diseases, 19(1):65–74, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Press:1995:ARE

[PRL95] C. Mcl. Press, L. J. Reitan, and T. Landsverk. Antigen retention and enzyme reactivity in the spleen of Atlantic salmon, Salmo salar L., following administration of injectable furunculosis vaccines. Journal of Fish Diseases, 18(3):199–210, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Prasad:1993:HSC

[PS93] M. Prasad and M. Shil. Histopathology of skin of the catfish, Heteropneustes fossilis: short-term effects of mercuric chloride. Journal of Fish Diseases, 16(2):143–147, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pope:1998:CCV

[PS98] R. K. Pope and R. W. Scheetz. Channel catfish virus: interaction with the cytoskeleton. Journal of Fish Diseases, 21(6): 433–442, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Palmer:1992:MAS

[PSB^{+}92] R. Palmer, R. H. Soutar, E. J. Branson, P. J. Southgate, E. Drinan, R. H. Richards, and R. O. Collins. Mortalities in Atlantic salmon, Salmo salar L., associated with pathology of the melano-macrophage and haemopoietic systems. Journal of Fish Diseases, 15(2):207–210, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Park:1993:IHN

[PSL^{+}93] M. A. Park, S. G. Sohn, S. D. Lee, S. K. Chun, J. W. Park, J. L. Fryer, and Y. C. Hah. Infectious haematopoietic necrosis virus from salmonids cultured in Korea. Journal of Fish Diseases, 16(5):471–478, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pazos:1996:EMS

[PSM^{+}96] F. Pazos, Y. Santos, A. R. Macías, S. Núñez, and A. E. Toranzo. Evaluation of media for the successful culture of Flexibacter maritimus. Journal of Fish Diseases, 19(2):193– 197, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plumb:1993:VCC

[PV93] J. A. Plumb and S. Vinitnantharat. Vaccination of channel catfish, Ictalurus punctatus (Rafinesque), by immersion and oral booster against Edwardsiella ictaluri. Journal of Fish Diseases, 16(1):65–71, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:1993:ORL

[RD93] H. D. Rodger and E. M. Drinan. Observation of a rickettsialike organism in Atlantic salmon, Salmo salar L., in Ireland. Journal of Fish Diseases, 16(4):361–369, July 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rowley:1998:ISP

[RDMW98] H. M. Rowley, C. E. Doherty, M. F. McLoughlin, and M. D. Welsh. Isolation of salmon pancreas disease virus (SPDV) from farmed Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Diseases, 21(6):469–472, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rose:1990:SAS

[REM90] A. S. Rose, A. E. Ellis, and A. L. S. Munro. The survival of Aeromonas salmonicida subsp. salmonicida in sea water. Journal of Fish Diseases, 13(3):205–214, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodriguez:1993:PLA

[RFN93] L. A. Rodriguez, A. I. G. Fernandez, and T. P. Nieto. Production of the lethal acetylcholinesterase toxin by different Aeromonas hydrophila strains. Journal of Fish Diseases, 16 (1):73–78, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodriguez:1998:HAO

[RGA^{+}98] L. A. Rodríguez, C. S. Gallardo, F. Acosta, T. P. Nieto, B. Acosta, and F. Real. Hafnia alvei as an opportunistic pathogen causing mortality in brown trout, Salmo trutta L. Journal of Fish Diseases, 21(5):365–370, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reddacliff:1993:MGC

[RHC^{+}93] G. L. Reddacliff, M. Hornitzky, J. Carson, R. Petersen, and R. Zelski. Mortalities of goldfish, Carassius auratus (L.), associated with Vibrio cholerae (non-01) infection. Journal of Fish Diseases, 16(5):517–520, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Richards:1994:UOC

[RHL^{+}94] D. T. Richards, D. Hoole, J. W. Lewis, E. Ewens, and C. Arme. Ultrastructural observations on the cellular response of carp, Cyprinus carpio L., to eggs of the blood fluke Sanguinicola inermis Plehn, 1905 (Trematoda: Sanguinicolidae). Journal of Fish Diseases, 17(5):439–446, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ritchie:1997:HTA

[Rit97] G. Ritchie. The host transfer ability of Lepeophtheirus salmonis (Copepoda: Caligidae) from farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 20(2):153–157, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruangpan:1991:VBI

[RK91] L. Ruangpan and T. Kitao. Vibrio bacteria isolated from black tiger shrimp, Penaeus monodon Fabricius. Journal of Fish Diseases, 14(3):383–388, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:1997:SII

[RKMF97] H. D. Rodger, M. Kobs, A. Macartney, and G. N. Frerichs. Systemic iridovirus infection in freshwater angelfish, Pterophyllum scalare (Lichtenstein). Journal of Fish Diseases, 20 (1):69–72, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rand:1992:EII

[RM92] T. G. Rand and D. Munden. Enzyme involvement in the invasion of brook char, Salvelinus fontinalis (Mitchill), eggs by Saprolegnia diclina (Oomycotina: Saprolegniaceae). Journal of Fish Diseases, 15(1):91–94, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ravaglia:1995:MMC

[RM95a] M. A. Ravaglia and M. C. Maggese. Melano-macrophage centres in the gonads of the swamp eel, Synbranchus marmoratus Bloch, (Pisces, Synbranchidae): histological and histochemical characterization. Journal of Fish Diseases, 18(2):117–125, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ross:1995:SAV

[RM95b] K. Ross and A. L. S. Munro. Serological analysis of virulent infectious pancreatic necrosis virus strains. Journal of Fish Diseases, 18(4):377–380, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romalde:1994:IYR

[RMP^{+}94] J. L. Romalde, B. Magariños, F. Pazos, A. Silva, and A. E. Toranzo. Incidence of Yersinia ruckeri in two farms in Galicia (NW Spain) during a one-year period. Journal of Fish Diseases, 17(5):533–539, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruane:1995:ARC

[RMR95] N. Ruane, T. K. Mccarthy, and P. Reilly. Antibody response to crustacean ectoparasites in rainbow trout, Oncorhynchus mykiss (Walbaum), immunized with Argulus foliaceus L. antigen extract. Journal of Fish Diseases, 18(6): 529–537, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Randi:1996:HEC

[RMRR96] A. S. Randi, J. M. Monserrat, E. M. Rodriguez, and L. A. Romano. Histopathological effects of cadmium on the gills of the freshwater fish, Macropsobrycon uruguayanae Eigenmann (Pisces, Atherinidae). Journal of Fish Diseases, 19(4):311– 322, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rougier:1992:VEZ

[RMT^{+}92] F. Rougier, A. Menudier, D. Troutaud, C. Bosgiraud, A. Ndoye, J. A. Nicolas, and P. Deschaux. In vivo effect of zinc and copper on the development of listeriosis in zebrafish, Brachydanio rerio (Hamilton-Buchanan). Journal of Fish Diseases, 15(5):453–456, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodgers:1991:UVA

[Rod91] C. J. Rodgers. The usage of vaccination and antimicrobial agents for control of Yersinia ruckeri. Journal of Fish Diseases, 14(3):291–301, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1994:HEA

[Rou94a] F. R. Roubal. Histopathological and ecological aspects of Henneguya and Myxobolus (Myxosporea) infections in Acanthopagrus australis (Günther) (Pisces: Sparidae) from Moreton Bay, Australia. Journal of Fish Diseases, 17(5):495–512, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1994:HCC

[Rou94b] F. R. Roubal. Histopathology caused by Caligus epidemicus Hewitt (Copepoda: Caligidae) on captive Acanthopagrus australis (Günther) (Pisces: Sparidae). Journal of Fish Diseases, 17(6):631–640, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodak:1993:ELI

[RPT^{+}93] L. Rodák, Z. Pospíšil, J. Tománek, T. Veselý, T. Obr, and L. Valíček. Enzyme-linked immunosorbent assay (ELISA) for the detection of spring viraemia of carp virus (SVCV) in tissue homogenates of the carp, Cyprinus carpio L. Journal of Fish Diseases, 16(2):101–111, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1992:ICP

[RQW92] F. R. Roubal, N. Quartararo, and A. West. Infection of captive Pagrus auratus (Bloch & Schneider) by the monogenean, Anoplodiscus cirrusspiralis Roubal, Armitage & Rohde (Anoplodiscidae) in Australia. Journal of Fish Diseases, 15 (5):409–415, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:1998:OSE

[RR98] Rodger and Richards. Observational study of erythrocytic inclusion bodies in farmed Atlantic salmon, Salmo salar L., in the British Isles. Journal of Fish Diseases, 21(2):101–111, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Robertsen:1990:ENS

[RRER90] B. Robertsen, G. Rørstad, R. Engstad, and J. Raa. Enhancement of non-specific disease resistance in Atlantic salmon, Salmo salar L., by a glucan from Saccharomyces cerevisiae cell walls. Journal of Fish Diseases, 13(5):391–400, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramasamy:1996:LCC

[RRJ^{+}96] P. Ramasamy, P. R. Rajan, R. Jayakumar, S. Rani, and G. P. Brennan. Lagenidium callinectes (Couch, 1942) infection and its control in cultured larval Indian tiger prawn, Penaeus monodon Fabricius. Journal of Fish Diseases, 19(1):75–82, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roth:1993:CPC

[RRS93] M. Roth, R. H. Richards, and C. Sommerville. Current practices in the chemotherapeutic control of sea lice infestations in aquaculture: a review. Journal of Fish Diseases, 16(1):1– 26, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rosjo:1993:GCA

[RSTE93] C. Røsjø, R. Salte, M. S. Thomassen, and G. Eggset. Glycerophospholipid:cholesterol acyltransferase complexed with lipopolysaccharide (GCAT-LPS) of Aeromonas salmonicida produces lysophospholipids in salmonid red cell membranes: a probable haemolytic mechanism. Journal of Fish Diseases, 16(2):87–99, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reitan:1994:VRS

[RT94] L. J. Reitan and A. Thuvander. In vitro responsiveness of subpopulations of leucocytes from Atlantic salmon, Salmo salar L., to Aeromonas salmonicida. Journal of Fish Diseases, 17 (5):549–553, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:1994:CBG

[RTEC94] H. D. Rodger, T. Turnbull, C. Edwards, and G. A. Codd. Cyanobacterial (blue-green algal) bloom associated pathology in brown trout, Salmo trutta L., in Loch Leven, Scotland. Journal of Fish Diseases, 17(2):177–181, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rintamaki:1991:ASF

[RV91] P. Rintamäki and E. T. Valtonen. Aeromonas salmonicida in Finland: pathological problems associated with atypical and typical strains. Journal of Fish Diseases, 14(3):323–331, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rajendran:1998:CMP

[RVA98] K. V. Rajendran, K. K. Vijayan, and S. V. Alavandi. Cardiac myxosporiosis of pearl-spot, Etroplus suratensis (Bloch), due to Myxobolus etropli sp. nov. Journal of Fish Diseases, 21 (3):169–176, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rajendran:1999:EHR

[RVSK99] K. V. Rajendran, K. K. Vijayan, T. C. Santiago, and R. M. Krol. Experimental host range and histopathology of white spot syndrome virus (WSSV) infection in shrimp, prawns, crabs and lobsters from India. Journal of Fish Diseases, 22 (3):183–191, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1993:MAE

[RWC93] R. J. Roberts, L. G. Willoughby, and S. Chinabut. Mycotic aspects of epizootic ulcerative syndrome (EUS) of Asian fishes. Journal of Fish Diseases, 16(3):169–183, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stoffregen:1990:MPS

[SA90] D. A. Stoffregen and W. I. Anderson. A myxosporidian parasite in the skeletal muscle of a black-tip reef shark, Carcharhinus melanopterus (Quoy and Gaimard, 1824). Journal of Fish Diseases, 13(6):549–552, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shankar:1994:EAI

[SA94] K. M. Shankar and T. Yam Amoto. Epitope analysis of infectious pancreatic necrosis virus (IPNV) isolated from lake trout, Salvelinus namaycush (Walbaum), by monoclonal antibodies. Journal of Fish Diseases, 17(5):471–482, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Starliper:1995:RRS

[SA95] C. E. Starliper and J. D. Tesk A. Relevance of Renibacterium salmoninarum in an asymptomatic carrier population of brook trout, Salvelinus fontinalis (Mitchill). Journal of Fish Diseases, 18(5):383–387, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saeed:1990:HPP

[SAAH90] M. O. Saeed, M. M. Alamoudi, and A. H. Al-Harbi. Histopathology of Pseudomonas putrefaciens associated with disease in cultured rabbitfish, Siganus rivulatus (Forsskål). Journal of Fish Diseases, 13(5):417–422, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stensvaag:1999:LPS

[SABòJ99] K. Stensvåg, S. M. Arnesen, J. Bøgwald, and T. ò. Jørgensen. Localization of purified surface molecules (lipopolysaccharide and A-protein) from Aeromonas salmonicida in the kidney and spleen from Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 22(5):337–349, September 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1991:DSA

[SAK91] M. Sakai, S. Atsuta, and M. Kobatachi. The detection of serum antibody to Renibacterium salmoninarum in pencultured Coho salmon, Oncorhynchus kisutch (Walbaum). Journal of Fish Diseases, 14(2):243–246, March 1991. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1997:BLA

[SAMP97] D. J. Speare, G. Arsenault, N. MacNair, and M. D. Powell. Branchial lesions associated with intermittent formalin bath treatment of Atlantic salmon, Salmo salar L., and rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 20(1):27–33, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sima:1994:HIC

[SASG94] A. R. Sima, M. L. Aguirre, T. Scholz, and R. J. Guemez. Histopathology of the intestine of Cichlasoma urophthalmus (Günther) infected with metacercariae of Oligogonotylus manteri Watson, 1976 (Digenea: Cryptogonimidae). Journal of Fish Diseases, 17(5):523–526, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Slotved:1993:ARE

[SB93] H. C. Slotved and K. Buchmann. Acquired resistance of the eel, Anguilla anguilla L., to challenge infections with gill monogeneans. Journal of Fish Diseases, 16(6):585–591, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sarcione:1994:ESL

[SB94] E. J. Sarcione and J. J. Black. Elevated serum levels of alpha foetoprotein (AFP)-like immunoreactivity in rainbow trout, Oncorhynchus mykiss (Walbaum), with aflatoxin B_{1} induced hepatocellular carcinoma. Journal of Fish Diseases, 17(3):219–226, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sitja-Bobadilla:1990:STS

[SBAP90] A. Sitja-Bobadilla and P. Alvarez-Pellitero. Sphaerospora testicularis sp. nov. (Myxosporea: Sphaerosporidae) in wild and cultured sea bass, Dicentrarchus labrax (L.), from the Spanish Mediterranean area. Journal of Fish Diseases, 13(3):193– 203, May 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanz:1993:MAA

[SBB^{+}93] F. Sanz, B. Basurco, M. Babin, J. Dominguez, and J. M. Coll. Monoclonal antibodies against the structural proteins of viral haemorrhagic septicaemia virus isolates. Journal of Fish Diseases, 16(1):53–63, January 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:1992:IPN

[SBD^{+}92] D. A. Smail, D. W. Bruno, G. Dear, L. A. Mcfarlane, and K. Ross. Infectious pancreatic necrosis (IPN) virus Sp serotype in farmed Atlantic salmon, Salmo salar L., postsmolts associated with mortality and clinical disease. Journal of Fish Diseases, 15(1):77–83, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1999:EWT

[SBD99] D. J. Speare, H. J. Beaman, and J. Daley. Effect of water temperature manipulation on a thermal unit predictive model for Loma salmonae. Journal of Fish Diseases, 22(4):277–283, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sitja-bobadilla:1992:SMB

[SbFSAP92] A. Sitjà-bobadilla, A. Franco-Sierra, and P. Alvarez-Pellitero. Sphaerospora (Myxosporea: Bivalvulida) infection in cultured gilt head sea bream, Sparus aurata L.: a preliminary report. Journal of Fish Diseases, 15(4):339–343, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1998:IRR

[SBJ^{+}98] Speare, Beaman, Jones, Markham, and Arsenault. Induced resistance in rainbow trout, Oncorhynchus mykiss (Walbaum), to gill disease associated with the microsporidian gill parasite Loma salmonae. Journal of Fish Diseases, 21(2):93– 100, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1995:ETS

[SBMB95] D. J. Speare, J. Brocklebank, N. Macnair, and K. A. Bernard. Experimental transmission of a salmonid Rhodococcus sp. isolate to juvenile Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(6):587–597, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Santos:1991:PTS

[SBN^{+}91] Y. Santos, I. Bandin, S. Nuñez, K. Gravningen, and A. E. Toranzo. Protection of turbot, Scophthalmus maximus (L.), and rainbow trout, Oncorhynchus mykiss (Richardson), against vibriosis using two different vaccines. Journal of Fish Diseases, 14(3):407–411, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sitja-Bobadilla:1994:ELI

[SBW94] A. Sitja-Bobadilla and P. T. K. Woo. An enzyme-linked immunosorbent assay (ELISA) for the detection of antibodies against the pathogenic haemoflagellate, Cryptobia salmositica Katz, and protection against cryptobiosis in juvenile rainbow trout, Oncorhynchus mykiss (Walbaum), inoculated with a live vaccine. Journal of Fish Diseases, 17(4):399–408, July 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stoffregen:1993:STF

[SCBB93] D. A. Stoffregen, A. J. Chako, S. Backman, and J. G. Babish. Successful therapy of furunculosis in Atlantic salmon, Salmo salar L., using the fluoroquinolone antimicrobial agent enrofloxacin. Journal of Fish Diseases, 16(3):219–228, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sharples:1994:FEF

[SCE94] A. D. Sharples, D. N. Campin, and C. W. Evans. Fin erosion in a feral population of goldfish, Carassius auratus (L.), exposed to bleached kraft mill effluent. Journal of Fish Diseases, 17(5):483–493, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1993:ECE

[SD93] P. Smith and S. Davey. Evidence for the competitive exclusion of Aeromonas salmonicida from fish with stress-inducible furunculosis by a fluorescent pseudomonad. Journal of Fish Diseases, 16(5):521–524, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saulnier:1997:PCR

[SD97a] D. Saulnier and P. De Kinkelin. Polymerase chain reaction primers for investigations on the causative agent of proliferative kidney disease of salmonids. Journal of Fish Diseases, 20 (6):467–470, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Strand:1997:AIG

[SD97b] H. K. Strand and R. A. Dalmo. Absorption of immunomodulating β (1,3)-glucan in yolk sac larvae of Atlantic halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 20 (1):41–49, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Svendsen:1999:TLA

[SDB99] Y. S. Svendsen, R. A. Dalmo, and J. Bøgwald. Tissue localization of Aeromonas salmonicida in Atlantic salmon, Salmo salar L., following experimental challenge. Journal of Fish Diseases, 22(2):125–131, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1998:LSA

[SDM^{+}98] D. J. Speare, J. Daley, R. J. F. Markham, J. Sheppard, H. J. Beaman, and J. G. Sanchez. Loma salmonae-associated growth rate suppression in rainbow trout, Oncorhynchus mykiss (Walbaum), occurs during early onset xenoma dissolution as determined by in situ hybridization and immunohistochemistry. Journal of Fish Diseases, 21(5):345–354, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scotland:1990:FAS

[SDS90] L. M. Scotland, P. Dobos, and R. M. W. Stevenson. Fry age and size effects on immersion immunization of brook trout, Salvelinus fontinalis Mitchell, against infectious pancreatic necrosis virus. Journal of Fish Diseases, 13(2):113– 125, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scotland:1995:IAB

[SDS95] L. M. Scotland, P. Dobos, and R. M. W. Stevenson. Immunization of adult brook trout, Salvelinus fontinalis (Mitchill), fails to prevent the infectious pancreatic necrosis virus (IPNV) carrier state. Journal of Fish Diseases, 18(5):449– 458, September 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samuelsen:1999:SDP

[SE99] O. B. Samuelsen and A. Ervik. A single-dose pharmacokinetic study of oxolinic acid and vetoquinol, an oxolinic acid ester, in Atlantic halibut, Hippoglossus hippoglossus L., held in sea water at 9^{o}C. Journal of Fish Diseases, 22(1):13–23, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simko:1999:AIZ

[SEMB^{+}99] E. Simko, A. El-Mowafi, W. J. Bettger, V. E. Ostland, H. W. Ferguson, and M. A. Hayes. Alterations in iron, zinc and major plasma proteins of rainbow trout, Oncorhynchus mykiss (Walbaum), and brook trout, Salvelinus fontinalis (Mitchill), with sterile peritonitis induced by oil-adjuvanted multivalent bacterin vaccination. Journal of Fish Diseases, 22(2):81–90, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1991:PBGb

[SFB^{+}91a] D. J. Speare, H. W. Ferguson, F. W. M. Beamish, J. A. Yager, and S. Yamashiro. Pathology of bacterial gill disease: sequential development of lesions during natural outbreaks of disease. Journal of Fish Diseases, 14(1):21–32, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:1991:PBGa

[SFB^{+}91b] D. J. Speare, H. W. Ferguson, F. W. M. Beamish, J. A. Yager, and S. Yamashiro. Pathology of bacterial gill disease: ultrastructure of branchial lesions. Journal of Fish Diseases, 14 (1):1–20, January 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1993:PPL

[SGNG93] R. Salte, H. M. Gjøen, K. Norberg, and T. Gjedrem. Plasma protein levels as potential marker traits for resistance to furunculosis. Journal of Fish Diseases, 16(6):561–568, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spanggaard:1994:GFP

[SGOH94] B. Spanggaard, L. Gram, N. Okamoto, and H. H. Huss. Growth of the fish-pathogenic fungus, Ichthyophonus hoferi, measured by conductimetry and microscopy. Journal of Fish Diseases, 17(2):145–153, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sandvik:1992:BIA

[SH92] O. Sandvik and A.-B. Hvaal. Bacterial inhibitors affecting proteases produced by Aeromonas salmonicida, and the occurrence of such inhibitors in furunculosis vaccines. Journal of Fish Diseases, 15(2):131–141, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spanggaard:1995:MIH

[SHB95] B. Spanggaard, H. H. Huss, and J. Bresciani. Morphology of Ichthyophonus hoferi assessed by light and scanning electron microscopy. Journal of Fish Diseases, 18(6):567–577, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sterud:1998:IGS

[SHB98] Sterud, Harris, and Bakke. The influence of Gyrodactylus salaris Malmberg 1957 (Monogenea) on the epidermis of Atlantic salmon, Salmo salar L., and brook trout, Salvelinus fontinalis (Mitchill): experimental studies. Journal of Fish Diseases, 21(4):257–263, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stangeland:1996:EII

[SHT96] K. Stangeland, S. Høie, and T. Taksdal. Experimental induction of infectious pancreatic necrosis in Atlantic salmon, Salmo salar L., post-smolts. Journal of Fish Diseases, 19 (4):323–327, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommerville:1991:PIM

[SI91] C. Sommerville and N. A. M. Iqbal. The process of infection, migration, growth and development of Sanguinicola inermis Plehn, 1905 (Digenea: Sanguinicolidae) in carp, Cyprinus carpio L. Journal of Fish Diseases, 14(2):211–219, March 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Steiro:1998:OCC

[SJGB98] K. Steiro, A. Johansen, A. Gildberg, and J. Bøgwald. Optimising of culture conditions and stimulation of head kidney macrophages from Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 21(5):335–344, September 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stensvaag:1993:PPC

[SJH^{+}93] K. Stensvåg, T. Ø. Jørgensen, J. Hoffman, K. Hjelmeland, and J. Bøgwald. Partial purification and characterization of extracellular metalloproteases with caseinolytic, aminopeptidolytic and collagenolytic activities from Vibrio anguillarum. Journal of Fish Diseases, 16(6):525–539, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saint-Jean:1991:DIP

[SJMPP91] S. Rodriguez Saint-Jean, Ma. Pilar Vilas Minondo, M. Angel Palacios, and S. Perez Prieto. Detection of infectious pancreatic necrosis in a carrier population of rainbow trout, Oncorhynchus mykiss (Richardson), by flow cytometry. Journal of Fish Diseases, 14(5):545–553, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stadtlander:1990:GOC

[SK90] C. Stadtländer and H. Kirchhoff. Gill organ culture of rainbow trout, Salmo gairdneri Richardson: an experimental model for the study of pathogenicity of Mycoplasma mobile 163 K. Journal of Fish Diseases, 13(1):79–86, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shoemaker:1997:PIA

[SK97] C. A. Shoemaker and P. H. Klesius. Protective immunity against enteric septicaemia in channel catfish, Ictalurus punctatus (Rafinesque), following controlled exposure to Edwardsiella ictaluri. Journal of Fish Diseases, 20(5):361–368, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shaw:1999:ITC

[SKA99] R. W. Shaw, M. L. Kent, and M. L. Adamson. Iodophor treatment is not completely efficacious in preventing Loma salmonae (Microsporidia) transmission in experimentally challenged Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 22(4):311–313, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Steinhagen:1990:SHO

[SKK90] D. Steinhagen, P. Kruse, and W. Körting. Some haematological observations on carp, Cyprinus carpio L., experimentally infected with Trypanoplasma borreli Laveran & Mesnil, 1901 (Protozoa: Kinetoplastida). Journal of Fish Diseases, 13(2): 157–162, March 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simko:1999:IAS

[SKQ^{+}99] E. Simko, T. E. Kocal, B. A. Quinn, V. E. Ostland, H. W. Ferguson, and M. A. Hayes. Influences of Aeromonas salmonicida lipopolysaccharide, prednisolone and water temperature on plasma protein composition in salmonids. Journal of Fish Diseases, 22(2):91–100, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samuelsen:1997:PES

[SLJ97] O. B. Samuelsen, B. T. Lunestad, and A. Jelmert. Pharmacokinetic and efficacy studies on bath-administering potentiated sulphonamides in Atlantic halibut, Hippoglossus hippoglossus L. Journal of Fish Diseases, 20(4):287–296, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sin:1996:CMI

[SLL96] Y. M. Sin, K. H. Ling, and T. J. Lam. Cell-mediated immune response of goldfish, Carassius aurafus (L.), to Ichthyophthirius multifiliis. Journal of Fish Diseases, 19(1):1–7, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommer:1996:PIS

[SM96] A.-I. Sommer and S. Mennen. Propagation of infectious salmon anaemia virus in Atlantic saimon, Salmo salar L., head kidney macrophages. Journal of Fish Diseases, 19(2): 179–183, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:1995:PIS

[SMBM95] D. A. Smail, L. Mcfarlane, D. W. Bruno, and A. H. Mcvicar. The pathology of an IPN-Sp sub-type (Sh) in farmed Atlantic salmon, Salmo solar L., post-smolts in the Shetland Isles, Scotland. Journal of Fish Diseases, 18(6):631–638, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sima:1996:PEE

[SMG96] A. R. Simá, F. Mmoravec, and R. J. Güemez. Pathogenic effects of the eggs of Capillaria cichlasomae (Nematoda: Capillariidea) in the liver of the cichlid Cichlasoma urophthalmus (Günther) in Mexico. Journal of Fish Diseases, 19(2):175– 178, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1993:HCS

[SMHS93] N. Sano, M. Moriwake, R. Hondo, and T. Sano. Herpesvirus cyprini: a search for viral genome in infected fish by infected fish by in situ hybridization. Journal of Fish Diseases, 16(5): 495–499, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1991:HCL

[SMSA91] T. Sano, N. Morita, N. Shima, and M. Akimoto. Herpesvirus cyprini: lethality and oncogenicity. Journal of Fish Diseases, 14(5):533–543, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1991:DIC

[SN91] R. Salte and K. Norberg. Disseminated intravascular coagulation in farmed Atlantic salmon, Salmo salar L.: evidence of consumptive coagulopathy. Journal of Fish Diseases, 14 (4):475–480, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1992:SPG

[SNA^{+}92] R. Salte, K. Norberg, J. A. Arnesen, O. R. Ødegaard, and G. Eggset. Serine protease and glycerophospholipid:cholesterol acyltransferase of Aeromonas salmonicida work in concert in thrombus formation; in vitro the process is counteracted by plasma antithrombin and α 2-macroglobulin. Journal of Fish Diseases, 15(3):215–227, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1991:DEP

[SNØ91] R. Salte, K. Norberg, and O. R. Ødegaard. Do extracellular products of Aeromonas salmonicida induce thrombosis by entering the fish coagulation system at factor X? Journal of Fish Diseases, 14(3):401–406, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1993:EIC

[SNØ^{+}93] R. Salte, K. Norberg, O. R. Ødegaard, J. A. Arnesen, and J. J. Olli. Exotoxin-induced consumptive coagulopathy in Atlantic salmon, Salmo salar L.: inhibitory effects of exogenous antithrombin and α_{2}-macroglobulin on Aeromonas salmonicida serine protease. Journal of Fish Diseases, 16(5):425–435, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1995:EOC

[SNó95] R. Salte, K. Norberg, and O. R. óDegaard. Evidence of ongoing coagulation and fibrinolysis in early-maturing Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 18(6):623– 626, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1993:ERB

[SOAK93] M. Sakai, T. Otubo, S. Atsuta, and M. Kobayashi. Enhancement of resistance to bacterial infection in rainbow trout, Oncorhynchus mykiss (Walbaum), by oral administration of bovine lactoferrin. Journal of Fish Diseases, 16(3):239–247, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1992:VIP

[SOF^{+}92] M. Sano, N. Okamoto, H. Fukuda, M. Saneyoshi, and T. Sano. Virulence of infectious pancreatic necrosis virus is associated with the larger RNA segment (RNA segment A). Journal of Fish Diseases, 15(4):283–293, July 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sveinbjornsson:1996:ILL

[SOP96] B. Sveinbjornsson, R. Olsen, and S. Paulsen. Immunocytochemical localization of lysozyme in intestinal eosinophilic granule cells (EGCs) of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 19(5):349–355, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1994:IPN

[SOS94] M. Sano, N. Okamoto, and T. Sano. Infectious pancreatic necrosis virus plaque size depends on the larger RNA segment and is not linked to virulence. Journal of Fish Diseases, 17(6): 657–659, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sovenyi:1992:EFF

[Söv92] J. F. Sövényi. Effect of fumagillin feeding on the haematopoietic organs and susceptibility of common carp, Cyprinus carpio L., to erythrodermatitis. Journal of Fish Diseases, 15 (3):255–260, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shafer:1990:EES

[SPC90] T. H. Shafer, D. E. Padgett, and D. A. Celio. Evidence for enhanced salinity tolerance of a suspected fungal pathogen of Atlantic menhaden, Brevoortia tyrannus Latrobe. Journal of Fish Diseases, 13(5):335–344, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skurdal:1990:YST

[SQTF90] J. Skurdal, T. Qvenild, T. Taugbøl, and E. Fjeld. A 6-year study of Thelohania contejeani parasitism of the noble crayfish, Astacus astacus L., in Lake Steinsfjorden, S.E. Norway. Journal of Fish Diseases, 13(5):411–415, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skliris:1999:NIE

[SR99] G. P. Skliris and R. H. Richards. Nodavirus isolated from experimentally infected tilapia, Oreochromis mossambicus (Peters). Journal of Fish Diseases, 22(4):315–318, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sousa:1996:HST

[SRL^{+}96] J. A. Sousa, J. L. Romalde, A. Ledo, J. C. Eiras, J. L. Barja, and A. E. Toranzo. Health status of two salmonid aquaculture facilities in North Portugal: characterization of the bacterial and viral pathogens causing notifiable diseases. Journal of Fish Diseases, 19(1):83–89, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Salte:1994:WUS

[SRRN94] R. Salte, K.-A. Rørvik, E. Reed, and K. Norberg. Winter ulcers of the skin in Atlantic salmon, Salmo salar L.: pathogenesis and possible aetiology. Journal of Fish Diseases, 17 (6):661–665, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shchelkunov:1990:IEC

[SS90] I. S. Shchelkunov and T. I. Shchelkunova. Infectivity experiments with Cyprinus carpio iridovirus (CCIV), a virus unassociated with carp gill necrosis. Journal of Fish Diseases, 13 (6):475–484, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sorensen:1997:ICC

[SSD^{+}97] K. K. Sørensen, B. Sveinbjørnsson, R. A. Dalmo, B. Smedsrød, and K. Bertheussen. Isolation, cultivation and characterization of head kidney macrophages from Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 20(2):93–107, March 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanchez:1997:MHA

[SSJ97] J. G. Sanchez, D. J. Speare, and G. J. Johnson. Morphometric and histochemical assessment of the branchial tissue response of rainbow trout, Oncorhynchus mykiss (Walbaum), associated with chloramine-T treatment. Journal of Fish Diseases, 20(5):375–381, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soltani:1995:CCF

[SSM95] M. Soltani, S. Shanker, and B. L. Munday. Chemotherapy of Cytophaga/Flexibacter-like bacteria (CFLB) infections in fish: studies validating clinical efficacies of selected antimicrobials. Journal of Fish Diseases, 18(6):555–565, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stone:1999:EEB

[SSS^{+}99] J. Stone, I. H. Sutherland, C. S. Sommerville, R. H. Richards, and K. J. Varma. The efficacy of emamectin benzoate as an oral treatment of sea lice, Lepeophtheirus salmonis (Krøyer), infestations in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 22(4):261–270, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1992:HCD

[SSSH92] N. Sano, M. Sano, T. Sano, and R. Hondo. Herpesvirus cyprini: detection of the viral genome by in situ hybridization. Journal of Fish Diseases, 15(2):153–162, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spisni:1998:HSA

[STP^{+}98] E. Spisni, M. Tugnoli, A. Ponticelli, T. Mordenti, and V. Tomasi. Hepatic steatosis in artificially fed marine teleosts. Journal of Fish Diseases, 21(3):177–184, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Subasinghe:1993:ECI

[Sub93] R. P. Subasinghe. Effects of controlled infections of Trichodina sp. on transmission of epizootic ulcerative syndrome (EUS) to naive snakehead, Ophicephalus striatus Bloch. Journal of Fish Diseases, 16(2):161–164, March 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1990:LTG

[SW90] A. C. Smith and T.-Y. Wang. A large tumour from the gray snapper, Lutjanus griseus (L.). Journal of Fish Diseases, 13 (4):311–316, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shankar:1994:PPI

[SY94] K. M. Shankar and T. Yamamoto. Prevalence and pathogenicity of infectious pancreatic necrosis virus (IPNV) associated with feral lake trout, Salvelinus namaycush (Walbaum). Journal of Fish Diseases, 17(5):461–471, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1995:ERV

[SYAK95] M. Sakai, T. Yoshida, S. Atsuta, and M. Kobayashi. Enhancement of resistance to vibriosis in rainbow trout, Oncorhynchus mykiss (Walbaum), by oral administration of Clostridium butyricum bacterin. Journal of Fish Diseases, 18(2):187–190, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thiery:1999:TIS

[TAD99] R. Thiéry, C. Arnauld, and C. Delsert. Two isolates of sea bass, Dicentrarchus labrax L., nervous necrosis virus with distinct genomes. Journal of Fish Diseases, 22(3):201–207, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taraschewski:1990:HPI

[Tar90] H. Taraschewski. Host-parasite interface of Neoechinorhynchus rutili (Eoacanthocephala) in naturally infected salmonids. Journal of Fish Diseases, 13(1):39–48, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tatner:1991:MEP

[Tat91] M. F. Tatner. Modified extracellular product antigens of Aeromonas salmonicida as potential vaccines for the control of furunculosis in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 14(3):395–400, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thornton:1994:DLV

[TGK94] J. C. Thornton, R. A. Garduño, and W. W. Kay. The development of live vaccines for furunculosis lacking the A-layer and O-antigen of Aeromonas salmonicida. Journal of Fish Diseases, 17(3):195–204, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tarrab:1996:ADE

[THBL96] E. Tarrab, J. Heppell, L. Berthiaume, and J. Lecomte. Antigenic diversity of eastern Canadian isolates of infectious pancreatic necrosis virus. Journal of Fish Diseases, 19(1):23–31, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thoney:1990:ETP

[Tho90] D. A. Thoney. The effects of trichlorfon, praziquantel and copper sulphate on various stages of the monogenean Benedeniella posterocolpa, a skin parasite of the cownose ray, Rhinoptera bonasus (Mitchill). Journal of Fish Diseases, 13 (5):385–389, September 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Takahashi:1996:PCR

[TIM^{+}96] Y. Takahashi, T. Itami, M. Maeda, N. Suzuki, J. Kasornchandra, K. Supamattaya, R. Khongpradit, S. Boonyaratpalin, M. Kondo, K. Kawai, R. Kusuda, I. Hirono, and T. Aoki. Polymerase chain reaction (PCR) amplification of bacilliform virus (RV-PJ) DNA in Penaeus japonicus Bate and systemic ectodermal and mesodermal baculovirus (SEMBV) DNA in Penaeus monodon Fabricius. Journal of Fish Diseases, 19(5): 399–403, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Treasurer:1994:ASI

[TL94] J. W. Treasurer and L. A. Laidler. Aeromonas salmonicida infection in wrasse (Labridae), used as cleaner fish, on an Atlantic salmon, Salmo salar L., farm. Journal of Fish Diseases, 17(2):155–161, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Turnbull:1994:HDS

[TLMM94] J. F. Turnbull, P. E. Larkins, C. Mcpadden, and R. Matondang. A histopathological disease survey of cultured shrimp in North East Sumatra, Indonesia. Journal of Fish Diseases, 17(1):57–65, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taveekijakarn:1996:SVK

[TMMA96a] P. Taveekijakarn, T. Miyazaki, M. Matsumoto, and S. Arai. Studies on vitamin K deficiency in amago salmon, Oncorhynchus rhodurus (Jordan & McGregor). Journal of Fish Diseases, 19(3):209–214, May 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taveekijakarn:1996:HHC

[TMMA96b] P. Taveekijakarn, T. Miyazaki, M. Matsumoto, and S. Aral. Histopathological and haematological changes in amago salmon, Oncorhynchus rhodurus (Jordan & McGregor), fed a vitamin-D-free diet. Journal of Fish Diseases, 19(4):289– 294, July 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tafalla:1999:VVE

[TNAF99] C. Tafalla, B. Novoa, J. M. Alvarez, and A. Figueras. In vivo and in vitro effect of oxytetracycline treatment on the immune response of turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 22(4):271–276, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Toranzo:1993:HRT

[TNB^{+}93] A. E. Toranzo, B. Novoa, A. M. Baya, F. M. Hetrick, J. L. Barja, and A. Figueras. Histopathology of rainbow trout, Oncorhynchus mykiss (Walbaum), and striped bass, Morone saxatilis (Walbaum), experimentally infected with Carnobacterium piscicola. Journal of Fish Diseases, 16(3):261–267, May 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Takahashi:1994:RBF

[TOK^{+}94] K. Takahashi, N. Okamoto, A. Kumagai, M. Maita, J. S. Rohovec, and Y. Ikeda. Relationship between fish growth rate and progression of artificially induced erythrocytic inclusion body syndrome in Coho salmon, Oncorhynchus kisutch (Walbaum). Journal of Fish Diseases, 17(1):77–83, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Turnbull:1996:GHS

[TRR96] J. F. Turnbull, R. H. Richards, and D. A. Robertson. Gross, histological and scanning electron microscopic appearance of dorsal fin rot in farmed Atlantic salmon, Salmo salar L., parr. Journal of Fish Diseases, 19(6):415–427, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taksdal:1998:IIP

[TRSD98] T. Taksdal, A. Ramstad, K. Stangeland, and B. H. Dannevig. Induction of infectious pancreatic necrosis (IPN) in covertly infected Atlantic salmon, Salmo salar L., post-smolts by stress exposure, by injection of IPN virus (IPNV) and by cohabitation. Journal of Fish Diseases, 21(3):193–204, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tojo:1994:PTA

[TSL^{+}94] J. L. Tojo, M. T. Santamarina, J. Leiro, F. M. Ubeira, and M. L. Sanmartin. Pharmacological treatments against Ichthyobodo necator (Henneguy, 1883) in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 17 (2):135–143, March 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taksdal:1999:ERC

[TT99] T. Taksdal and K. Thorud. Evaluation of a rapid coagglutination (COA) test for the detection of infectious pancreatic necrosis virus (IPNV) in tissue samples of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 22(2):117– 124, March 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1990:DMH

[TW90a] P. T. Thomas and P. T. K. Woo. Dietary modulation of humoral immune response and anaemia in rainbow trout, Oncorhynchus mykiss (Walbaum), infected with Cryptobia salmositica Katz, 1951. Journal of Fish Diseases, 13(6):435– 446, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1990:VVC

[TW90b] P. T. Thomas and P. T. K. Woo. In vivo and in vitro cellmediated immune responses of rainbow trout, Oncorhynchus mykiss (Walbaum), against Cryptobia salmositica Katz, 1951 (Sarcomastigophora: Kinetoplastida). Journal of Fish Diseases, 13(6):423–433, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1992:ART

[TW92] P. T. Thomas and P. T. K. Woo. Anorexia in rainbow trout, Oncorhynchus mykiss (Walbaum), infected with Cryptobia salmositica (Sarcomastigophora: Kinetoplastida): its onset and contribution to the immunodepression. Journal of Fish Diseases, 15(5):443–447, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tung:1999:HEM

[TWC99] C. W. Tung, C. S. Wang, and S. N. Chen. Histological and electron microscopic study on Macrobrachium muscle virus (MMV) infection in the giant freshwater prawn, Macrobrachium rosenbergii (de Man), cultured in Taiwan. Journal of Fish Diseases, 22(4):319–323, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Urawa:1991:HMC

[UUNY91] S. Urawa, N. Ueki, T. Nakai, and H. Yamasaki. High mortality of cultured juvenile Japanese flounder, Paralichthys olivaceus (Temminek & Sehlegel), eaused by the parasitic flagellate Ichthyobodo sp. Journal of Fish Diseases, 14(4):489–494, July 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Voronin:1993:IGP

[VC93] V. N. Voronin and N. B. Chernysheva. An intracellular gill parasite as the possible causative agent of mortality during swim-bladder inflammation in common carp, Cyprinus carpio L. Journal of Fish Diseases, 16(6):609–611, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valdron:1994:PNL

[VCC94] K. A. Valdron, D. K. Cone, and R. Cusack. Persistent nodular lesions of bacterial origin on the gills of gasterosteid fish inhabiting an estuarine bay in Nova Scotia, Canada. Journal of Fish Diseases, 17(6):673–679, November 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vaagsholm:1998:RFSa

[VD98a] Vågsholm and Djupvik. Risk factors for spinal deformities in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21 (1):47–53, January 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vaagsholm:1998:RFSb

[VD98b] I. Vågsholm and H. O. Djupvik. Risk factors for skin lesions in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 21(6):449–454, November 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vaagsholm:1999:RFA

[VD99] I. Vågsholm and H. O. Djupvik. Risk factors for abdominal adhesions in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 22(1):53–58, January 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

vandenIngh:1996:ASC

[vdIOK96] T. S. G. A. M. van den Ingh, J. J. Olli, and Å. Krogdahl. Alcohol-soluble components in soybeans cause morphological changes in the distal intestine of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 19(1):47–53, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vogelbein:1999:HMF

[VFCV99] W. K. Vogelbein, J. W. Fournie, P. S. Cooper, and P. A. Van Veld. Hepatoblastomas in the mummichog, Fundulus heteroclitus (L.), from a creosote-contaminated environment: a histologic, ultrastructural and immunohistochemical study. Journal of Fish Diseases, 22(6):419–431, December 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Velez-Hernandez:1998:SCL

[VHCCGMOS98] E. M. Vélez-Hernández, F. Constantino-Casas, L. J. García-Márquez, and D. Osorio-Sarabia. Short communicationGill lesions in common carp, Cyprinus carpio L., in Mexico due to the metacercariae of Centrocestus formosanus. Journal of Fish Diseases, 21(3):229–232, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Virgona:1992:EFI

[Vir92] J. L. Virgona. Environmental factors influencing the prevalence of a cutaneous ulcerative disease (red spot) in the sea mullet, Mugil cephalus L., in the Clarence River, New South Wales, Australia. Journal of Fish Diseases, 15(5):363–378, September 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vickers:1993:ISM

[VPSL93] J. E. Vickers, J. L. Paynter, P. B. Spradbrow, and R. J. G. Lester. An impression smear method for rapid detection of Penaeus monodon-type baculovirus (MBV) in Australian prawns. Journal of Fish Diseases, 16(5):507–511, September 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valtonen:1992:OPY

[VRK92] E. Tellervo Valtonen, P. Rintamäki, and M. Koskivaara. Occurrence and pathogenicity of Yersinia ruckeri at fish farms in northern and central Finland. Journal of Fish Diseases, 15 (2):163–171, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Verity:1996:CMA

[VW96] C. K. Verity and P. T. K. Woo. Characterization of a monoclonal antibody against the 47- kDa antigen of Cryptobia salmositica Katz and its use in an antigen-capture enzymelinked immunosorbent assay for detection of parasite antigen in infected rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 19(1):91–109, January 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wakabayashi:1991:EEC

[Wak91] H. Wakabayashi. Effect of environmental conditions on the infectivity of Flexibacter columnaris to fish. Journal of Fish Diseases, 14(3):279–290, May 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:1994:CTD

[WB94] R. Wang and M. Belosevic. Cultivation of Trypanosoma danilewskyi (Laveran & Mesnil, 1904) in serum-free medium and assessment of the course of infection in goldfish, Carassius auratus (L.). Journal of Fish Diseases, 17(1):47–56, January 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waterstrat:1991:UEB

[WBA91] P. R. Waterstrat, J. Brazil, and A. J. Ainsworth. Use of an ELISA-based assay for the detection of antibody-secreting cells in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 14(6):669–675, November 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waagbo:1996:NSA

[WBS^{+}96] R. Waagbø, E. Bjerkås, H. Sveier, O. Breck, E. Bjørnestad, and A. Maage. Nutritional status assessed in groups off smolting Atlantic salmon, Salmo salar L., developing cataracts. Journal of Fish Diseases, 19(5):365–373, September 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whyte:1990:PRT

[WCS90] S. K. Whyte, L. H. Chappell, and C. J. Secombes. Protection of rainbow trout, Oncorhynchus mykiss (Richardson), against Diplostomum spathaceum (Digenea): the role of specific antibody and activated macrophages. Journal of Fish Diseases, 13(4):281–291, July 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whitby:1992:OKS

[WDCM92] P. W. Whitby, S. G. Delany, G. Coleman, and A. L. S. Munro. The occurrence of a 70-kDa serine protease gene in typical and atypical strains of Aeromonas salmonicida. Journal of Fish Diseases, 15(6):529–535, November 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiegertjes:1993:DRC

[WDV93] G. F. Wiegertjes, J. G. Daly, and W. B. Van Muiswinkel. Disease resistance of carp, Cyprinus carpio L.: identification of individual genetic differences by bath challenge with atypical Aeromonas salmonicida. Journal of Fish Diseases, 16(6): 569–576, November 1993. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wahli:1995:CES

[WFS^{+}95] T. Wahli, R. Frischknecht, M. Schmitt, J. Gabaudan, V. Verlhac, and W. Meier. A comparison of the effect of silicone coated ascorbic acid and ascorbyl phosphate on the course of ichthyophthiriosis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 18(4):347–355, July 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:1998:GCA

[WGFF98] Wolf, Ginn, and Francis-Floyd. Goitre in a colony of African cichlids. Journal of Fish Diseases, 21(2):139–143, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waagbo:1994:IDC

[WGSH94] R. Waagbø, J. Glette, K. Sandnes, and G. I. Hemre. Influence of dietary carbohydrate on blood chemistry, immunity and disease resistance in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 17(3):245–258, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:1998:ISP

[WHC^{+}98] J-L. Wu, J-R. Hong, C-Y. Chang, C-F. Hui, C-F. Liao, and Y-L. Hsu. Involvement of serine proteinase in infectious pancreatic necrosis virus capsid protein maturation and NS proteinase cleavage in CHSE-214 cells. Journal of Fish Diseases, 21(3):215–220, May 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waagbo:1995:TFA

[WHHL95] R. Waagbó, G. I. Hemre, J. Chr. Holm, and Ó. Lie. Tissue fatty acid composition, haematology and immunity in adult cod, Gadus morhua L., fed three dietary lipid sources. Journal of Fish Diseases, 18(6):615–622, November 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiklund:1990:AAS

[Wik90] T. Wiklund. Atypical Aeromonas salmonicida isolated from ulcers of pike, Esox lucius L. Journal of Fish Diseases, 13(6): 541–544, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wagner:1997:DCM

[WLHD97] U. Wagner, I. Lachmann, D. Hädge, and K. Drössler. Development and characterization of monoclonal antibodies specific for two different exoproteases of Aeromonas salmonicida. Journal of Fish Diseases, 20(3):223–235, May 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:1994:EEH

[WPRM94] R. J. Whittington, A. Philbey, G. L. Reddacliff, and A. R. Macgown. Epidemiology of epizootic haematopoietic necrosis virus (EHNV) infection in farmed rainbow trout, Oncorhynchus mykiss (Walbaum): findings based on virus isolation, antigen capture ELISA and serology. Journal of Fish Diseases, 17(3):205–218, May 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1992:TSU

[WR92] L. G. Willoughby and R. J. Roberts. Towards strategic use of fungicides against Saprolegnia parasitica in salmonid fish hatcheries. Journal of Fish Diseases, 15(1):1–13, January 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1994:IMI

[WR94] L. G. Willoughby and R. J. Roberts. Improved methodology for isolation of the Aphanomyces fungal pathogen of epizootic ulcerative syndrome (EUS) in Asian fish. Journal of Fish Diseases, 17(5):541–543, September 1994. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

West:1998:ELF

[WR98] West and Roubal. Experiments on the longevity, fecundity and migration of Anoplodiscus cirrusspiralis (Monogenea) on the marine fish Pagrus auratus (Bloch & Schneider) (Sparidae). Journal of Fish Diseases, 21(4):299–303, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1995:AIS

[WRC95] L. G. Willoughby, R. J. Roberts, and S. Chinabut. Aphanomyces invaderis sp. nov., the fungal pathogen of freshwater tropical fish affected by epizootic ulcerative syndrome. Journal of Fish Diseases, 18(3):273–276, May 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woo:1990:LCC

[WS90] P. T. K. Woo and M. Shariff. Lernea cyprinacea L. (Copepoda: Caligidea) in Helostoma temmincki Cuvier & Valenciennes: the dynamics of resistance in recovered and naive fish. Journal of Fish Diseases, 13(6):485–493, November 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Williams:1996:CRT

[WST96] J. Williams, P. Stewart, and P. Timpson. Characterization of rainbow trout, Oncorhynchus mykiss (Walbaum), serum creatine kinase isoenzymes and isoforms by means of electrophoresis and isoelectric focusing. Journal of Fish Diseases, 19(6):441–447, November 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woo:1992:CVS

[WT92] P. T. K. Woo and P. T. Thomas. Comparative in vitro studies on virulent and avirulent strains of Cryptobia salmositica (Sarcomastigophora: Kinetoplastida). Journal of Fish Diseases, 15(3):261–266, May 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:1997:LEM

[WTKC97] C. S. Wang, K. F. J. Tang, G. H. Kou, and S. N. Chen. Light and electron microscopic evidence of white spot disease in the giant tiger shrimp, Penaeus monodon (Fabricius), and the kuruma shrimp, Penaeus japonicus (Bate), cultured in Taiwan. Journal of Fish Diseases, 20(5):323–331, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wahli:1998:ICV

[WVG^{+}98] Wahli, Verlhac, Gabaudan, Schüep, and Meier. Influence of combined vitamins C and E on non-specific immunity and disease resistance of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 21(2):127–137, March 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:1995:EPC

[YA95] C. Z. Yang and L. J. Albright. Elevation of plasma cortisol and hypertrophic response of inter-renal and chromaffin tissues of rainbow trout, Oncorhynchus mykiss (Walbaum), to the harmful diatom Chaetoceros concavicornis. Journal of Fish Diseases, 18(2):165–174, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yokoyama:1997:VST

[YDOW97] H. Yokoyama, T. Danjo, K. Ogawa, and H. Wakabayashi. A vital staining technique with fluorescein diacetate (FDA) and propidium iodide (PI) for the determination of viability of myxosporean and actinosporean spores. Journal of Fish Diseases, 20(4):281–286, July 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoshikoshi:1990:VNN

[YI90] K. Yoshikoshi and K. Inoue. Viral nervous necrosis in hatchery-reared larvae and juveniles of Japanese parrotfish, Oplegnathus fasciatus (Temminck & Schlegel). Journal of Fish Diseases, 13(1):69–77, January 1990. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yadav:1992:CEA

[YIA92] M. Yadav, G. Indira, and A. Ansary. Cytotoxin elaboration by Aeromonas hydrophila isolated from fish with epizootic ulcerative syndrome. Journal of Fish Diseases, 15(2):183– 189, March 1992. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoshida:1997:VAP

[YIM^{+}97] T. Yoshida, V. Inglis, N. Misawa, R. Kruger, and M. Sakai. In vitro adhesion of Pasteurella piscicida to cultured fish cells. Journal of Fish Diseases, 20(1):77–80, January 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoshida:1995:ANS

[YKI95] T. Yoshida, R. Kruger, and V. Inglis. Augmentation of non-specific protection in African catfish, Clarias gariepinus (Burchell), by the long-term oral administration of immunostimulants. Journal of Fish Diseases, 18(2):195–198, March 1995. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yokoyama:1999:EFA

[YLSW99] H. Yokoyama, Y. S. Liyanage, A. Sugai, and H. Wakabayashi. Efficacy of fumagillin against haemorrhagic thelohanellosis caused by Thelohanellus hovorkai (Myxosporea: Myxozoa) in coloured carp, Cyprinus carpio L. Journal of Fish Diseases, 22(3):243–245, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yano:1991:PIP

[YMM91] T. Yano, H. Matsuyama, and R. E. P. Mangindaan. Polysaccharide-induced protection of carp, Cyprinus carpio L., against bacterial infection. Journal of Fish Diseases, 14 (5):577–582, September 1991. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zorrilla:1999:ICC

[ZBM^{+}99] I. Zorrilla, M. C. Balebona, M. A. Mori·igo, C. Sarasquete, and J. J. Borrego. Isolation and characterization of the causative agent of pasteurellosis, Photobacterium damsela ssp. piscicida, from sole, Solea senegalensis (Kaup). Journal of Fish Diseases, 22(3):167–172, May 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zuo:1997:VIP

[ZFW97] X. Zuo, S. Feng, and P. T. K. Woo. The in vitro inhibition of proteases from Cryptobia salmositica Katz by a monoclonal antibody (MAb-001) against a glycoprotein on the pathogenic haemoflagellate. Journal of Fish Diseases, 20(6): 419–426, November 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:1996:RCC

[ZH96] H. G. Zhang and L. A. Hanson. Recombinant channel catfish virus (Ictalurid herpesvirus 1) can express foreign genes and induce antibody production against the gene product. Journal of Fish Diseases, 19(2):121–128, March 1996. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zintl:1997:NOT

[ZPVH97] A. Zintl, W. R. Poole, H. P. Voorheis, and C. V. Holland. Naturally occurring Trypanosoma granulosum infections in the European eel, Anguilla anguilla L. from County Mayo, western Ireland. Journal of Fish Diseases, 20(5):333–341, September 1997. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zintl:1999:CIB

[ZPVH99] A. Zintl, W. R. Poole, H. P. Voorheis, and C. V. Holland. Cryptobia infections in brown trout, Salmo trutta L., caught as bycatch in the Burrishoole catchment, County Mayo, western Ireland. Journal of Fish Diseases, 22(4):325–329, July 1999. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zuo:1998:VSM

[ZW98] Zuo and Woo. In vitro secretion of metallo-protease (200 kDa) by the pathogenic piscine haemoflagellate, Cryptobia salmositica Katz, and stimulation of protease production by collagen. Journal of Fish Diseases, 21(4):249–255, July 1998. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).